### v1

# Learning Program Behavioral Models from Synthesized Input-Output Pairs

Abstract.
We introduce Modelizer—a novel framework that, given a black-box program, learns a model from its input/output behavior using neural machine translation. The resulting model mocks the original program: Given an input, the model predicts the output that would have been produced by the program. However, the model is also reversible—that is, the model can predict the input that would have produced a given output. Finally, the model is differentiable and can be efficiently restricted to predict only a certain aspect of the program behavior. Modelizer uses grammars to synthesize inputs and to parse the resulting outputs, allowing it to learn sequence-to-sequence associations between token streams. Other than input and output grammars, Modelizer only requires the ability to execute the program. The resulting models are small, requiring fewer than 6.3 million parameters for languages such as Markdown or HTML; and they are accurate, achieving up to 95.4% accuracy and a BLEU score of 0.98 with standard error 0.04 in mocking real-world applications. We foresee several applications of these models, especially as the output of the program can be any aspect of program behavior. Besides mocking and predicting program behavior, the model can also synthesize inputs that are likely to produce a particular behavior, such as failures or coverage.
Software Testing, Mocking, Deep Learning
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Computing methodologies Learning from demonstrationsccs: Software and its engineering Model-driven software engineeringccs: Security and privacy Software reverse engineeringccs: Software and its engineering Software prototypingccs: Software and its engineering Software testing and debugging
1. Introduction
Figure 1. Example of Modelizer in action.
[…]
Example of Modelizer in action.
Understanding and predicting the behavior of a given program is a core challenge of Computer Science. Many programs people use are black-box systems that conceal their internal workings, while allowing users to interact with them by providing inputs and observing outputs. In this work, we present Modelizer—a novel approach to extract a behavior model from a black-box program under test (PUT)—by synthesizing inputs, observing outputs, and learning the relationships between input and output features. The resulting neural machine translation model $M$ can be used for several purposes. First, $M$ is able to mock the behavior of the PUT: Given an input, $M$ predicts the output that the PUT would have produced, with high accuracy (Figure 1). $M$ can then replace the PUT in case the PUT can no longer be executed—for instance, because it requires a special environment, or is expensive to run, or because it may be suspected to contain malware (which will not be extracted into $M$). As the output can be any aspect of program behavior (such as execution traces or coverage), the model $M$ can also predict such aspects for any given input.
The most interesting point, however, is that the learned model $M$ is reversible: it can predict an input that likely produces a given output. Assuming the PUT translates Markdown inputs to HTML, then a reversed model $M^{-1}$ would take HTML inputs and translate them back to the original Markdown code, again with high accuracy. And if the output is, say, a particular execution trace, then the reversed model $M^{-1}$ will produce an input that is likely to produce this very trace. The learned model $M$ is differentiable, and can thus be efficiently restricted to particular output features.
Modelizer learns the model from the PUT by executing it again and again, observing how its inputs translate to its outputs. However, like all machine learning models, Modelizer depends on good training. This raises two central problems, for each of which Modelizer has a solution:
Synthesizing inputs.:
The first central problem thus is the availability of input data for training. In Software Engineering, the field of test generation discusses how to obtain input data for a program. The test data must be valid, such that it gets accepted by the PUT, as well as diverse, such that a wide variety of input features is covered. In a black-box setting, obtaining valid and diverse input data requires a dedicated generator or producer, leveraging knowledge about the input language of the program.
Tokenization.:
The second problem is obtaining features from inputs and outputs. Modelizer uses a sequence-to-sequence translation model to translate token sequences from inputs into outputs. Hence, we need a decomposition of inputs and outputs into individual tokens, so we can learn how token sequences map into each other. While this is a mostly solved problem for natural languages, tokenizing formal inputs correctly is a challenge, notably as we want the tokenizer to be general-purpose, and suitable for all sorts of input and output languages.
Modelizer addresses both problems by leveraging grammars for both input and output language. The input grammar is used to produce syntactically valid and diverse inputs; it can also parse existing inputs into derivation trees. In both cases, we know the structure of the input. The output grammar is used to parse the output, again delivering a tree structure allowing decomposition into tokens. Using the input grammar as a producer, we can synthesize valid and diverse input data automatically and thus obtain as many input-output pairs from the PUT as we want. In cases when output language grammar is unavailable, the program output can be decomposed by unsupervised tokenizers. Modelizer distinguishes structure from content in the decomposed sequences and can abstract content by replacing them with placeholders to improve the learning quality.
Developers who want to use Modelizer thus only need to provide the PUT in an executable form, as well as grammars that describe its input and output languages; such grammars can also be inferred from the PUT or input samples. Modelizer synthesizes inputs for the PUT, eventually extracting its behavior model $M$. The model can then be used for a variety of purposes, including:
  • as a mock for the PUT, predicting its behavior—for instance, for testing or development when the PUT is not available for execution;
  • as a reverse engineering tool—for instance, allowing developers to simulate and analyze the behavior of the PUT without requiring its code; or for plain model stealing, replicating the behavior of an existing model;
  • as an anomaly detector monitoring the PUT and detecting behavior changes—for instance, flagging situations where the PUT and model deviate;
  • as a reverse predictor, producing likely inputs that result in given outputs—for instance, to produce inputs that trigger behaviors of interest; this is the core challenge of software test generation.
These scenarios only scratch the surface of possible usages. In this paper, however, we focus on the mock and reverse predictor scenarios—that is, predicting outputs from inputs and vice versa. We evaluate Modelizer across a range of challenging PUTs and achieve excellent accuracy. Notably, the models produced by Modelizer showed high accuracy on non-trivial programs despite having a small number of neurons (and thus efficient to learn). On the Pandoc markup converter, for instance, Modelizer achieved over 90% accuracy in converting Markdown to HTML and vice-versa with a BLEU score of 0.71. While complex computations as performed by or many commercial programs remain out of reach for our current methods, our experiments indicate that increased model and training sizes may make such learning possible in the future.
In summary, our contributions are as follows:
A novel, generic means to extract behavior models from programs.:
To the best of our knowledge, Modelizer is the first approach to automatically learn sequence-to-sequence models from arbitrary programs.
Using grammars to produce inputs for training.:
Modelizer uses grammars to produce inputs (and parse given inputs), and can thus learn from an unbounded number of training inputs and resulting outputs.
Abstract placeholders for tokenizers.:
Tokenizers that implement the suggested interface can automatically replace the real content with abstract placeholders, thus focusing on translating text structure rather than text content.
Predictive behavior models with high accuracy.:
The models produced by the Modelizer framework show high accuracy when predicting outputs for given inputs.
Reversible behavior models.:
The models produced by Modelizer can be reversed, producing inputs that are likely to result in given outputs.
The remainder of this paper is organized as follows: Section 2 describes in detail how Modelizer works, including input generation, dataset preprocessing, model learning, and model deployment. Section 3 covers the implementation aspects of Modelizer. In Section 4, we evaluate Modelizer on a variety of text-to-text translators, such as Markdown to HTML, query languages, and more. Section 5 discusses the limitations of Modelizer. After discussing related work (Section 6), Section 7 closes with a conclusion and future work. Modelizer and all experimental data are available as open source; see Section 7 for details.
2. Approach
Figure 2. How Modelizer works. Modelizer tests program with synthesized inputs and automatically learns reversible program behavior model from extracted input-output pairs.
[…]
How Modelizer works. Modelizer tests program with synthesized inputs and automatically learns reversible program behavior model from extracted input-output pairs.
Modelizer is a generic framework for learning program behavior. It uses synthesized data to automatically learn reversible generative models that can replicate the behavior of real-world programs with high accuracy. Figure 2 shows a general overview of the framework. Given the input specification for the selected PUT, Modelizer automatically generates the required number of inputs and passes them to the PUT for processing. Once the PUT produces the corresponding output, Modelizer extracts these input-output pairs and uses them to train a reversible neural machine translation model. The trained model can be further deployed on modern computer architectures. We have automated program behavior mocking using the following steps: Input Generation, Dataset Preprocessing, Model Learning, and Model Deployment. The following sections will cover each step in detail.
2.1. Input Generation
While developing our framework, we aimed to make it interoperable with different types of programs. Even though all programs are written in different programming languages, differ in implementation logic, and target different hardware architectures and operating systems, most programs are made to process input coming from users or other programs. Thus, we have decided to learn program behavior by training our behavior mocking models from the program inputs and processed outputs. Of course, such a strategy will enforce certain limitations:
  • The program behavior should be deterministic. At every execution, the program should produce the same output for the given input.
  • We will not be able to model the behavior of programs that process extremely long inputs or outputs at once. The computational complexity, memory, and time constraints scale quadratically with the input-output length.
  • Learning accurate inverse models that predict inputs for the given output is challenging when PUT computes multiple inputs to the same output. An intermediate representation for inputs could assist in such cases, which leads to context loss.
To learn an accurate model, we need sufficiently many diverse input-output pairs. Since public test datasets are typically biased towards commonly used features, and we do not know in advance how many inputs are required for our approach to learn the behavior of an unseen program, Modelizer assumes that inputs are automatically generated.
2.1.1. Input Generation with Grammars
To cover the entire input space (and hence increase our chances to also cover the behavior and output spaces), we use grammar-based input producers (Zeller et al., 2023). These encode input specifications as context-free grammars (CFG) (Chomsky, 1956)—a finite set of recursive string generation rules. For some input formats, such CFGs already exist and are publicly available. For example, the Grammar Zoo (Zaytsev, 2015) repository contains definitions of 1756 grammars of mainstream programming languages mined from standard documentation or community-created resources.
Figure 3. Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification
[…]
Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification
The first major component of the Modelizer framework is the Input Generator Module. This module functions as a data processing pipeline, which is sketched in Figure 3. It contains four steps: (1) Synthesis, (2) Post-Processing, (3) Hashing, and (4) Validation.
At the synthesis step, the input producer synthesizes new input samples according to the rules encoded in grammar. Then, the synthesized samples get refined to better align with real-world data. Such a refinement helps to map synthesized input features to real-world data features without increasing the size of grammars and, thus, simplifies the model learning process. The refined samples are then checked for uniqueness by computing a hash of the synthesized sample. If the hash value is not present in the Hashes Dataset, the sample is passed to the PUT for further processing. Otherwise, all steps are repeated till a unique sample is generated. The PUT produces the corresponding output, which is saved in a persistent storage along with the input. The collected input-output pairs are later used to train a program behavior model.
1<start> ::= <blocks>
[…]
3<element> ::= <heading> | <paragraph_block> | <blockquote> |
4 <horizontal_rules> | <ordered_list> | <unordered_list>
5<heading> ::= # <paragraph_text> | ## <paragraph_text> | ### <paragraph_text> |
6 #### <paragraph_text> | ##### <paragraph_text> | ###### <paragraph_text> |
[…]
9<paragraph_text> ::= <formatted_text> | [<formatted_text>](<url>) |
10 <code_block> | <paragraph_text> <paragraph_text>
[…]
14<code_block> ::= ‘<text>‘ | ‘‘‘<text>‘‘‘ | "<empty_space><text>"
15<quoted_text> ::= ’<text>’ | "<text>"
16<text> ::= TEXT
17<url> ::= URL
Listing 1: Markdown grammar used for input generation (excerpt)
Let us reuse our handwritten Markdown grammar (LABEL:lst:markdown-grammar) to demonstrate the synthetic input generation process. We will follow the steps that Modelizer takes to produce inputs. You can find the high-level implementation of the input generation pipeline written in Python in LABEL:lst:generate. Once the input producer is configured, it is ready to synthesize as many inputs as needed (Step 1). Below we give an example of synthesized input that was generated by the producer:
[…]
1def generate_data(grammar, min_terminals, max_terminals, total_elements, program_handler):
2 dataset = list()
3 seen_inputs = set()
4 processor = PlaceholderProcessor(grammar)
5 fuzzer = GrammarFuzzer(grammar, min_terminals, max_terminals)
6 while(len(dataset) < total_elements):
7 input_data = fuzzer.fuzz() # Step 1
8 input_data = processor.deduplicate(input_data) # Step 2
9 input_hash = sha384(input_data.encode("utf-8")).hexdigest() # Step 3.1
10 if input_hash not in seen_inputs:
11 seen_inputs.add(input_hash) # Step 3.2
12 output_data = exec(program_handler, input_data) # Step 4
13 dataset.append((input_data, output_data))
14 return dataset
Listing 2: Input generation in Modelizer (pseudocode)
The producer has synthesized one top-level Markdown element, which is a paragraph block. The paragraph block was expanded to a single paragraph text element which respectively contains a sequence of formatted text, hyperlink, and codeblock elements. One of the formatted text elements was further expanded to a bold text element. The synthesized input contains TEXT and URL placeholders instead of real character sequences. We have decided to define placeholder tokens for replacing the real-world data instead of encoding the whole language vocabulary into our grammar and synthesizing natural language sentences. We will elaborate more about placeholders in the follow-up section.
2.1.2. Placeholders
Synthesizing the data that is equivalent to real-world counterparts is a challenging task. Different programs can process the same data in various ways. This means that the grammar must be specified or refined according to program expectations. Depending on the complexity of input specification, the number of required input samples for efficient behavior learning can also become extremely large. We have figured out that for certain data types and programs, some fragments that are not affected or used by underlying implementation logic can be abstracted. For example, many text format converters only process the document structure, leaving the document content untouched. To benefit from such effects, we suggested and implemented two strategies in our input generation pipeline:
  • We suggest generalizing certain low-level non-terminal expansion rules by replacing a fraction of non-terminal to terminal expansions with generic placeholders.
  • We have added an optional post-processing step that augments placeholders with unique identifiers.
Both strategies will help to capture large portions of possible input space with few samples and correctly map real-world inputs to generic representations and vice-versa during the inference phase. Instead of learning the behavior from real-world examples, we learn how the PUT processes placeholders.
Let us refer back to the Markdown synthesis example and follow the input generation phases. The Placeholder Processor at Step 2 will replace these placeholders with equivalents that additionally carry unique identifiers. The output of the Placeholder Processor is shown below:
[…]
We assume that a TEXT placeholder can represent any character sequence in a real Markdown document. Similarly, a URL placeholder can substitute an arbitrary definition. By introducing such an abstraction layer, we can cover the whole space of structures of real-world documents with a relatively small number of synthesized samples.
While the input producer’s output is already well-formed and will pass the input parsing guards, it can be challenging to align real-world Markdown documents with the synthesized structure because it is impossible to distinguish the generated placeholders from each other. Therefore, we have included a post-processing phase in the input generation pipeline, for which we have implemented a generic Placeholder Processor. It must be initialized with the list of Placeholders that are encoded into selected grammar, like [TEXT, URL], and can automatically assign a unique identifier to every placeholder that it finds in the given sequence. Of course, it is possible to inject identifiers during the input generation. However, grammar-based input generators do not synthesize integers incrementally. They cannot guarantee that the generated identifier will be used in a synthesized sequence of characters only once. It is much faster and easier to generate a simple candidate sequence and then refine it using custom string-replacement or constraint-solving rules. The presence of unique identifiers improves the accuracy of both input-to-output and output-to-input translations, specifically in the cases when elements must be reordered during translation. During the tokenization process, which we will cover in the following sections, the respective data portions are assigned with a matching placeholder type that already includes a uniquely generated identifier. Optionally, the assignment of unique identifiers can be disabled.
Since grammar-based producers do not maintain a list of generated samples, we have implemented a sample validation procedure which first computes an HMAC-SHA-384 (Hansen and 3rd, 2011) value (Step 3.1) for the synthesized sample and then checks whether the computed hash value is stored in the Hashes Dataset (Step 3.2). If a computed hash value is absent in the dataset, the dataset gets updated with a new record (Step 4) and the synthesized input is passed to the PUT to compute output. Otherwise, the synthesized sample is dropped, and all input generation steps are repeated till the unique sample is synthesized. Corresponding input-output pairs are saved in a persistent storage for later processing. The collected input-output pairs are saved in a persistent storage for later processing.
2.2. Dataset Pre-processing
Machine translation is the task of automatically translating a text from one language to another. Neural machine translation models are the latest iteration of machine translation approaches that rely on deep neural networks to learn dependencies and patterns between sentences in the source and target languages. The ability to learn these dependencies from matching input-output pairs inspired us to reason about the program behavior modeling as a sequence-to-sequence translation task. In principle, the current generation of neural machine translation architectures that we use in our implementation can operate with the whole input at once. The computational complexity of these systems scales quadratically concerning the length of the input, and practical implementations do not follow this principle. Instead, text-to-text translations are performed by partitioning the whole input into sequences of small elements called tokens. Then, neural text translation models learn how to generate sequences of tokens in a target language by mapping the cross-token dependencies from a source language to a target language. Similarly, the synthesized input and computed output must be converted into batched tensors for the Learner to be able to process them. While text partitioning is a well-studied problem in the natural language processing domain (Manning et al., 2008), and there are already tools that can perform this task (Honnibal and Montani, 2017), we find the tokenization of program input and output to be a challenging task.
2.2.1. Pre-trained and Unsupervised Tokenizers
Instead of implementing tokenizers from scratch, one can potentially rely on existing generic tokenizers. For example, the HuggingFace repository (Wolf et al., 2020) provides access to many community-developed natural language tokenizers through the HuggingFace API. Such tokenizers cannot properly split the program data for several reasons. First, they ignore the data-specific or computation-specific features. Both factors impact the tokenization granularity, which is covered in the following paragraphs. Also, some tokenizers, like the pre-trained Llama3 tokenizer, modify the input by prepending or appending additional characters to the produced tokens. This can break the syntactical correctness when it comes to program data and can impact the quality of learned models.
While any input and output can be encoded as a sequence of characters, semantically it does not necessarily need to be textual data. Programs can accept or produce structural data. Such a structure contains knowledge that is relevant for behavior learning, but many existing tokenizers cannot distinguish the data structure from the content. Since, for certain computation tasks, like converting Markdown-to-HTML or SQL-to-KQL, the program is going to process the structure of the input while leaving content unchanged, a tokenizer needs to be able to map the fragments of the input to an abstract representation while processing. It is possible to augment the tokenization process of pre-trained tokenizers with additional pre and post-processing steps, where inputs can be transformed into semantically similar representations with placeholders replacing the fragments containing content. To build such a pipeline, the parser for the given data type has to exist and align with the input parsing rules of PUT.
Still, the program behavior learning approach is tokenizer-agnostic, and an unsupervised tokenizer like Google SentencePiece (Kudo and Richardson, 2018) is sufficient to quickly start the behavior learning process. However, the implementation of a data-specific tokenizer will help to further improve the quality of the learned model.
2.2.2. Masked Tokenization
While our approach works with generic tokenizers, we obtain better results with custom tokenizers adapted to the languages at hand. Our HTMLTokenizer adapted to HTML, for instance, takes an HTML string such as
<h1>Foobar</h1><p>Foobar is a <em>Python</em> library</p>
The result of the tokenization looks like following:
[ <h1>, TEXT_1, </h1>, <p>, TEXT_2, <em>, TEXT_3, </em>, TEXT_4, </p> ] { TEXT_1: Foobar, TEXT_2: Foobar is a, TEXT_3: Python, TEXT_4: library }
If we include the placeholder mapping as input to the reconstruction routine for the selected example, it will output the original HTML string. Otherwise, we will get the following output:
<h1>TEXT_1</h1><p>TEXT_2<em>TEXT_3</em>TEXT_4</p>
The reconstruction routine is inverse to tokenization, and for many data formats, it can be performed as a simple string concatenation. However, Modelizer provides several implementations for mapped string reconstruction. We support both early and late token instantiations, meaning that abstract tokens will be replaced with real-world values before or after the token concatenation. Furthermore, we support different mapping policies; see Section 3.2 for details of our implementation, including the pseudocode for HTML tokenization.
Full control over the tokenization process allows us to specify the tokenization granularity, impacting the prediction accuracy and model learning time. The vocabulary size grows corresponding to the number of unique tokens in the training set, which impacts the overall number of model parameters. It is useful to combine tokens that sequentially appear in the same context. With this measure, we reduce
  1. (1)
    the chances of breaking the syntactical correctness of the generated sequences;
  2. (2)
    the time required for learning the model; and
  3. (3)
    the time required for sequence generation.
How does one implement a domain-specific tokenizer? The simplest way is by to extend the parsing functionality, by deriving a parser from the grammar. In this case, we can reuse the grammar that was used for synthesizing inputs. Still, we cannot infer the output parser this way yet. Once we can automatically and correctly extract output grammars from PUTs, we will be able to infer domain-specific parsers and tokenizers for them.
2.3. Model Learning
Our Learner is based on a sequence-to-sequence neural machine translation model called Transformer (Vaswani et al., 2017). It learns a program-specific behavior mocking model from selected input-output pairs.
We present the high-level representation of the learning pipeline in LABEL:lst:learn and cover the steps taken in detail in the following paragraphs. Model Learner accepts the input-output data formats as input and loads the corresponding synthesized training sample pairs from the dataset. The loaded samples get automatically split into training, validation, and test sets.
1def learn_model(dataset, input_type, output_type, train_fraction, num_epochs):
2 dataloader = DataLoader(dataset, input_type, output_type)
3 train_split, validation_split, test_split = dataloader.random_split(train_fraction)
4 src_vocab, trg_vocab = dataloader.get_vocabularies()
5 param_searcher = HyperparameterSearcher(src_vocab, trg_vocab,
6 model_trials=250, lr_trials=100)
7 model_params = param_searcher.search(test_split)
8 model = Learner(model_params, src_vocab, trg_vocab)
9 for _ in range(num_epochs):
10 model.train(train_split)
11 model.validate(validation_split)
12 return model
Listing 3: The model learning pipeline in Modelizer (pseudocode)
Modelizer automates the hyperparameter tuning process. We have implemented a Hyperparameter Optimizer that tries to automatically find the best combination of configurable parameter values using a small disjoint set of test samples. Such a test set consists of synthesized 10,000 samples. The optimizer is built upon the Optuna (Akiba et al., 2019) framework, which is a hyperparameter optimization framework that can be used for any machine learning model. Parameter search is performed in two phases. The first phase is Model Parameter Search, which looks for the best configuration of the neural network for the fixed number of iterations (Vaswani et al., 2017). Using a fixed learning rate, the optimizer computes the value of the cross entropy loss metric for the different combinations of the Linear Layer size, Embedding size, the number of Attention Heads, and the number of Encoder and Decoder layers. The combination that has the lowest value of cross-entropy criterion is then used in the Learning Rate Search phase, where the starting learning rate (LR) value, as well as weight decay and learning rate adjustment policy, are determined. For our experiments, we have configured the Hyperparameter Optimizer to execute the Model Parameter Search phase for 250 trials and the Learning Rate Search phase for 100 trials.
Then the Learner gets initialized using the discovered model configuration and the model learning process gets executed for the specified number of iterations, called epochs. We use the cross-entropy loss function (Shannon, 1948), which is a generalization of the Logarithmic Loss function for multi-class classification problems, as our training evaluation criterion. It measures the dissimilarity of probability distributions between the expected and predicted labels. Also, we use the Adaptive Moment Estimation with Decoupled Weight Decay Regularization optimization algorithm (AdamW) (Loshchilov and Hutter, 2019) to automatically adjust the model’s parameters to minimize the training loss value. The validation split is used to evaluate the model’s performance on unseen data to prevent overfitting.
With our “lightweight” models, which, for example, require less than 6.3 million parameters to learn the Markdown to HTML conversion behavior of the Pandoc markup converter (MacFarlane, 2006), we achieve high accuracy in mocking program-specific behavior in a short amount of time on a commodity hardware. For the rest of the subjects that we used in the evaluation, the forward models contained 3-to-7.7 million parameters and inverse models required 2.8 to 10.3 million parameters to learn the hypothetical inverse behavior. The training time varies from several minutes to a couple of hours, depending on the subject.
2.4. Model Deployment
Figure 4. Model Deployment. An example scenario of the Learned Behavior Model predicting Program Input given the Program Output. The model prediction gets validated for correctness in the optional validation phase.
[…]
Model Deployment. An example scenario of the Learned Behavior Model predicting Program Input given the Program Output. The model prediction gets validated for correctness in the optional validation phase.
The model Deployment phase, which is summarized in Figure 4, includes the Prediction and optional Validation phases. Depending on the source data type, which can be either program input or output, the respective model, source tokenizer, and target tokenizer get initialized. From this point, model inference operates similarly to the model training phase. The source data gets tokenized (1) and passed to the Program Model (2), which predicts a target token sequence. Then, the predicted sequence gets lifted from token representation to real-world format by the target reconstructor component (3). The translation routine supports an optional max output sequence length parameter. If the framework user does not specify this argument, our default implementation dynamically estimates the maximum output sequence length by applying a 1.25 multiplier to the length of the currently processed input sequence.
Lastly, the predicted data can be passed to an optional Validation phase (4), which requires access to the respective PUT. It operates according to the model’s operation mode. For the behavior-mocking mode (predicting program output from a given program input), predicted output gets compared with computed output, which additionally gets tokenized and is marked as a reference sequence. Using both token sequences, we compute the Levenshtein distance (Levenshtein, 1965) and the BLEU (Papineni et al., 2002) score. Levenshtein distance, also named the edit distance in literature, is a metric for measuring the difference between two sequences, computed as a minimal number of insertions, deletions, and substitutions required to change the predicted sequence into a reference sequence. More on the BLEU score and other sequence similarity metrics will be covered in the next section.
In the Inverse-mocking mode (finding the program input used to compute the given program output), the computed output collected by passing the predicted input to the PUT gets compared with the initial program output. Since the actual program input (ground truth) is unknown, currently we can compute sequence similarity score metrics in cases when the program accepts the predicted input and does not fail while processing it. A failure of course will indicate the incorrectness of the predicted input. It is also challenging to validate the correctness of the prediction in the inverse-mocking mode if the PUT computes multiple valid inputs to the same output.
1def deploy(model, grammar, test_case, program_handler=None):
2 # Test the Model
3 in_tokenizer = Tokenizer(grammar)
4 out_tokenizer = CustomTokenizer()
5 test_tokens, test_mapping = in_tokenizer.mapped_feed(test_case) # Step 1
6 prediction = model.translate(test_tokens) # Step 2
7 reconstructed = out_tokenizer.mapped_reconstruct(prediction, test_mapping) # Step 3
8
9 # Evaluate the Translation
10 if program_handler is not None: # Step 4
11 reference = exec(program_handler, test_case)
12 if reconstructed == reference:
13 print("Model correctly predicted", reconstructed)
14 else:
15 reference_tokens, _ = out_tokenizer.mapped_feed(reference)
16 bleu_score = compute_bleu(reference_tokens, prediction)
17 edit_distance = compute_distance(reference, reconstructed)
18 print("Model failed to correctly predict the output")
[…]
20 print("BLEU Score:", bleu_score, "Edit Distance:", edit_distance")
21 else:
22 print("Model predicted", reconstructed")
Listing 4: Model deployment with the help of Modelizer (pseudocode)
Such programs have to be modeled by introducing an additional abstraction layer that defines the set of semantically equivalent inputs using formal rules.
We want to demonstrate the results of the model deployment process depicted in LABEL:lst:deploy using the example from Table 1. We used the HTML to Markdown conversion model learned from the Pandoc program, which was trained on 1,000,000 synthesized document pairs. The input HTML string is passed to the model, which predicts the corresponding Markdown string. At first, the input HTML string is tokenized and the content gets replaced with placeholders (Step 1). The model uses the produced input tokens to predict a sequence of Markdown tokens (Step 2). During the reconstruction phase (Step 3) the placeholder tokens in the predicted sequence get replaced with the real-world data. Since the Pandoc program was locally installed on the same machine, which was used to run the HTML to Markdown conversion model, we can validate the correctness of the predicted output by comparing it with the computed output. We measured the BLEU score at 1.0 and the Levenshtein distance at 0.0, which indicates that the predicted output is identical to the computed output.
Table 1. A conversion example from HTML to Markdown.
Input HTML Reconstructed Markdown
<p>Note that <a href="/docs/
transformers/v4.34.0/en/
model_doc/tapas#transformers.
TapasTokenizer">TapasTokenizer</a>
expects the data of the table
to be <strong>text-only</strong>.
You can use <code>.astype(str)
</code> on a dataframe to turn it
into text-only data. Of course,
this only shows how to encode a
single training example. It is
advised to create a dataloader
to iterate over batches.</p>
Note that [TapasTokenizer]
(/docs/transformers/v4.34.0/en/
model_doc/tapas#transformers.
TapasTokenizer) expects the data
of the table to be **text-only**.
You can use ‘.astype(str)‘ on a
dataframe to turn it into
text-only data. Of course, this
only shows how to encode a single
training example. It is advised
to create a dataloader to
iterate over batches.
Input Tokens Predicted Tokens
<p> TEXT_1 <a  href="URL_1" >
TEXT_2 </a> TEXT_3 <strong>
TEXT_4 </strong> TEXT_5 <code>
TEXT_6 </code> TEXT_7 </p>
TEXT_1 [ TEXT_2 ]( URL_1 ) TEXT_3 **
[…]
The conversion was performed by a model of the Pandoc program trained on 1,000,000 synthesized sequence pairs. BLEU score = 1.0, Levenshtein distance = 0.0
3. Implementation
Modelizer implements grammar-based input generators using the grammar fuzzers from FuzzingBook(Zeller et al., 2023). While the Fuzzing Book GrammarFuzzer randomly chooses the expansion rule from the set of available rules, its ProbabilisticGrammarFuzzer allows the developer to control the input generation by assigning probabilities to certain expansion rules. In this case, the non-attributed expansion rules will be assigned by equally distributed remaining probabilities.
Such a technique allows us the following:
  • Automatically generating a dataset of syntactically valid inputs.
  • Customizing the feature selection procedure in the sample generation process.
  • Ensuring the representation of specific and corner samples in the data corpus.
3.1. System Requirements
We implemented the model learning functionality of Modelizer in Python using the PyTorch 2.0 Machine Learning Framework (Paszke et al., 2019). Our framework currently supports the following hardware backends:
  • NVIDIA GPU that supports CUDA 11.7 or higher (Fastest)
  • Apple ARM-based system-on-a-chip with at least 16GB of system memory
  • System with a x64 processor and 32 GB of system memory (Slowest)
While model learning is a highly parallelizable task, and it can benefit from the availability of dedicated hardware accelerators, the rest of the tasks can also be scaled according to the number of available CPU cores. For example, to reduce the total runtime, our generators support the execution of multiple grammar-based fuzzers with the same configuration in parallel. The amount of Fuzzers that got initialized depends on the number of available CPU cores. Modelizer is responsible for distributing the total workload among all fuzzers by assigning every fuzzer with its own budget. Then each fuzzer produces samples in batches until it runs out of the budget. Modelizer systematically collects the generated samples and checks for their uniqueness. If there was no previously unseen sample generated within a single batch or a configurable parameter of failing attempts is reached, the value of the minimum and maximum non-terminals gets automatically incremented. To simplify the usability and applicability of our framework, we include a base class generator that can be extended to support new subjects.
3.2. Tokenization with Placeholders
Even though our models are trained on synthesized data, they must still be capable of operating with real-world data. However, there can be a difference in the representation between real-world and synthesized data due to the latter optionally including placeholders instead of real-world values, which impacts the implementation of our tokenizers. As a part of our contribution, we supplied the Modelizer with the base implementation and abstract interfaces for the mapped tokenization and mapped reconstruction routines. For every new data type, the framework user can wrap the functionality of unsupervised tokenizers with our mapped tokenization algorithm or implement a format-specific tokenizer in the subclass.
3.2.1. Abstract Tokenizers
By assuming that for both input and output formats, a suitable parser can be obtained, we implemented AbstractTokenizer—a generic mapped tokenization-reconstruction algorithm that can benefit from format structure knowledge and replications of input content fragments in the output. While it is not a fully functional tokenizer, it provides a high-level generic and extendible implementation of mapped tokenization-reconstruction routines, which assist in mapping real-world content to abstract placeholders and vice-versa. LABEL:lst:tokenize-base contains the fragment of the mapped tokenization routine.
1class AbstractTokenizer:
2 def mapped_tokenize(self, data: str):
3 self.clear_mappings()
4 self.mask_mapping = True
5 self.tokenize(data)
6 self.mask_mapping = False
7 return self.buffer, self.mappings
8
9 def mask_token(self, token: str, pldr: str) -> str:
10 tag = pldr + "_" + str(len(self.mappings[pldr]) + 1)
11 if self.mapping_policy == MappingPolicy.SIMPLIFIED:
12 self.mappings[pldr][tag] = token
13 tag = pldr
14 elif self.mapping_policy == MappingPolicy.OPTIMIZING:
15 if token in self.inverse_mapping:
16 tag = self.inverse_mapping[token]
17 else:
18 self.mappings[pldr][tag] = token
19 self.inverse_mapping[token] = tag
20 elif self.mapping_policy == MappingPolicy.EXHAUSTIVE:
21 self.mappings[pldr][tag] = token
22 self.inverse_mapping[token] = tag
23 return tag
24
25 def set_mapping_policy(self, policy: MappingPolicy):
26 self.mapping_policy = policy
27
28 def tokenize(self, data: str):
29 raise NotImplementedError
Listing 5: Tokenization in Modelizer (pseudocode)
To be able to correctly map real-world content to abstract placeholders (we use the term mask tokens, which you can see in the pseudocode) whether a parser-tokenizer processes content fragment of data the provided mask_token method must be called. Thus, the framework user must provide the format-specific implementation of the tokenize routine in a subclass. Then algorithm can correctly replace the content with an abstract placeholder during tokenization and successfully recover the original sequence of characters during the reconstruction. During initialization, the tokenizer, can learn the placeholders from the grammar definition automatically or operate with the user-defined collection of placeholders. The proposed algorithm allows reusing the same tokenizer for tokenizing both synthesized and real-world data.
3.2.2. A HTML Tokenizer
Below we demonstrate a partial implementation of the HTMLTokenizer class, which demonstrates the procedure of masking HTML content by extending the functionality of the AbstractTokenizer class.00footnotetext: The full implementation of HTMLTokenizer, which decomposes data using Python built-in HTMLParser, can be found in our GitHub repository (see Section 7 for details). This repository also provides additional examples of tokenizer implementations that rely on PUT’s functionality to parse inputs or implement a custom parser using regular expressions.
1from html.parser import HTMLParser
2
3class HTMLTokenizer(AbstractTokenizer, HTMLParser):
4 def tokenize(self, data: str):
5 self.buffer.clear()
6 HTMLParser.feed(self, data)
7
8 def handle_data(self, data: str):
9 if self.mask_mapping == True: # True only if called via mapped_tokenize(data)
10 token = self.mask_token(data, self.placeholders[1])
11 self.buffer.append(token)
12 else:
13 tokens = self.split_and_align(data)
14 self.buffer.extend(tokens)
Listing 6: Tokenization in Modelizer (pseudocode)
Once the mapped_tokenize method is called, the tokenizer clears all internal buffers and starts processing the given string with the selected Parser. In this implementation, the HTMLParser requires the end-user to implement handler methods to process HTML tags together and document content. Here we present the simplified implementation of the document content handler handle_data. Depending on the operation mode of the tokenizer, which is determined by the mask_mapping flag, the tokenizer either replaces the document content with a matching placeholder or performs a sequence of format-specific string splitting operations, which are abstracted in this example.
3.2.3. Token Masking Strategies
The token masking algorithm behaves differently according to one of the selected token mapping policy strategies:
  • The SIMPLIFIED policy reduces the vocabulary size of the model by removing identifiers from placeholder tokens.
  • The OPTIMIZING policy, which is the default policy in our implementation, uses a minimal number of placeholder tokens to encode data content.
  • The EXHAUSTIVE policy assigns every content fragment with a unique placeholder token.
We use the extended version of the HTML document from Figure 1 to illustrate the effect of using different token masking strategies in tokenization:
<h1>Foobar</h1><p>Foobar is a <em>Python</em> library. Find more on <a href="127.0.0.1"><i>Foobar</i>.com</a></p>
When the tokenizer is running in SIMPLIFIED token masking mode, then it removes every unique identifier that had been assigned to a placeholder. It significantly reduces the model’s vocabulary size, but increases the chances of producing errors during the reconstruction phase. The tokenized sequence looks like the following:
[ <h1>, TEXT, </h1>, <p>, TEXT, <em>, TEXT, </em>, TEXT,
<a, href="URL">, <i>, TEXT, </i>, TEXT, </a>, </p> ]
The tokenizer tries to use a minimal number of placeholders to encode the real-world data when it is configured to use OPTIMIZING token masking strategy. The repetitive content is encoded by the same placeholder. This is the default operation mode and evaluation experiments were arranged only using this policy. The tokenized sequence looks like the following:
[ <h1>, TEXT_1, </h1>, <p>, TEXT_2, <em>, TEXT_3, </em>, TEXT_4,
<a, href="URL_1">, <i>, TEXT_1, </i>, TEXT_5, </a>, </p> ]
The EXHAUSTIVE token masking strategy increases the model’s vocabulary size and negatively affects the model learning time. The tokenizer maps every content piece in the processed input to a unique identifier. However, this policy is useful for programs that reorder elements during processing. The tokenized sequence looks like the following:
[ <h1>, TEXT_1, </h1>, <p>, TEXT_2, <em>, TEXT_3, </em>, TEXT_4,
<a, href="URL_1">, <i>, TEXT_5, </i>, TEXT_6, </a>, </p> ]
4. Evaluation
We evaluated Modelizer by learning the execution behavior of real-world programs. For every PUT, we collected two datasets of input-output pairs by executing the program with synthesized and real-world inputs respectively. For every evaluation subject, Modelizer additionally learned an inverse model  $M^{-1}$ that corresponds to a hypothetical program  $\textnormal{PUT}^{-1}$, which is expected to implement an inverse of the behavior of the original program  PUT.
4.1. Research Questions
In the evaluation, we address the following research questions:
  1. RQ1.
    How accurate are the Modelizer models that mock the behavior of the  PUT? This question is at the heart of Modelizer: How well do the learned models perform in predicting outputs from inputs?
  2. RQ2.
    How accurate are the Modelizer models that are trained on synthesized data in processing real-world data? This question evaluates the quality of the learned models when tested with real-world data.
  3. RQ3.
    How accurate are the inverse models that mock the behavior of  $\textnormal{PUT}^{-1}$? With RQ3, we evaluate an important application of Modelizer, namely its ability to predict inputs from outputs.
  4. RQ4.
    How fast Modelizer can synthesize a training dataset? This question evaluates the efficiency of the input generation pipeline implemented in Modelizer.
  5. RQ5.
    How fast Modelizer can learn program models? This question evaluates the efficiency of Modelizer core functionality, which is model learning.
  6. RQ6.
    How well do the synthesized samples represent the real-world data? With RQ6, we investigate the quality of the synthesized training set.
  7. RQ7.
    How much synthetic data is required for mocking behavior at which accuracy? With RQ7, we investigate the impact of the size of the synthesized training set on the accuracy of learned models.
  8. RQ8.
    How should synthetic data be generated to achieve the highest accuracy of the learned models? With RQ8, we investigate the impact of the synthetic data generation strategies on the accuracy of learned models.
  9. RQ9.
    How does Modelizer learn the hyperparameters for each data pair combination? This question assesses the impact of hyperparameters (Section 2.3) on the performance of Modelizer.
  10. RQ10.
    Should Modelizer re-use learned hyperparameters for inverse modeling? With RQ10, we check the reusability of found hyperparameters (Section 2.3) with forward models for training inverse models.
  11. RQ11.
    What is the effect of placeholders on behavior learning? This question assesses the impact of placeholders (Section 2.1.2) on the performance of Modelizer.
  12. RQ12.
    What is the effect of tokenizers on learning performance? With RQ12, we investigate the impact of tokenizers (Section 2.2) on the performance of Modelizer.
4.2. Evaluation Subjects
We conducted the evaluation using the four programs listed in Table 2:
Table 2. The list of evaluated programs and their respective input-output types.
Program Input Type Output Type Version
Pandoc Markdown HTML 2.19.2
msticpy SQL KQL 2.9.0
latexify-py Python Math Expression 0.2.0
py-asciimath MathML 0.3.0
py-asciimath AsciiMath 0.3.0
py-asciimath AsciiMath MathML 0.3.0
Pandoc:
(MacFarlane, 2006) is an open-source library and command-line tool for converting markups from one format to another. While the library natively supports the translation in both directions(Markdown to HTML and HTML to Markdown), we have tested it only with Markdown to HTML translation functionality. We manually extracted a Markdown Grammar from the CommonMark Spec Version 0.29 (MacFarlane, 2019) specification. We used Pandoc’s functionality to convert Markdown to HTML for our case study, first, to validate the feasibility of neural behavior modeling of black-box systems, second, to check the applicability of the same technique for inverse behavior modeling, when a model has to predict the program input given the output. The promising results of our first experiments encouraged us to test the approach with other subjects.
Microsoft Threat Intelligence Python Security Tools:
(msticpy) (Hellen, 2019) is a library for InfoSec investigations, data analysis, and hunting in Jupyter Notebooks authoring for Azure Sentinel, which is a cloud-native security information and event management solution offered by Microsoft for intelligent security analytics and threat intelligence. Modelizer was evaluated on the SQL to KQL Conversion feature of MSTIC, which accepts SELECT SQL queries constructed using a subset of ANSI SQL-92 (ISO/IEC 9075:1992, 1992) as input and produces Kusto Query Language (KQL) queries. KQL is a proprietary query language developed by Microsoft, which is used for querying and analyzing data in Microsoft Azure services. The supported subset of SQL-92 specification was encoded as a CFG grammar by us.
latexify-py:
(Oda, 2020) is a Python library that converts Python math expressions to code. The grammar for synthesizing Python math expressions encodes the math module (Foundation, 2021) functionality from the Python 3.10 standard library.
py-asciimath:
(Belotti, 2020) is a Python library for converting mathematical expressions from AsciiMath representation to Mathematical Markup Language (MathML) and representation. The grammar for synthesizing MathML was manually extracted from the MathML 2.0 Specification (Ausbrooks et al., 2003). The training pairs were collected by converting the synthesized MathML formulas first into format, and then the collected formulas were translated once more into AsciiMath format. This way, we managed to collect three different sets of input-output pairs and test the framework’s capabilities in performing the cross-protocol translations in directions that the given PUT does not currently support.
4.3. Training Setup
For every subject, we generated training sets of different sizes (5k, 10k, 50k, 100k, 250k, 500k, 1M). The input generation starts with the minimal number of non-terminal expansions (MinNonTerminals) set to 10 and maximal (MaxNonTerminals) set to 20. Once configured, the generator has 100,000 attempts to generate unique samples using a fixed configuration. If the generator runs out of attempts the values of MinNonTerminals and MaxNonTerminals parameters are increased by 10 and the generator is restarted. The generation process will be running till the generator produces the required number of unique samples. Also, we configured the data generators to produce a separate set of input samples that had been synthesized with a continuously increasing complexity of the generated samples. It was achieved by implementing the sliding window modification of MinNonTerminals and MaxNonTerminals values. In this case, the generator uses at least four different non-terminal expansion configurations to generate the required number of samples.
Pandoc, msticpy, and py-asciimath subjects were evaluated using real-world data. We evaluated the SQL to KQL translation models on 1,000 samples that were collected from the “sql-create-context” open-access dataset (brianm, 2023). This dataset contains natural language to SQL SELECT query pairs. During testing the msticpy subject, we ensured that it produced a non-empty output for every extracted SQL query extracted from the dataset. For the py-asciimath subject, we also extracted 1000 real-world MathML formulas from the NTCIR-12 dataset (Zanibbi, 2016). For testing the models of the Pandoc program we extracted 12,783 top-level elements (tree-like structures nesting other markdown definitions) from 100 markdown files that were collected from the open-access Markdown dataset (Schmid, 2023) containing readme documentation files of projects hosted at the HuggingFace Hub (Face, 2024). We have not managed to find a dataset containing mathematical expressions written in Python using the math module from the Python Standard Library. Instead, we synthesized 1000 samples using our Python Expression grammar and replaced the placeholders with random values of corresponding data types. Additionally, we evaluated every model on 10,000 synthesized samples that were not included in the training set. The synthesized dataset was generated by fixing the minimal number of non-terminal expansions to 10 and the maximal number of non-terminal expansions to 20.
While the decoder of Modelizer’s translation routines supports a configurable beam size parameter, we collected all of our experiments with a fixed beam size = 1. All models were trained for 10 iterations(epochs) with a context-window size equal to 5000 tokens. In every training configuration, 80% of the synthesized data had been used for model training. After every training iteration, the training quality was validated using the remaining 20% of the training set. Additionally, the learning quality across models was evaluated with 10,000 previously unseen data pairs which were collected using synthesized inputs, that were generated by setting the maximal number of non-terminal expansion to 20. We trained all models using a single NVIDIA Geforce RTX 4090 GPU with 24 GB of Memory. During the model learning phase Modelizer was allocating from 650 MB to 22 GB of memory depending on the configuration of hyperparameters and the sizes of collected vocabularies. Once trained, models can be deployed on any modern computer architecture that is equipped with at least 8 GB of system memory and can execute a Python interpreter version 3.10 or newer. In contrast to real-world programs the learned models do not require any external dependencies or permissions for system-wide installation. The required Python interpreter and the set of Python modules can be pre-configured and shipped with the model instance. A single environment can be used to execute multiple instances of the model in parallel. Because the size of trained models is relatively small, we monitored a low memory footprint when they were running in the inference mode.
4.4. Evaluation Metrics
To evaluate the quality of the learned models, we have used the following evaluation metrics:
BLEU score:
(bilingual evaluation understudy) (Papineni et al., 2002) is a commonly used metric for the evaluation of machine translation systems, that calculates the geometric average precision of single, double, triple, and quadruple token overlap cases between reference and hypothesis token sequences multiplied by brevity penalty. The score output is always a number between 0 and 1, with 1 indicating the perfect translation case. The BLEU score was computed on the whole corpus of test cases using the implementation provided by the Python NLTK package (Team, 2007).
BLEU Error:
stands for Standard Error computed using BLEU score.
NIST score:
(Doddington, 2002) is a metric developed by the US National Institute of Standards and Technology for evaluating the quality of machine-translated texts, which is also based on the BLEU score and additionally calculates the frequency of n-grams and gives a higher score to n-grams that are considered to less likely occur. We calculated the metric using the implementation provided by the Python NLTK package (Team, 2007).
Word Error Rate:
(WER) (Morris et al., 2004) is a metric that calculates the accuracy of a model prediction by computing a ratio between the sum of token insertions, deletions, and substitutions required to transform the predicted token sequence into the reference token sequence and the total number of tokens in the reference token sequence. The lower the value the better the model performs with a score of 0 indicating that the model has not made any mistakes during the translation process.
World Information Lost:
(WIL) (Morris et al., 2004) is a metric that computes the percentage of total tokens that were incorrectly predicted by the model between sets of reference token sequences and sets of hypothesis token sequences. The lower the value the better the model performs with a score of 0 indicating that the model has not lost any information during the translation process.
Exact Match:
is a metric that reflects the percentage of evaluation samples that were translated correctly by the model. The higher the value, the better the model performs with a score of 100 indicating that the model has not made any mistakes during the translation process. This metric relies on the Levenshtein distance to compare the similarity between the predicted and reference sequences.
Close Match:
is a metric that reflects the percentage of evaluation samples that were translated by the model with a single error. We treat such cases as potentially recoverable at a post-processing stage. This metric relies on the Levenshtein distance to compare the similarity between the predicted and reference sequences.
4.5. Results
Let us now turn to the evaluation results.
Table 3. Evaluation results for models mocking Pandoc, msticpy, latexify-py, and py-asciimath behaviors.
Program Source Target
Model
Type
Size BLEU
BLEU
Error
NIST WER WIL
Exact
Match
Close
Match
1M 0.9986 0.0033 9.9102 0.04 0.05 98.69 99.39
100k 0.2923 0.3025 4.2689 54.2 72.13 51.72 56.03
Markdown HTML Forward 500k 0.7128 0.2337 7.5700 23.19 35.30 90.57 94.79
1M 0.9988 0.0045 10.9328 0.08 0.11 99.27 99.55
250k 0.2352 0.2319 4.0123 52.62 66.00 55.78 57.51
Pandoc HTML Markdown Inverse 1M 0.7036 0.1554 7.5202 23.74 34.48 93.27 93.58
500k 0.9997 0.0035 11.0359 0.03 0.03 99.58 99.69
100k 0.5040 0.0497 5.0374 21.48 38.07 1.60 1.90
SQL KQL Forward 100k 0.9791 0.0397 6.4760 1.51 2.09 95.40 96.90
250k 0.9206 0.0374 11.0730 6.41 7.48 27.51 32.64
50k 0.9566 0.0388 5.0318 2.92 4.51 89.00 93.90
msticpy KQL SQL Inverse 100k 0.9725 0.0379 5.1913 1.94 2.82 94.60 95.80
1M 0.9969 0.0543 12.9138 0.45 0.48 88.47 88.86
50k 0.5380 0.1796 6.2148 23.30 23.60 0.00 0.00
PyExpres. Forward 100k 0.9219 0.2806 9.8394 8.48 9.59 48.50 53.50
250k 0.9612 0.1039 15.9279 3.39 4.12 40.69 43.98
10k 0.8595 0.0908 10.7279 9.43 12.48 31.60 52.40
latexify-py PyExpres. Inverse 100k 0.8747 0.1240 11.0931 10.20 12.21 44.60 56.00
1M 0.9999 0.0008 11.6252 0.01 0.01 99.82 98.92
250k 0.3764 0.3522 4.7470 48.28 58.93 0.00 1.40
AsciiMath Forward 1M 0.7553 0.2994 8.4978 19.24 24.15 35.70 35.00
1M 0.9983 0.0030 12.1796 0.11 0.12 98.86 99.86
50k 0.3025 0.4990 4.4004 55.25 67.26 0.10 0.40
AsciiMath Inverse 1M 0.6899 0.3470 8.1353 24.57 32.19 30.40 42.50
Outliers
500k 0.9846 0.0086 11.9758 0.93 1.36 93.33 93.81
100k 0.0568 3.6482 1.6700 84.26 94.50 0.00 0.00
MathML Forward 5k 0.2674 1.8474 3.3514 58.75 71.58 3.00 1.40
500k 0.9782 0.0107 9.2103 0.88 1.71 61.09 93.44
500k 0.0000 0.5871 0.0171 97.16 99.69 0.00 0.20
MathML Inverse 10k 0.5336 0.5286 4.1001 39.32 50.64 1.00 7.40
1M 0.9780 0.0101 9.2108 0.92 1.74 60.81 94.01
500k 0.0000 3.6580 0.0134 97.40 99.72 0.00 0.00
AsciiMath MathML Hypothet. Forward 250k 0.3084 2.7773 1.1503 53.34 59.63 0.70 0.70
500k 0.9844 0.0112 11.4014 0.94 1.37 93.31 93.41
5k 0.1028 0.6621 1.9223 74.02 84.91 0.00 0.20
py-asciimath MathML AsciiMath Hypothet. Inverse 5k 0.2963 0.5779 3.6842 54.06 67.69 3.40 8.80
Yellow cells indicate an evaluation run with synthetic test data, blue cells indicate evaluation runs with the real-world data, and purple cells indicate the evaluation runs on the real-world data performed after fine-tuning the baseline model on a few real-world data pairs.
RQ1: Mocking Accuracy
We start with addressing RQ1: How well do Modelizer models predict outputs from inputs? To answer this question, we evaluated our models on sets of 10,000 synthesized test cases per subject that have no intersection with any of the training sets. The results of the evaluation are summarized in Table 3. The evaluation results show that the models trained on the synthetic data can achieve a high prediction accuracy for most subjects from the evaluation set. For most models and subjects, the BLEU score is above 0.98, the Word Information Lost score is below 0.02 and the Exact Match is above 93%.
Table 4. Evaluation results of Markdown to HTML conversions done with the learned models of Pandoc converter.
Dataset Trained on Synthetic Data Fine-Tuned on Real Data
Size Partition BLEU WIL
Exact
Match
Close
Match
BLEU WIL
Exact
Match
Close
Match
0.2576 76.16 51.40 55.93 0.6763 36.97 89.80 94.01
5k 0.2710 72.94 51.64 55.96 0.6813 42.80 89.35 93.61
0.2746 74.28 51.40 55.93 0.6934 40.00 88.53 92.73
10k 0.2723 75.36 44.43 52.11 0.6747 42.95 89.07 93.24
0.2851 75.21 51.71 56.01 0.6735 43.97 89.42 93.80
50k 0.2849 73.25 51.69 56.03 0.6874 40.49 89.72 93.84
0.2923 72.13 51.72 56.03 0.6966 36.79 88.66 92.88
100k 0.2704 73.26 51.61 55.89 0.6592 42.85 88.99 93.27
0.2898 75.39 51.72 56.03 0.6725 38.03 89.30 93.56
250k 0.2824 74.49 49.67 52.17 0.6753 37.61 89.03 93.23
0.2900 70.90 40.49 55.78 0.7128 35.30 90.57 94.79
500k 0.2950 68.56 51.71 56.02 0.6867 35.64 89.31 93.76
0.2535 72.90 24.38 28.68 0.6886 33.95 90.25 94.56
1000k 0.2940 70.49 51.61 55.91 0.7208 36.64 90.35 94.69
Models were trained on the various sizes of datasets containing synthesized Markdown to HTML structures. Blue cells indicate models trained on synthetic data generated with a fixed number of non-terminal expansions, and purple cells indicate application of the synthetic data with an increasing number of non-terminal expansions.
Table 5. Evaluation results of HTML to Markdown conversions done with the learned inverse models of Pandoc.
Dataset Trained on Synthetic Data Fine-Tuned on Real Data
Size Partition BLEU WIL
Exact
Match
Close
Match
BLEU WIL
Exact
Match
Close
Match
0.1988 76.01 55.08 56.83 0.6441 43.73 81.51 90.85
5k 0.1886 76.63 28.36 57.40 0.6550 42.45 91.14 91.97
0.2006 74.47 55.71 57.46 0.6441 44.38 90.39 91.26
10k 0.2316 72.49 28.39 57.47 0.6280 45.16 87.64 88.59
0.2115 70.59 53.72 55.25 0.6282 43.94 92.15 92.93
50k 0.2503 65.47 55.75 56.84 0.6170 46.18 88.61 88.97
0.2096 69.33 50.58 57.31 0.6119 45.50 88.60 88.95
100k 0.2112 70.16 28.62 56.03 0.6471 37.28 89.24 91.39
0.2218 65.89 55.78 57.31 0.6499 40.72 90.57 90.97
250k 0.2352 66.00 55.78 57.51 0.6778 39.30 92.20 92.52
0.2251 66.88 55.78 55.89 0.6793 36.75 91.51 92.72
500k 0.2145 68.21 19.78 47.65 0.6646 40.97 89.02 89.56
0.2326 67.26 54.97 56.31 0.6749 39.36 91.08 91.36
1000k 0.2408 64.17 55.75 57.31 0.7036 34.48 93.27 93.58
Models were trained on the various sizes of datasets containing synthesized Markdown to HTML structures. Blue cells indicate models trained on synthetic data generated with a fixed number of non-terminal expansions, and purple cells indicate application of the synthetic data with an increasing number of non-terminal expansions.
According to the prediction accuracy scores depicted in Table 3 and representative results from both Table 4 and Table 5, the medium size of training sets can already achieve a high prediction accuracy for most subjects from the evaluation set. Using the given grammar definitions, the data generator can produce the vast majority of data feature combinations within the 100,000–250,000 samples range. By further increasing the number of generated samples, while eliminating duplicates, the generator starts to produce a more unnatural combination of features, which tends to negatively affect the model accuracy, makes the learning more costly, and hits the prediction performance.
Modelizer models trained on synthetic data have a high accuracy in predicting outputs from inputs.
RQ2: Mocking Performance with Real Data
To answer this question, we evaluated our models on sets of 1,000 real-world test cases per subject, except Markdown and HTML, where, as already mentioned, we took 100 real-world readme documents and extracted 12,783 top-level elements from them. These test sets have no intersection with any of the train sets. While the models achieve great accuracy when tested with synthetic data, it is noticeably lower when tested with real-world data.
For the input generation, we did not learn the feature distribution from the particular set of test cases to avoid overfitting to the test data and check Modelizer’s capabilities in learning the program behavior while having no additional knowledge of the test subject and the data it processes. Consecutively, the feature distribution between the real-world and synthesized samples did not align well resulting in a significant performance hit. The availability of additional data pairs for the few-shot fine-tuning of the learned models can significantly boost the prediction accuracy and partially compensate for the feature distribution inconsistency at a small cost. After the fine-tuning, we recorded the edit distance value between expected and predicted output to be greater than 1 mostly for a small fraction of test samples. The inspection of these test cases has shown that they contain features that were not part of the grammar that we used to synthesize input/output pairs, thus samples with similar sets of features were not present in the training data. The only way to fix this issue is to populate the grammar definition with new expansion rules. In contrast to the majority of test data, a few test cases contained an unusual number of placeholder tokens of a particular type. More on this will be discussed in the next research question.
Modelizer requires that the training data (and hence, the synthesizing grammar) covers the features found in the testing data.
In this evaluation, all models that had been trained on MathML Formulas used as input or output (MathML to , to MathML, MathML to AsciiMath, AsciiMath to MathML) have shown the lowest prediction accuracy. We treated them as outliers and investigated the cause of failure. The reason for this is the complexity of the MathML language and the lack of the corresponding test-case features in the grammar we used for input generation. The MathML language supports several operators and operands, including the different encodings for the semantically same element. The grammar we used for the input generation covers only a fraction of these features. Due to this fact, the custom tokenizer also fails to parse these special cases. The better mapping of operators and operands during the tokenization will help to resolve this issue. While the MathML grammar was used for synthetic input generation and the py-asciimath library was utilized for sequential conversion of first MathML to and then to AsciiMath during the training dataset formation, the accuracy of AsciiMath to and to AsciiMath models were less impacted by the inconsistency of the vocabularies. The accuracy drop was only caused by the lack of the specific unseen features, while the encoding variations of the same operators got mapped to the same command by the py-asciimath library in the and AsciiMath representations, thus the model’s vocabulary had already contained them. However, as the Table 3 also depicts, the MathML models manage to learn the conversion behavior for the synthesized test cases. We believe that alterations to the grammar and the tokenization process will help to improve the accuracy of the learned models for MathML Formulas.
In the performance evaluation, MathML is an outlier, due to the restricted grammar used for input generation.
RQ3: Inverse Accuracy
We turn to RQ3 and evaluate the accuracy of Modelizer when predicting inputs from outputs. We again refer to the prediction accuracy scores depicted in Table 3 and representative results from both Table 4 and Table 5. In particular, we are interested in HTML to Markdown, KQL to SQL, to PyExpression, and AsciiMath to models. The numbers show that the inverse model achieves a similar Exact Match accuracy, as their counterpart forward conversion models. The BLEU Score values that we recorded for both forward and inverse models are high and have insignificant margins between them. While the models of the Pandoc and msticpy library achieve more than 90% accuracy in both directions, the latexify-py and py-asciimath can produce fully correct results in half of the cases. Between 20% to 35% of predictions require up to 4 edit operations, including token insertions and deletions, to correct the prediction. Currently, Modelizer can automatically detect these failures using the prediction validation mechanism discussed in Section 2.4. We plan to integrate the automatic prediction repair mechanism in our future work.
Learning inverse models is as performant as learning mock models.
RQ4: Dataset Generation Time
Figure 5. Speed of Modelizer training data generation. This graph represents the data generation time distribution for different subjects and training set sizes. The x-axis represents the input types, and the y-axis represents the time in minutes. Every training set size is represented by a different color. For every input type, the generator was producing samples with fixed and increasing complexity (partitioned).
[…]
Speed of Modelizer training data generation. This graph represents the data generation time distribution for different subjects and training set sizes. The x-axis represents the input types, and the y-axis represents the time in minutes. Every training set size is represented by a different color. For every input type, the generator was producing samples with fixed and increasing complexity (partitioned).
We measured the performance of our input generation pipeline in forming a tokenized training set of generated positive samples of data pairs. Our generators can be configured to ignore all PUT runs when PUT fails to process the synthesized input. In this evaluation, our generators we configured to ensure that the number of requested and produced data pairs was equal. Generation of MathML inputs was a special case, where the input generator was executing the PUT twice to first produce from MathML, and then AsciiMath from . The dataset generation time plotted on Figure 5 includes input generation with grammar fuzzer, check for the uniqueness of produced input, input processing with PUT, tokenization of both synthesized input and produced output, and flushing the formatted dataset to the persistent storage. As Figure 5 highlights, for the majority of subject and input generation strategy combination pairs, it required 2-to-4 minutes to synthesize 10,000 samples, 1.5 to 2 hours to produce 250k samples, and more than 6 to 8 hours to produce a dataset with 1,000,000 samples using an Ubuntu Linux workstation equipped with AMD Ryzen Threadripper 3960X 24-Core processor and 256 GB of RAM. The dataset generation time increases almost linearly according to the number of included records. We also see that input generation with an increasing number of non-terminal expansions takes from 0.3 to 2.5 times longer, and is on average 44% slower. Also, we want to mention that by increasing the complexity of synthesized inputs for SQL and MathML we started to get fewer input samples that PUT can successfully process. While this can be seen as a desired effect for stress testing the application, it significantly slows down the dataset generation process if the end goal is to form a dataset containing only positive samples.
1.5 to 2 hours are required to produce a mid-size dataset containing 250,000 samples using a multicore processor.
RQ5: Model Learning Time
Figure 6. Speed of Modelizer learning models from samples with a fixed and increasing complexity (Partitioned). All model learning was done using a single NVIDIA Geforce RTX 4090 GPU.
[…]
Speed of Modelizer learning models from samples with a fixed and increasing complexity (Partitioned). All model learning was done using a single NVIDIA Geforce RTX 4090 GPU.
According to timing records from the model learning log, it takes Modelizer less than half an hour to learn a model from 10,000 synthesized training pairs using NVIDIA RTX 4090 GPU. The similar models will be trained within 7 to 12 hours using a mid-size dataset with 250,000 training samples. Lastly, it takes Modelizer 22 to 49 hours to train a model using a dataset with 1,000,000 records. Figure 6 demonstrates the model learning time distribution across subjects from the evaluation set, training set sizes, and input generation strategies. While an early program mock can be produced already within 30 minutes using a GPU, a significantly more accurate model will be trained within 10 hours.
We also measured from 1% to 59% slowdown, with an average of 11%, while training models containing inputs generated with fixed or increasing complexity. The training time grows almost linearly with the increase of the training set size, given the automatically discovered hyperparameter configurations per subject from the evaluation set.
Training a Modelizer model can be done on a common PC in a few hours.
RQ6: Real-World Data
Our next research question is RQ6: How well do the synthesized samples represent the real-world data?
The quality of synthesized samples depends on the presence and correctness of the corresponding expansion rules that encode data-relevant features in the grammar. We evaluated the quality of the synthesized samples by checking for the inclusion of tokens collected from real-world samples into the vocabularies of learned models. We have collected the following statistics per data type used in our evaluation:
Markdown.:
One top-level element out of 12,783 contained a high number of URL placeholders. Unfortunately, our generators, given the selected non-terminal expansion configuration, have not produced sequences containing more than 51 URL declarations within 1 generation cycle. Thus, the corresponding placeholders were not included in the model’s vocabulary, resulting in encoding these tokens as “UNKNOWN” and the model’s inability to fully correctly predict the output.
HTML.:
Similar to Markdown, in one of the top-level HTML elements we have observed the absence of the same set of URL placeholders in the model’s vocabulary. Additionally, from 9 more top-level elements a few extra HTML tags, like <head> </head> <title> </title>, and meta attributes, like data-cites class, were not seen during the data generation. During the data generation, we filter out attributes that do not affect the semantics of the output data. But, we have not manipulated the test data. Thus the model prediction could be semantically correct but insignificantly differ from the reference sequence.
SQL.:
The test data contained 3 test cases that had been accessing more than 3 tables in a single query. The synthesized queries did not contain such complex samples, which resulted in the model’s inability to produce a fully matching prediction for these test cases.
KQL.:
2 out of 1,000 queries have accessed more than 3 tables in a single request and used 1 additional feature not seen during the data generation. The model cannot predict the output for these queries.
MathML Formulas.:
In 115 cases out of 1,000 formulas, the token sequence contained the unseen operands and operators. This problem originates from the ability of the custom tokenizer to parse and process these elements due to encoding and lack of the corresponding parser rules. The grammar we have defined for the synthesis of MathML Formulas does not cover all possible encodings for operators and contains only a fraction of the special operand space. Due to this fact, the custom tokenizer also fails to parse these special cases. While our tokenizer already abstracts the numeric values and string literals with the corresponding placeholders, we have not implemented the same abstraction mechanism for the special operands. The better encoding and mapping of operators will also help to represent such uncommon inputs without the necessity to store them in the model’s vocabulary. The last issue is the utilization of more than 18 identifiers within a single formula. To synthesize such patterns the data generation configuration has to be changed.
and AsciiMath Formulas.:
Similar to the MathML case we have detected the both larger than synthesized number of enumerated placeholders and the presence of the unseen operands and operators in the test data in 70 and 73 token sequences respectively.
Python and Expressions.:
No unseen tokens were detected.
MathML to AsciiMath behavior mocking and vice-versa, and MathML to  behavior mocking and vice-versa are representative demonstrations of a correlation between the quality of the grammar used for input generation and the accuracy of learned models. The more features are covered by the grammar, the more program behavior gets tested and leaked, and the better model is capable of accurately mocking the program behavior. A fully trained model continuously failing to predict the output acts as a marker of the missing features in the training data, thus in the input grammar. The failure can be detected using a mechanism described in Section 2.4.
At this point, two scenarios are possible:
  • First, the grammar that has been used in the data generation process was an estimate of the actual protocol specification, and it has to be refined according to newly discovered features.
  • Second, the grammar follows the protocol specification, but the program implementation does not. Then the learned model can be used as an oracle to detect the program’s behavior deviation from the protocol specification.
In the case of predicting conversions of MathML Formulas to and AsciiMath representations, we demonstrated the necessity of the population of the grammar with new features and improvements in the tokenization process. Our study of the Pandoc markup converter with Modelizer detected a deviation from the CommonMark Spec Version 0.29 (MacFarlane, 2019). Once we extended our grammar declaration with Extended Markup Syntax, the model accuracy significantly improved. Thus, we think that our models can detect behavior deviations and can assist in the protocol specification refinement.
Altogether, we can conclude that the synthesizing data is suitable to form a representative model vocabulary for the significant part of the real-world data. However, according to the evaluation results that we observed, the absence of feature distribution information encoded as part of grammar or mined from real-world data at the synthesized time cannot guarantee accurate model training. Still, the training with synthesized data allows learning basic patterns and relationships between the input and output tokens, which can be further fine-tuned on the real-world data using just a few representative examples (few-shot training) to achieve the desired accuracy.
The performance of Modelizer depends on the accuracy and completeness of the producing grammar.
RQ7: Usage of Synthetic Data
Our next research question is RQ7: How much synthetic data is required for the behavior mocking of black-box programs? To answer these questions, we need to distribute the training set sizes into three clusters:
Small:
consisting of 5k, 10k, and 50k samples;
Medium:
containing 100k, and 250k samples; and
Large:
with 500k and 1M samples.
The results are summarized in Table 4 and Table 5. We see that the best results for models trained only on synthetic data were achieved with Medium-sized models. They scored the highest number of Exact Matches while also maintaining higher values for Close Matches and BLEU scores. However, after a few-shot fine-tuning with real data, the best-performing models were trained with Large datasets. Such a pattern is representative of half of the subjects we tested. More aggregated statistics for the remaining subjects can be found in Table 3.
Small and medium data sets suffice to obtain good Modelizer models.
RQ8: Generating Synthetic Data
Let us now go to RQ8: How should synthetic data be generated to achieve the highest accuracy of the learned models? To answer this question, we are going to explore the impact of the non-terminal expansion strategy on the accuracy of the learned models and refer to the representative results of Markdown to HTML conversion shown in Table 4 and HTML to Markdown conversion shown in Table 5. As we can see, increasing the complexity of synthesized tokens does not improve the conversion accuracy when the model is trained only on synthetic data. When we fine-tuned the model on a small set number of additional real-world samples that were not part of the test set, we measured an insignificant increase in prediction accuracy.
Without providing the feature distribution by assigning expansion probabilities, we assume the fuzzing budget gets wasted on randomly chosen expansions that do not align with the feature patterns observed in the real-world data. We can conclude that the efficient learning strategy would be to generate the mid-size synthetic dataset with a fixed number of non-terminal expansions, which, according to our results, is sufficient to cover the feature mappings between input and output, and then perform a few-shot fine-tuning on the real-world data, if available. This will achieve the highest or almost the highest possible accuracy of the learned models while maintaining the lowest possible time and resource consumption.
The highest Modelizer performance is obtained by combining a medium-sized synthetic data set with fine-tuning on real-world data.
RQ9: Hyperparameters
Our next research question concerns hyperparameters: How does Modelizer learn the hyperparameters for each data pair combination?
While Modelizer makes no assumptions on the availability of the real-world data at the training time, it still can perform hyperparameter optimization. For that, it needs a fixed set of test cases that are used to evaluate the model’s performance by computing the cross-entropy loss function after every search trial. The optimizer tries to minimize the value of the loss criterion and prunes non-promising trials to further speed up the search process. In our experiments, we reused the same set of test cases for forward and inverse models, which allowed us to study the reusability of discovered hyperparameters for the inverse models. Given the behavior mocking accuracy we have observed during the evaluation of real-world data, we think that running a hyperparameter parameter search on synthesized data is a suitable approach for cases when sufficient quality and quantity of real-world inputs are not available at the training time. According to the results of the hyperparameter search in Table 6, the optimizer managed to find an individual configuration for every translation pair type and the resulting number of model parameters is small, ranging between 2.7 and 10 million parameters for the given subjects.
Hyperparameter optimization is automatic and efficient.
Table 6. Automatically configured hyperparameter values using the Hyperparameter Search.
Model
Source Vocab Size
Target Vocab Size
Encoder Layers
Decoder Layers
Embedding Size
Feedforward Size
Attention Heads
Learning Policy
Learning Rate
Weight Decay
Dropout
Clip Gradients
Parameter Count
Markdown HTML 292 309 1 3 256 2048 32 cosine 0.0005 0.0100 0.0 False 6285621
HTML to Markdown 309 292 1 3 256 2048 32 step 0.0001 0.0100 0.0 False 6281252
SQL to KQL 674 675 1 2 256 2048 64 cosine 0.0001 0.0005 0.1 True 4992419
KQL to SQL 675 674 1 1 256 2048 64 cosine 0.0001 0.0001 0.0 False 3413410
Expression to 1501 656 1 3 256 2048 16 cosine 0.0005 0.0005 0.1 True 6773136
to Expression 656 1501 1 1 256 1024 16 multipl. 0.0005 0.0005 0.0 True 2782173
AsciiMath 122 131 1 1 256 2048 32 step 0.0001 0.0100 0.1 False 2993283
to AsciiMath 131 122 1 4 256 2048 64 step 0.0001 0.0100 0.1 False 7727226
MathML to 174 131 1 4 256 1024 64 cosine 0.0001 0.0005 0.0 False 5116291
to MathML 131 174 1 3 256 4096 64 cosine 0.0001 0.0100 0.1 False 10377646
MathML to AsciiMath 174 122 1 3 256 2048 64 step 0.0001 0.0100 0.0 False 6159482
AsciiMath to MathML 122 174 1 2 256 1024 64 step 0.0001 0.0100 0.0 False 3018158
RQ10: Reusability of Hyperparameters
In this research question, we want to check the reusability of the hyperparameters found with forward models for training the inverse models. To answer this question, we need to take a look at the results of Hyperparameter Search presented in Table 6. While the optimal embedding size value equal to 256 and the number of encoder layers equal to 1 were common for all input-output pairs, the rest of the hyperparameters were different. Additionally, we can mention that the best models were mainly trained with the lowest Learning Rate value; Linear Layer Size has to be at least four times larger than the Embedding Size; and, for the same data type pairs, inverse models require mostly two times more decoder layers than forward models. We can conclude that the hyperparameters for the forward models can serve as a lower bound for the hyperparameter search for the inverse models.
Hyperparameters from forward models can be reused for inverse models.
RQ11: Effect of Placeholders
Let us now turn to the effect of design decisions, and start with RQ11: How does the inclusion of additional enumerated placeholders into the model’s input/output vocabularies impact behavior learning?
The inclusion of the additional enumerated tokens into both model’s input/output vocabularies will impact the model’s configuration setup and increase the number of required neurons. Since the training data does not contain these injected tokens, the model will not be able to predict them. Modelizer will deactivate dead neurons, and they will not be used for the prediction. In cases, when the real-world samples contain the injected placeholders, the corresponding neurons still can be revived using the model fine-tuning. Such a technique is a cheap workaround to the problem of the incompatible feature distribution between the training and prediction inputs, when the input/output feature distribution is unknown at data generation time, or the protocol specifications that PUT implements get changed over time affecting only the quantity, but not the variety of these features. It helps to either completely avoid or delay the necessity to retrain the model, which is a time and resource-consuming process.
For several subjects that we tested like bidirectional conversion of the AsciiMath formulas to the representation and vice versa, the MathML formulas to the AsciiMath representation and vice versa, MathML formulas to the representation, the quantity of the synthesized placeholders given the used configuration of non-terminal nodes expansion for data generation was insufficient to represent all symbolic values that can be found in the test set. The preliminary oversizing of the vocabularies with the additional placeholders helps to avoid the model’s failure to predict test cases that contain unseen placeholders. Thus, if the selected hardware used for model learning can tolerate the additional memory allocation caused by the inclusion of additional placeholder tokens into the model’s vocabulary, we recommend using this technique to handle the processing of previously unseen input and output data, which is encoded with additional features, in a combination with fine-tuning.
Placeholders can significantly improve Modelizer performance.
RQ12: Effect of Tokenizers
Let us turn to tokenizers. How should the model input/output be tokenized to achieve the best learning performance? Should the pre-trained tokenizers used instead of domain-specific tokenizers, considering the higher development effort and cost?
The performance of the model is highly dependent on the quality of the generated tokens. The more tokens encode the same input/output data the more complex the model becomes, the more time it requires to produce the output, and the higher the chance it has to fail predicting a correct token sequence. Complex models require more powerful resources for functioning. At the same time, having a very small vocabulary also limits the model’s capability to learn a mapping between input and output correctly. In our work, we suggested using domain-specific tokenizers, which are developed by the domain experts and are capable of tokenizing the input/output data with the matching granularity. The custom tokenization pipeline gives full control over the tokenization process and allows to encoding of the input/output data in the most suitable way. The custom tokenizers are also capable of handling domain-specific data types and formats, which are not currently supported by the pre-trained tokenizers.
Table 7. Comparison of the tokenization quality between implemented domain-specific tokenizers and selection of pre-trained tokenizers.
Subject Tokenizer Average Length Vocab Size Length Ratio Vocab Ratio
Custom 12 176 - -
BERT 22 115 1.75 0.65
Markdown Llama3 22 136 1.84 0.77
Custom 16 245 - -
BERT 49 178 3.10 0.73
HTML Llama3 39 191 2.43 0.78
Custom 11 42 - -
BERT 20 38 1.93 0.90
SQL Llama3 20 42 1.93 1
Custom 14 42 - -
BERT 26 41 1.89 0.98
KQL Llama3 27 50 1.97 1.19
Custom 25 81 - -
BERT 47 102 1.87 1.26
Expression Llama3 42 103 1.69 1.27
Custom 57 87 - -
BERT 90 108 1.58 1.24
Expression Llama3 66 148 1.16 1.7
Custom 151 101 - -
BERT 342 77 2.30 0.76
MathML Llama3 296 101 1.99 1
Custom 27 104 - -
BERT 56 145 2.20 1.39
Formula Llama3 42 219 1.69 2.11
Custom 26 86 - -
BERT 51 111 2.14 1.29
AsciiMath Llama3 39 229 1.66 2.66
The pre-trained tokenizers are shipped with the state-of-the-art NLP models BERT and Llama3. Yellow cells indicate an evaluation run with a custom domain-specific tokenizer, blue cells indicate the evaluation runs of the BERT tokenizer, and purple cells indicate the evaluation runs of the Llama3 Tokenizer.
We compared the tokenization quality of the pre-trained tokenizers that are shipped with the state-of-the-art NLP models BERT (Devlin et al., 2019) and Llama3 (AI@Meta, 2024) with the domain-specific tokenizers that were implemented for each subject. We have computed the average length of the tokenized sequences and the vocabulary size per subject. To simplify the comparison we derived additional metrics, the length ratio and vocabulary ratio, which are computed as the ratio of the average length and vocabulary size of the pre-trained tokenizer to the domain-specific tokenizer. The results of our experiments are presented in Table 7. The evaluation was conducted on 1000 real-world samples for each subject, except for Markdown and HTML, where the tokenization was performed on all top-level elements extracted from 100 real-world documents. All inputs that were processed by each tokenizer were augmented by the PlaceholderProcessor module, which replaced the numeric and string literals with the matching placeholders.
While the pre-trained tokenizers managed to formulate a smaller vocabulary for half of the evaluated subjects, the average length of the tokenized sequences was significantly higher than the one produced by the domain-specific tokenizers. We demonstrate an example of the tokenization of the HTML element by the domain-specific and pre-trained tokenizers in Table 8. In this case, the domain-specific tokenizer produces a sequence consisting of 18 tokens, BERT produces 58 tokens, and Llama3 produces 48 tokens. Both BERT and Llama3 split the HTML tags and placeholders into multiple sub-tokens, which increases the length of the tokenized sequence. The domain-specific tokenizer, on the other hand, produces a shorter sequence of tokens that are more meaningful and easier to interpret.
The results of the evaluation in Table 7 show that the domain-specific tokenizers outperform the pre-trained tokenizers in terms of the average length of the tokenized sequences. The pre-trained tokenizers, on the other hand, do not require additional development effort, can tokenize multiple subjects, and are easily interchangeable. We still think that, currently, the domain-specific tokenizers can achieve better tokenization granularity, thus they are more suitable for better-quality mocking. At the same, we think that once a new type of pre-trained tokenizer that focuses on the tokenization of the program-specific data is released, program behavior modeling will become more accessible and easier to implement. Still, we currently suggest using the pre-trained tokenizers as a lower bound for the domain-specific tokenizers.
By repeatedly analyzing the token vocabulary for the test set and using the knowledge of domain experts and feedback from the PUT, one can potentially quicker come up with a domain-specific tokenizer implementation that will outperform the generic pre-trained tokenizers. We are currently working on the automation of this process and plan to reduce the development cost and effort of such tokenizers while keeping the tokenization quality high.
Table 8. Example of the tokenization of the HTML element by the domain-specific and pre-trained tokenizers.
Input
<p>TEXT_1<a href="URL_1">TEXT_2</a>TEXT_3<strong>
   TEXT_4</strong>TEXT_5<code>TEXT_6</code>TEXT_7</p>
Tokenizer Tokens
Custom
  <p>   TEXT_1   <a    href="URL_1"   >   TEXT_2   </a>
  TEXT_3   <strong>   TEXT_4   </strong>   TEXT_5
[…]
BERT
  <   p   >   text   _   1   <   a   hr   ##ef   =   "
  ur   ##l   _   1   "   >   text   _   2   <   /   a   >
  text   _   3   <   strong   >   text   _   4   <   /
  strong      >   text   _   5   <   code   >   text   _
  6   <   /   code      >   text   _   7   <   /   p   >
Llama3
  <p   >   TEXT   _   1   <a   Ġhref   ="   URL   _   1
  ">   TEXT   _   2   </   a   >   TEXT   _   3   <strong   >
  TEXT   _   4   </   strong   >   TEXT   _   5   <   code   >
  TEXT   _   6   </   code   >   TEXT   _   7   </   p   >Ċ
Domain-specific tokenizers (as provided by output grammars) are better than generic tokenizers.
5. Limitations
In this section, we want to discuss the limitations of the Modelizer framework and the current generation of sequence-to-sequence translation models. In particular, we want to address the following limitations:
Stateless and Stateful Programs.:
In the current study, we have explored programs that implement deterministic and stateless algorithms, including data format converters with recursive descent parsers. Such programs fit the nature and functionality of sequence-to-sequence models. In contrast, the modeling of stateful systems will require tracing and encoding additional program states as a part of the input data. Such a task implies further research and implementation of extensions to the framework to identify the meaningful fraction of the program’s state representation that should be used to model its behavior and the development of efficient state encoding and tokenization algorithms. We assume that the ability of the Modelizer framework to automatically validate and refine the learned models in combination with reinforcement learning algorithms (Sutton and Barto, 2018) will help to solve this problem. We plan to address this research topic in our future work.
Computational Complexity.:
Modelizer, will not be able to create models for all sorts of programs. The complexity of the whole program logic or its individual capabilities can be too high to be modeled by the sequence-to-sequence models. The transformer models themselves have a quadratic computational complexity. For the input sequence length $n$, hidden dimension size $h$, and number of layers in the neural network $L$, we get the following computational complexity:
Time $\textit{Time Complexity}:O\bigl{(}L\times(n^{2}d+nd^{2})\bigr{)}$
Space $\textit{Space Complexity}:O\bigl{(}L\times(n^{2}+nd^{2})\bigr{)}$
This means that modeling of extremely long input-output will be challenging and impractical. The models that are learned by Modelizer are not intended to replace programs. We see Modelizer as a tool that can help to understand the behavior of the program and assist in the program testing and debugging.
Context Window Size.:
Similar to many Transformer-based systems, Modelizer is sensitive to the length of the processed sequences. One of the factors that impact translation accuracy is the synthesis of training sequences. The grammar-based fuzzers used in the implementation of our Generators, by default, tend to produce short sequences. To overcome this problem we have introduced the sliding non-terminal expansion window, which partially resolves this issue. While the total token quantity per input and output sequences reached values similar to real-world samples, it does not guarantee a correlation for the input/output feature distributions between synthesized and real-world samples. At the same time, such a distribution can be learned from existing input samples (Soremekun et al., 2022). This approach introduces three requirements: such inputs have to be (1) available before input generation, (2) parsable with selected grammar, and (3) the correct values have to be set for the minimal and maximal number of non-terminal expansions in the generator. In the current study, we intentionally have not followed this approach to check Modelizer’s capabilities in learning behavior models of arbitrary systems given only the formal specification of such systems.
Program-Specific Input Output Formats.:
Certain subjects, specifically the output data types like KQL, required the implementation of custom parsers due to the unavailability of third-party parsers at the time this publication was prepared. The implementation of such parsers is a time-consuming and error-prone task. The third-party pre-trained tokenizers suboptimally partition the corresponding sequences, by unnecessarily splitting atomic tokens into sub-tokens similar to the constructions that are found in natural language sequences, misinterpreting certain characters like punctuation symbols, which cover part of the format syntax. Partially, such an issue can be mitigated by defining custom post-processing routines that will refine the outcome of the tokenization phase, but, such a list of regular expressions and string replacement rules can quickly become very long, create dependencies and conflicts in the order or application, what makes them hard to maintain. Such measures do not simplify the tokenization process but rather move the complexity from the tokenization phase to the post-processing phase. While the custom subject-specific parser-tokenizers are robust and produce high-quality tokens, we do not think that automated model learning problems can benefit from this approach in the long run, specifically for cases when complex system behavior has to be modelized by chaining the behavior models of sub-systems. We think that such a domain-specific tokenizer can be automatically created by wrapping existing tokenization strategies like byte-pair encoding, word-piece, or sentence-piece tokenization with a reinforcement learning loop. Alternatively, we are looking forward to solutions that would be able to mine the input-output grammars from the PUTs and automatically infer the parser-tokenizers from them.
Semantic Validity of Synthesized Inputs.:
Given the correct grammar definition, our generators which rely on grammar-based fuzzers, are very efficient in producing syntactically correct inputs. However, the PUT still can reject these inputs because grammar-based fuzzers cannot guarantee that the generated input will be semantically valid and pass all parsing guards in the PUT. Thus, an additional refinement of the synthesized data might be required. Steinhöfel et al. (Dominic Steinhöfel and Andreas Zeller, 2022) have proposed a technique and a formal language that allows manual specification of the semantic constraints inferred from the system specification, domain knowledge, and human expertise. Their approach, which builds on a modern Z3 SMT solver, can automatically refine the semantically invalid inputs using encoded constraints. However, fuzzing with semantic constraints is a computationally expensive task, and it slows down the training dataset formation on a large scale. One type of inputs that we synthesized, and which required such a refinement, were SQL queries with a JOIN condition. Using our SQL grammar, the GrammarFuzzer was able to produce SQL queries similar to the one given below:
SELECT COLUMN FROM TABLE JOIN TABLE ON TABLE.COLUMN = TABLE.COLUMN
Once PlaceholderProcessor updates all placeholders with unique identifiers the query starts to look like this:
SELECT COLUMN_1 FROM TABLE_1 JOIN TABLE_2 ON TABLE_3.COLUMN_2 = TABLE_4.COLUMN_3
The refined query is semantically invalid. To overcome this issue, we implemented an additional string refinement step, which parses the synthesized input and corrects the identifiers. While such an additional refinement step requires a bit of manual effort, for simple cases like this, it adds an insignificant performance penalty to the input generation process. Thus, the current implementation of the Modelizer framework does not support fuzzing with constraint solving, but we plan to support this functionality in future releases.
Some of the identified limitations can be resolved in the future releases of Modelizer.
6. Related Work
The evolution of language models and their integration into existing projects and solutions have opened new horizons in automating routine tasks and processes (Chen et al., 2021; Ziegler et al., 2022). The so-called general-purpose large language models (LLMs) like Llama 2 (Touvron et al., 2023) or T5 (Raffel et al., 2020) have been adapted to process code-related tasks by fine-tuning on a huge corpus of publicly available source codes and documentation (Code Llama (Rozière et al., 2024) or CodeT5 (Wang et al., 2021)). Such optimized code-specific models can be further fine-tuned on code generation and understanding downstream tasks. Below we discuss the most relevant applications of neural models in the context of program behavior modeling and the challenges they face:
Behavior Mocking.:
There have already been previous attempts to perform behavior mocking by training sequence-to-sequence neural networks. TransCoder (Roziere et al., 2020) represents two sequence-to-sequence models for cross-language translations of code written in C++, Java, and Python with additional support for back translations. Armengol-Estape et al. (Armengol-Estapé and O’Boyle, 2022) train transformer models to replicate the behavior of the GCC compiler. The training data is obtained by selecting C functions from the Anghabench benchmark and compiling them to x86 assembly code. The learned models can correctly compile the C code in less than half of the cases, making replacing compilers with neural machine Translation models invisible in practice. Zaremba et al., in their “Learning To Execute” publication (Zaremba and Sutskever, 2015), try to predict the execution results of short Python programs performing simple arithmetic operations and echo printing by training an LSTM (Hochreiter and Schmidhuber, 1997) model of a Python interpreter on a set of synthesized programs. These works have shown that the current generation of neural network architectures cannot correctly learn the full behavior of such complex programs as compilers or interpreters. While learned models cannot fully replace complex programs, our approach has demonstrated promising results in modeling linear or linearithmic programs. In contrast to previous approaches that rely on sequence-to-sequence models to achieve behavior mocking, Shen. et al and their Konure (Shen and Rinard, 2021) system relies on the active learning approach to mine the behavior model of applications that interact with relational databases. Konure synthesizes inputs with database content simultaneously, monitors the traffic exchange between the database and the application, and inspects the results of the database queries to continuously refine the application behavior hypothesis. Using the refined application model, Konure can generate the Python code for database querying that is semantically equivalent to the original application code, which might be written in a different programming language. Similar to Konure, our approach learns the program models from the input-output pairs. While Konure is designed to model the database interaction behavior of applications, our approach does not have such a limitation and can be used to either model the behavior of the whole system or its components and individual functionality.
Reversing Computation.:
Neural machine translation techniques have also shown promising results in low-level code decompilation tasks. Such neural decompilers are either implemented as sequence-to-sequence (Katz et al., 2018; Liang et al., 2021) sequence-to-tree (Fu et al., 2019; Cao et al., 2022) models. While existing works focus on converting the program’s binary representation to high-level program code, our framework allows users to reverse the direction of computation (predict the program input that was used to produce a given output) by training a behavioral model of the inverse program. This result is achieved by simply swapping training data pairs, i.e., using program output as model input and program input as model output. Even though such an inverse program may not exist, our approach still allows framework users to estimate the result of an inverse computation. Meanwhile, the approaches proposed by Katz et al. in the TRAFIX framework (Katz et al., 2019) and Fu et al., in the N-Bref framework (Fu et al., 2021), suggest training a neural decompiler from generated data. We think that these works are the closest to ours. In contrast, our framework supports behavior modeling of arbitrary programs, learning both direct and inverse models of the selected PUT, and supports further generalization of the learning process from mapping finite input-output pairs to extracting common input properties and constraints.
Data and Model Poisoning.:
Language models trained from crowd-sourced data suffer from data poisoning problems. In their research, Pearce et al. (Pearce et al., 2022) have studied the impact of training dataset poisoning on the quality of language model predictions. Code-generation models still tend to generate buggy or malicious code snippets because they have been trained on very similar corpora of publicly available documents. Schuster et al. (Schuster et al., 2021) have demonstrated two types of attacks on generative models: data poisoning and model poisoning. In the first scenario, only a few specially crafted samples, which are added to the training corpus, are sufficient to enforce code generation models to produce malicious code snippets. Alternatively, such models can be attacked by manipulating the training parameters (weights) or fine-tuning them on a few malicious samples. Both attack strategies will result in a deviation of generated sequences, and existing defense measures against dataset poisoning attacks cannot mitigate the problem completely (Wan et al., 2022; He and Vechev, 2023). Due to the possibility of manipulating model behavior, the model outcome cannot be trusted, and their application in mission-critical systems is still questionable. In contrast to other approaches that rely on crowd-sourced training data, our framework allows full control over the data collection and model training processes through the specification of input synthesis rules, testing of generated samples with the PUT, and filtration or sanitization of faulty, undesired, and dangerous samples.
Domain-specific Adaptations.:
Due to the reason that such domain-specific models were fine-tuned only on the program source codes, they cannot capture the program behavior patterns without further fine-tuning. Such fine-tuning could be an expensive process due to the sizes of models and specific hardware requirements. Zheng et al. and their LlamaFactory citezheng2024llamafactory framework provide estimated memory requirements implied in fine-tuning open-source LLMs. For example, the small versions of commonly used LLMs, like Llama2, Llama3, and Mixtral, with 7-to-8 billion parameters require more than 60GB of memory for full-precision fine-tuning. Even though models are open-source, it is mostly impossible to fine-tune them locally without access to data-center-grade hardware. Instead, our framework can learn “lightweight” program behavior models consisting of only several million parameters and can be trained on commodity hardware.
Neural Program Induction.:
We find that the neural program induction and neural program synthesis fields are also closely related to our research. While neural program induction models (Joulin and Mikolov, 2015; Devlin et al., 2017a; Neelakantan et al., 2016; Graves et al., 2014; Zaremba et al., 2016; Kurach et al., 2016) try to train a sequence-to-sequence generative model that learns data patterns from input-output examples, neural program synthesis (Vijayakumar et al., 2018; Devlin et al., 2017b; Shin et al., 2019) use neural networks for predicting algorithms from corresponding input-output samples. Existing studies cannot capture the behavior of complex algorithms and programs because they had been relying on already outdated neural network architectures; they could not generalize to arbitrary complex inputs because the corresponding models had been trained on a limited number of training samples, which, for certain studies are in human-generated; require the additional manual modeling and encoding of the algorithm or program in a domain-specific language that has limited functionality. Recent advances in sequence-to-sequence prediction algorithms (Transformer architecture and its variations) make it possible to automatically infer complex data patterns from input-output pairs. Moreover, our approach can also automatically learn the inverse patterns, which allows the revision of the computation results of certain programs at a small cost.
AI-driven Testing.:
AI-driven, or specifically LLM-driven software testing is becoming more and more popular. We picked a few representative works from myriads of publications in this field to highlight the s TOGA (Dinella et al., 2022) project explores neural test oracle generation problem by fine-tuning the CodeBERT (Feng et al., 2020) model on automated inference of exceptional and assertion oracles from developer-written tests. Haluptzok et al. (Haluptzok et al., 2023) synthesize and solve Python programming puzzles using language models. The authors of DeepAnalyze (Shetty et al., 2022) learned a sequence labeling model to automatically mark memory frames in a stack trace that caused the crash of an application. The LIBRO (Kang et al., 2023) framework uses the Codex Large Language Model to generate and rank bug-reproducing test cases from bug reports. Tufano et al. (Tufano et al., 2019) train a neural machine translation model to automatically replicate in 36% of cases the changes between pull requests in code repositories like bug fixes or code refactoring. The DeepFix (Gupta et al., 2017) approach randomly mutates correct programs to introduce errors. It uses these data pairs to train the neural machine translation model for automated program repair, which succeeds in fixing buggy programs in 27% of cases. The Break-It-Fix-It (Yasunaga and Liang, 2021) approach achieves high accuracy in code repair tasks by chaining two generative neural machine translation models for synthesizing naturally looking code errors and training a fixer model from synthesized pairs. Godefroid et al., in their “Learn&Fuzz” publication (Godefroid et al., 2017), learn a generative model for PDF structures by training a recurrent neural network from a set of PDF objects. The input sequence is represented by a sequence of characters in PDF objects, while the output sequence is obtained by shifting the corresponding input sequence by one character. While every mentioned work follows its unique approach, they all require a significant amount of human effort for problem-solving. In contrast, Modelizer tries to automate the process of data collection (which is one of the most time-consuming parts of model preparation) and model training. The minimalistic set of requirements, in particular the ability to run a Python interpreter locally, simplifies the integration of the learned models into the existing projects and testing tools.
AI-generated Synthetic Data:
The capabilities of LLMs revealed several approaches of their application for synthetic data generation (Li et al., 2023). They can generate human-like text or dialogues given the prompt query, augment existing datasets by producing variations of the original data, and produce structured representations of existing data like JSON or CSV. Such AI-generated synthetic data is mostly used for training classifiers or question-answering systems. For example, general-purpose LLMs like OpenAI ChatGPT or Google Bard have shown promising results (İrfan AYGÜN, 2023) in generating synthetic data in a medical text production task. However, in certain tasks, like hateful speech detection and classification, the inclusion of syntactically generated samples to the training set does not improve the performance of classifiers (Kruschwitz and Schmidhuber, 2024). Still, it is an active and promising field of research, and industry giants like Nvidia invest in the development of LLMs, like Nemotron-4 (Nvidia et al., 2024), specifically designed for synthetic data generation. While synthetic data generation with LLMs implies specific hardware requirements to run models locally or lead to a privacy compromise while processing the data with a third-party cloud provider at an additional service cost, our approach that relies on synthesizing data with grammar fuzzers only requires a system with an ability to run a Python interpreter locally.
Data Generation by Mutating Inputs:
We also want to address the alternatives to grammar-based and AI-based input generation strategies. Potentially, an input set for the given PUT can be automatically constructed with the help of mutation-based fuzz-testing (Barton P. Miller et al., 1990). A mutation fuzzer, like AFL++ (Andrea Fioraldi et al., 2020), can accept one input for a PUT as an initial seed and produce infinitely many new inputs by applying pre-defined mutation strategies, like bit flips, addition, or subtraction of integers, on different parts of the given seed. Such a technique produces two types of mutants. The first type of mutants has a broken syntactical or semantic structure. These mutants will either crash the PUT if it contains a bug in its implementation, or the input parser guards will reject such input, and program execution will be terminated. The collected input-output pairs will only cover the behavior of the input parser rather than the whole program. In most cases, the second type of mutants, which would have managed to bypass the parser guards, will contain insignificant differences when compared to an initial input sample. Thus, testing the program with this kind of input will unlikely produce a diverse set of outputs. While fuzz testing is helpful for bug discovery in the implementations of input parsers, we do not think the mutation-based input generation technique is applicable to automate program behavior modeling.
Mining Input Grammars:
In certain cases, input grammars can also be automatically mined from a PUT. The Mimid approach by Gopinath et al. (Rahul Gopinath et al., 2020) relies on the program instrumentation in combination with string wrappers to identify the respective control flow nodes that access the relevant portions of a program input. Kulkarni et al. presented the Arvada algorithm (Kulkarni et al., 2021) that can automatically learn a highly recursive CFG from a given program using a set of positive examples and a Boolean-valued oracle. Currently, if input specifications are not publicly available or are not recoverable from the PUT, then the user of our framework has to encode it in CFG format manually.
7. Conclusion and Future Work
With Modelizer, we provide a principled approach to turning regular programs into models. This opens up the manifold capabilities of machine learning models for several Software Engineering tasks such as mocking, testing, and debugging. The maybe most promising feature of the learned translation models is their ability to predict inputs for given outputs, not only reversing executions but also offering new opportunities for test input generation. We see Modelizer as a pioneer in project-specific learning—rather than trying to generate expensive LLMs that may or may not evolve into a one-size-fits-all, Modelizer demonstrates that it is possible to learn small models for nontrivial programs that can successfully replicate their behavior—and their reverse.
Besides testing Modelizer with more complex programs and larger model sizes, our future work on Modelizer will focus on further applications, as well as addressing some of its current limitations. Specifically:
Targeted input generation.:
One of the most promising applications of Modelizer is its use for test input generation. By treating traces or coverage as output of the PUT (and supplying an output grammar that can parse trace logs), Modelizer can learn which input features correlate with which traces or coverage features; and can then again predict which input would be required to achieve a particular trace or coverage. These inputs could then be tested automatically, and if the trace or coverage deviates, such feedback could be used to fine-tune the model (see below).
Monitoring.:
Once Modelizer has extracted a model, one can also use it for monitoring the PUT, continuously comparing the program behavior against the extracted model. This could be used to detect and prevent behavior not seen during testing.
Debugging.:
By learning correlations between collected crash data (core dumps and stack traces) and input features, Modelizer could predict inputs and input features likely to have caused a given program state.
Smarter test generation.:
In our evaluation, we have found again and again that the performance of Modelizer very much depends on the quality and diversity of the synthesized training data, and thus on the quality of the test generator. In our future work, we will experiment with smarter black-box test generators, covering grammar features and combinations thereof (Havrikov and Zeller, 2020), but also smarter grey-box fuzzing, using code coverage in the PUT as guidance (Pham et al., 2019; Aschermann et al., 2019).
Stateful programs.:
At this point, the models extracted by Modelizer do not consider the internal state, as, say, a server could maintain. Let us assume the PUT starts in a known initial state, and the input to the PUT is a sequence of commands. Then, having Modelizer learn the effects of command sequences should also give it the ability to detect that certain outputs depend on specific subsequences of commands (implying a specific state). Such reasoning will likely require larger models; we plan to evaluate Modelizer on more complex and stateful programs.
Feature distribution.:
The feature distribution in the synthesized samples does not always correspond to real-world counterparts. Our attempt to add the sliding window strategy to the number of selected non-terminals and manually adjust probabilities of expansion rules only partially solves this problem. Without having the precise input specification in advance, it is not possible to initialize the generator to produce samples that fully cover the real-world input space. This problem can be resolved by first mining corresponding expansion probabilities from a few real-world samples and using them to fine-tune the model on additionally synthesized samples (Soremekun et al., 2022). Such samples will serve as guidance for the generator to focus on the uncovered areas of the input space.
Leveraging program feedback.:
We estimate that the model’s performance can be further improved by utilizing the feedback from the PUT, when available. In contrast to fundamental LLMs, Modelizer framework allows relatively cheap continuous re-training of the learned models. If the specifications of the target system have insignificantly changed, only affecting the system behavior, but not extending the population of the input-output vocabularies with new elements, the learned models can be fine-tuned with a few samples collected from system executions. In Section 2.4, we have represented a scenario when the PUT serves as an oracle to validate the model’s predictions. Once the model prediction does not align with execution results, the collected feedback can be passed to the model as additional training data to adjust the model’s parameters. The model has to be repeatedly re-evaluated and fine-tuned till the model’s predictions match the PUT’s execution results. We plan to study the capabilities of Modelizer and respective learned models to serve as an oracle for automatic detecting of the behavior deviation cases between various revisions and implementations of the same protocol. Also, we hope to explore the possibility of forging feedback-driven tokenizers, which will be able to automatically adjust the tokenization granularity and possibly mine the input-output grammars from black-box systems. They will rely not only on the feedback from the PUT but also leverage the feedback from the model evaluation.
1-bit models.:
Lastly, we are excited by the idea of replacing 16-bit parameter weights with a 1-bit representation (Ma et al., 2024), which significantly reduces the memory requirements, and thus allows training more cost-effective models. We plan to integrate this feature into Modelizer soon.
Modelizer and all experimental data are available at
https://github.com/2ral/Modelizer
Acknowledgments. We appreciate the thorough review and critical insights provided by our colleagues, which significantly improved the quality of this manuscript.
This work is funded by the European Union (ERC S3, 101093186). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@
 Abstract.
-We introduce Modelizer—a novel framework that, given a black-box program, learns a model from its input/output behavior using neural machine translation.
+We introduce Modelizer—a novel framework that, given a black-box program, learns a model from its input/output behavior using neural machine translation algorithms.
 The resulting model mocks the original program: Given an input, the model predicts the output that would have been produced by the program.
@@
 Finally, the model is differentiable and can be efficiently restricted to predict only a certain aspect of the program behavior.
-Modelizer uses grammars to synthesize inputs and to parse the resulting outputs, allowing it to learn sequence-to-sequence associations between token streams.
-Other than input and output grammars, Modelizer only requires the ability to execute the program.
+Modelizer uses grammars to synthesize and inputs and unsupervised tokenizers to decompose the resulting outputs, allowing it to learn sequence-to-sequence associations between token streams.
+Other than input grammars, Modelizer only requires the ability to execute the program.
 The resulting models are small, requiring fewer than 6.3 million parameters for languages such as Markdown or HTML; and they are accurate, achieving up to 95.4% accuracy and a BLEU score of 0.98 with standard error 0.04 in mocking real-world applications.
-We foresee several applications of these models, especially as the output of the program can be any aspect of program behavior.
-Besides mocking and predicting program behavior, the model can also synthesize inputs that are likely to produce a particular behavior, such as failures or coverage.
+As it learns from and predicts executions rather than code, Modelizer departs from the LLM-centric research trend, opening new opportunities for program-specific models that are fully tuned towards individual programs.
+Indeed, we foresee several applications of these models, especially as the output of the program can be any aspect of program behavior.
+Beyond mocking and predicting program behavior, the models can also synthesize inputs that are likely to produce a particular behavior, such as failures or coverage, thus assisting in program understanding and maintenance.
 Software Testing, Mocking, Deep Learning
@@
 However, like all machine learning models, Modelizer depends on good training.
-This raises two central problems, for each of which Modelizer has a solution:
-Synthesizing inputs.:
-The first central problem thus is the availability of input data for training.
+This raises data availability problem, for which we need to provide a solution.
 In Software Engineering, the field of test generation discusses how to obtain input data for a program.
@@
 In a black-box setting, obtaining valid and diverse input data requires a dedicated generator or producer, leveraging knowledge about the input language of the program.
-Tokenization.:
-The second problem is obtaining features from inputs and outputs.
-Modelizer uses a sequence-to-sequence translation model to translate token sequences from inputs into outputs.
-Hence, we need a decomposition of inputs and outputs into individual tokens, so we can learn how token sequences map into each other.
-While this is a mostly solved problem for natural languages, tokenizing formal inputs correctly is a challenge, notably as we want the tokenizer to be general-purpose, and suitable for all sorts of input and output languages.
-Modelizer addresses both problems by leveraging grammars for both input and output language.
+Modelizer addresses the data availability problem by synthesizing inputs with the help of grammars - formal languages for generating strings.
 The input grammar is used to produce syntactically valid and diverse inputs; it can also parse existing inputs into derivation trees.
 In both cases, we know the structure of the input.
-The output grammar is used to parse the output, again delivering a tree structure allowing decomposition into tokens.
+The output grammar can also be used to parse the output, again delivering a tree structure allowing decomposition into tokens; if the output language specification is unknown, the program output can be decomposed by automatically training a tokenizer from scratch using the SentencePiece (Kudo and Richardson, 2018) algorithm.
 Using the input grammar as a producer, we can synthesize valid and diverse input data automatically and thus obtain as many input-output pairs from the PUT as we want.
-In cases when output language grammar is unavailable, the program output can be decomposed by unsupervised tokenizers.
-Modelizer distinguishes structure from content in the decomposed sequences and can abstract content by replacing them with placeholders to improve the learning quality.
+Additionally, Modelizer distinguishes structure from content in the decomposed sequences and can abstract content by replacing them with placeholders to further improve the learning quality and efficiency.
 Developers who want to use Modelizer thus only need to provide the PUT in an executable form, as well as grammars that describe its input and output languages; such grammars can also be inferred from the PUT or input samples.
@@
 We evaluate Modelizer across a range of challenging PUTs and achieve excellent accuracy.
-Notably, the models produced by Modelizer showed high accuracy on non-trivial programs despite having a small number of neurons (and thus efficient to learn).
-On the Pandoc markup converter, for instance, Modelizer achieved over 90% accuracy in converting Markdown to HTML and vice-versa with a BLEU score of 0.71.
-While complex computations as performed by or many commercial programs remain out of reach for our current methods, our experiments indicate that increased model and training sizes may make such learning possible in the future.
+Notably, the models produced by Modelizer showed high accuracy on non-trivial programs despite having a small number of neurons (and thus efficient to learn) when compared to large language models (LLMs).
+For instance, on the Pandoc markup converter, Modelizer achieved over 90% accuracy in converting Markdown to HTML and vice-versa with a BLEU score of 0.71.
+Despite the potential of Modelizer, we also identify several limitations:
+  • Some complex programs remain out of reach for our current methods due to the computational complexity of the underlying learning algorithm which scales quadratically relative to the processed data length.
+The legnths of tokenized sequences with input-output data can vary significantly depending on the selected data decomposition and abstraction strategies.
+For example, intermediate abstraction layers can significantly shorten sequence lengths but can cause context loss, which will affect the prediction quality.
+Existing sequence-to-sequence translation models tend to make mistakes when predicting long sequences.
+Thus, lack of prior knowledge about the subject properties and hardware limits makes it hard to pre-specify all tasks where Modelizer will succeed.
+  • We still require minor developer effort in configuring the Modelizer framework.
+Our experiments suggest that improving synthetic input generation is crucial for automating behavior extraction.
+Feedback from real program runs can already help to automate grammar tweaks, reconfigure fuzzers, and adjust learners’ hyperparameters.
+Continuous growth of the computational power will soon enable complete behavior extraction of any subjects.
 In summary, our contributions are as follows:
 A novel, generic means to extract behavior models from programs.:
-To the best of our knowledge, Modelizer is the first approach to automatically learn sequence-to-sequence models from arbitrary programs.
+To the best of our knowledge, Modelizer is the first approach to automatically extract execution behavior from arbitrary programs by converting them into sequence-to-sequence translation models.
 Using grammars to produce inputs for training.:
-Modelizer uses grammars to produce inputs (and parse given inputs), and can thus learn from an unbounded number of training inputs and resulting outputs.
-Abstract placeholders for tokenizers.:
-Tokenizers that implement the suggested interface can automatically replace the real content with abstract placeholders, thus focusing on translating text structure rather than text content.
+Modelizer uses grammars to produce inputs (and parse given inputs), and, thus, can learn from an unbounded number of training inputs and resulting outputs.
 Predictive behavior models with high accuracy.:
@@
 In Section 4, we evaluate Modelizer on a variety of text-to-text translators, such as Markdown to HTML, query languages, and more.
-Section 5 discusses the limitations of Modelizer.
-After discussing related work (Section 6), Section 7 closes with a conclusion and future work.
+Section 5 discusses the limitations of Modelizer; Section 6 details the related work.
 Modelizer and all experimental data are available as open source; see Section 7 for details.
 2. Approach
-Figure 2. How Modelizer works. Modelizer tests program with synthesized inputs and automatically learns reversible program behavior model from extracted input-output pairs.
-[…]
-How Modelizer works. Modelizer tests program with synthesized inputs and automatically learns reversible program behavior model from extracted input-output pairs.
 Modelizer is a generic framework for learning program behavior.
@@
 Figure 2 shows a general overview of the framework.
-Given the input specification for the selected PUT, Modelizer automatically generates the required number of inputs and passes them to the PUT for processing.
+Given the input specification of the selected PUT encoded as grammar, Modelizer automatically generates the required number of inputs and passes them to the PUT for processing.
 Once the PUT produces the corresponding output, Modelizer extracts these input-output pairs and uses them to train a reversible neural machine translation model.
@@
 The following sections will cover each step in detail.
+Figure 2. How Modelizer works. Modelizer tests the program with synthesized inputs and automatically learns a reversible program behavior model from extracted input-output pairs.
+[…]
+How Modelizer works. Modelizer tests the program with synthesized inputs and automatically learns a reversible program behavior model from extracted input-output pairs.
 2.1. Input Generation
-While developing our framework, we aimed to make it interoperable with different types of programs.
+While developing the Modelizer framework, we aimed to make it interoperable with different types of programs.
 Even though all programs are written in different programming languages, differ in implementation logic, and target different hardware architectures and operating systems, most programs are made to process input coming from users or other programs.
-Thus, we have decided to learn program behavior by training our behavior mocking models from the program inputs and processed outputs.
-Of course, such a strategy will enforce certain limitations:
-  • The program behavior should be deterministic. At every execution, the program should produce the same output for the given input.
-  • We will not be able to model the behavior of programs that process extremely long inputs or outputs at once.
-The computational complexity, memory, and time constraints scale quadratically with the input-output length.
-  • Learning accurate inverse models that predict inputs for the given output is challenging when PUT computes multiple inputs to the same output.
-An intermediate representation for inputs could assist in such cases, which leads to context loss.
+Thus, we have decided to learn the program behavior by training our behavior-mocking models from the program inputs and processed outputs.
 To learn an accurate model, we need sufficiently many diverse input-output pairs.
-Since public test datasets are typically biased towards commonly used features, and we do not know in advance how many inputs are required for our approach to learn the behavior of an unseen program, Modelizer assumes that inputs are automatically generated.
+Since public test datasets are typically biased towards commonly used features, and we do not know in advance how many inputs are required for learning the behavior of an unseen program, Modelizer assumes that inputs are generated automatically.
 2.1.1. Input Generation with Grammars
+Figure 3.
+Input generation pipeline.
+Modelizer automatically synthesizes and validates unique inputs from given specification.
+Input generation steps: (1) Synthesis, (2) Post-Processing, (3) Hashing, and (4) Validation.
+[…]
+Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification.
 To cover the entire input space (and hence increase our chances to also cover the behavior and output spaces), we use grammar-based input producers (Zeller et al., 2023).
@@
 For example, the Grammar Zoo (Zaytsev, 2015) repository contains definitions of 1756 grammars of mainstream programming languages mined from standard documentation or community-created resources.
-Figure 3. Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification
-[…]
-Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification
-The first major component of the Modelizer framework is the Input Generator Module.
-This module functions as a data processing pipeline, which is sketched in Figure 3.
-It contains four steps: (1) Synthesis, (2) Post-Processing, (3) Hashing, and (4) Validation.
-At the synthesis step, the input producer synthesizes new input samples according to the rules encoded in grammar.
-Then, the synthesized samples get refined to better align with real-world data.
-Such a refinement helps to map synthesized input features to real-world data features without increasing the size of grammars and, thus, simplifies the model learning process.
+Developers could handwrite such grammars using the knowledge provided as protocol specifications such as RFCs.
+We cover alternative ways of automatically extracting grammars from existing systems in Section 6.10.
+The Input Generator Module is the first component of the Modelizer framework.
+It acts as a data processing pipeline, as shown in Figure 3, consisting of four steps: (1) Synthesis, (2) Post-Processing, (3) Hashing, and (4) Validation.
+At the synthesis step, the input producer synthesizes new input samples according to rules encoded in grammar.
+Then, the synthesized samples are refined to better align with real-world data.
+Such refinement helps map synthesized input features to real-world data features without increasing grammar size, thereby simplifying the model learning process.
 The refined samples are then checked for uniqueness by computing a hash of the synthesized sample.
-If the hash value is not present in the Hashes Dataset, the sample is passed to the PUT for further processing.
-Otherwise, all steps are repeated till a unique sample is generated.
-The PUT produces the corresponding output, which is saved in a persistent storage along with the input.
+If the hash is absent in the Hashes Dataset, the sample is passed to the PUT for further processing.
+Otherwise, all steps are repeated until a unique sample is generated.
+The PUT produces the corresponding output, which is saved in persistent storage with the input.
 The collected input-output pairs are later used to train a program behavior model.
@@
 […]
-14<code_block> ::= ‘<text>‘ | ‘‘‘<text>‘‘‘ | "<empty_space><text>"
-15<quoted_text> ::= ’<text>’ | "<text>"
-16<text> ::= TEXT
-17<url> ::= URL
+[…]
+15<code_block> ::= ‘<text>‘ | ‘‘‘<text>‘‘‘ | <empty_space><text>
+16<quoted_text> ::= ’<text>’ | "<text>"
+17<empty_space> ::= " "
+18<text> ::= TEXT
+19<url> ::= URL
 Listing 1: Markdown grammar used for input generation (excerpt)
@@
 We will follow the steps that Modelizer takes to produce inputs.
-You can find the high-level implementation of the input generation pipeline written in Python in LABEL:lst:generate.
+LABEL:lst:generate shows the high-level implementation of the input generation pipeline written in Python.
 Once the input producer is configured, it is ready to synthesize as many inputs as needed (Step 1).
-Below we give an example of synthesized input that was generated by the producer:
+Here is an example of synthesized input that was generated by the producer:
 […]
-1def generate_data(grammar, min_terminals, max_terminals, total_elements, program_handler):
+1def generate_data(grammar, min_non_term, max_non_term, total_elements, program_handler):
 2 dataset = list()
@@
 4 processor = PlaceholderProcessor(grammar)
-5 fuzzer = GrammarFuzzer(grammar, min_terminals, max_terminals)
+5 fuzzer = GrammarFuzzer(grammar, min_non_term, max_non_term)
 6 while(len(dataset) < total_elements):
@@
 Listing 2: Input generation in Modelizer (pseudocode)
-The producer has synthesized one top-level Markdown element, which is a paragraph block.
-The paragraph block was expanded to a single paragraph text element which respectively contains a sequence of formatted text, hyperlink, and codeblock elements.
+The producer has synthesized one top-level Markdown element, a paragraph block.
+The paragraph block was expanded to a single paragraph text element, which contains a sequence of formatted text, hyperlink, and code block elements.
 One of the formatted text elements was further expanded to a bold text element.
 The synthesized input contains TEXT and URL placeholders instead of real character sequences.
-We have decided to define placeholder tokens for replacing the real-world data instead of encoding the whole language vocabulary into our grammar and synthesizing natural language sentences.
-We will elaborate more about placeholders in the follow-up section.
+Instead of encoding the whole language vocabulary into our grammar and synthesizing natural language sentences, we have decided to define placeholder tokens to replace the real-world data.
+(We will elaborate more about placeholders in the follow-up section.)
 2.1.2. Placeholders
-Synthesizing the data that is equivalent to real-world counterparts is a challenging task.
-Different programs can process the same data in various ways.
-This means that the grammar must be specified or refined according to program expectations.
-Depending on the complexity of input specification, the number of required input samples for efficient behavior learning can also become extremely large.
-We have figured out that for certain data types and programs, some fragments that are not affected or used by underlying implementation logic can be abstracted.
+Synthesizing data that is equivalent to real-world counterparts is a challenging task.
+As different programs can process the same input in various ways, the input grammar must be specified or refined according to program expectations.
+Depending on the complexity of the input specification, the number of required input samples for efficient behavior learning can also become extremely large.
+In their work on abstracting failure-inducing inputs, Gopinath et al.
+(Gopinath et al., 2020a) found that some input fragments are not affected or used by underlying implementation logic and can be abstracted for certain data types and programs.
 For example, many text format converters only process the document structure, leaving the document content untouched.
-To benefit from such effects, we suggested and implemented two strategies in our input generation pipeline:
+To benefit from similar effects for behavior extraction, we integrated the following two strategies in our input generation pipeline:
   • We suggest generalizing certain low-level non-terminal expansion rules by replacing a fraction of non-terminal to terminal expansions with generic placeholders.
   • We have added an optional post-processing step that augments placeholders with unique identifiers.
-Both strategies will help to capture large portions of possible input space with few samples and correctly map real-world inputs to generic representations and vice-versa during the inference phase.
+Both strategies will help capture large portions of possible input space with few samples and correctly map real-world inputs to generic representations and vice versa during the inference phase.
 Instead of learning the behavior from real-world examples, we learn how the PUT processes placeholders.
+(In the current Modelizer implementation, developers have to specify how to abstract grammar rules; future revisions will automate this process.)
 Let us refer back to the Markdown synthesis example and follow the input generation phases.
@@
 Similarly, a URL placeholder can substitute an arbitrary definition.
-By introducing such an abstraction layer, we can cover the whole space of structures of real-world documents with a relatively small number of synthesized samples.
-While the input producer’s output is already well-formed and will pass the input parsing guards, it can be challenging to align real-world Markdown documents with the synthesized structure because it is impossible to distinguish the generated placeholders from each other.
-Therefore, we have included a post-processing phase in the input generation pipeline, for which we have implemented a generic Placeholder Processor.
-It must be initialized with the list of Placeholders that are encoded into selected grammar, like [TEXT, URL], and can automatically assign a unique identifier to every placeholder that it finds in the given sequence.
-Of course, it is possible to inject identifiers during the input generation.
-However, grammar-based input generators do not synthesize integers incrementally.
-They cannot guarantee that the generated identifier will be used in a synthesized sequence of characters only once.
-It is much faster and easier to generate a simple candidate sequence and then refine it using custom string-replacement or constraint-solving rules.
-The presence of unique identifiers improves the accuracy of both input-to-output and output-to-input translations, specifically in the cases when elements must be reordered during translation.
-During the tokenization process, which we will cover in the following sections, the respective data portions are assigned with a matching placeholder type that already includes a uniquely generated identifier.
-Optionally, the assignment of unique identifiers can be disabled.
-Since grammar-based producers do not maintain a list of generated samples, we have implemented a sample validation procedure which first computes an HMAC-SHA-384 (Hansen and 3rd, 2011) value (Step 3.1) for the synthesized sample and then checks whether the computed hash value is stored in the Hashes Dataset (Step 3.2).
-If a computed hash value is absent in the dataset, the dataset gets updated with a new record (Step 4) and the synthesized input is passed to the PUT to compute output.
-Otherwise, the synthesized sample is dropped, and all input generation steps are repeated till the unique sample is synthesized.
-Corresponding input-output pairs are saved in a persistent storage for later processing.
-The collected input-output pairs are saved in a persistent storage for later processing.
+Learning from the small number of synthesized structures, our models will still be capable of processing principally equivalent real-world inputs.
+While the input producer’s output is well-formed and passes parsing guards, aligning real-world Markdown documents with the synthesized structure is challenging because it is hard to distinguish the generated placeholders.
+We have implemented a post-processing phase in the input generation pipeline as a generic Placeholder Processor module.
+It initializes with a list of encoded placeholders like [TEXT, URL] and augments each placeholder with a unique identifier.
+Such identifiers enhance the accuracy of bidirectional translations, especially when elements need reordering.
+Since grammar-based producers do not maintain a list of generated samples, we implemented a sample validation procedure.
+This involves computing an HMAC-SHA-384 (Hansen and 3rd, 2011) value (Step 3.1) for the synthesized sample and checking if it’s stored in the Hashes Dataset (Step 3.2).
+If absent, the dataset gets updated with a new record (Step 4), and the synthesized input is passed to the PUT for output computation.
+Otherwise, the sample is dropped, and input generation repeats until a unique sample is created.
+Input-output pairs are saved in persistent storage for later processing.
 2.2. Dataset Pre-processing
-Machine translation is the task of automatically translating a text from one language to another.
-Neural machine translation models are the latest iteration of machine translation approaches that rely on deep neural networks to learn dependencies and patterns between sentences in the source and target languages.
-The ability to learn these dependencies from matching input-output pairs inspired us to reason about the program behavior modeling as a sequence-to-sequence translation task.
-In principle, the current generation of neural machine translation architectures that we use in our implementation can operate with the whole input at once.
-The computational complexity of these systems scales quadratically concerning the length of the input, and practical implementations do not follow this principle.
-Instead, text-to-text translations are performed by partitioning the whole input into sequences of small elements called tokens.
-Then, neural text translation models learn how to generate sequences of tokens in a target language by mapping the cross-token dependencies from a source language to a target language.
-Similarly, the synthesized input and computed output must be converted into batched tensors for the Learner to be able to process them.
-While text partitioning is a well-studied problem in the natural language processing domain (Manning et al., 2008), and there are already tools that can perform this task (Honnibal and Montani, 2017), we find the tokenization of program input and output to be a challenging task.
-2.2.1. Pre-trained and Unsupervised Tokenizers
-Instead of implementing tokenizers from scratch, one can potentially rely on existing generic tokenizers.
-For example, the HuggingFace repository (Wolf et al., 2020) provides access to many community-developed natural language tokenizers through the HuggingFace API.
-Such tokenizers cannot properly split the program data for several reasons.
-First, they ignore the data-specific or computation-specific features.
-Both factors impact the tokenization granularity, which is covered in the following paragraphs.
-Also, some tokenizers, like the pre-trained Llama3 tokenizer, modify the input by prepending or appending additional characters to the produced tokens.
-This can break the syntactical correctness when it comes to program data and can impact the quality of learned models.
-While any input and output can be encoded as a sequence of characters, semantically it does not necessarily need to be textual data.
-Programs can accept or produce structural data.
-Such a structure contains knowledge that is relevant for behavior learning, but many existing tokenizers cannot distinguish the data structure from the content.
-Since, for certain computation tasks, like converting Markdown-to-HTML or SQL-to-KQL, the program is going to process the structure of the input while leaving content unchanged, a tokenizer needs to be able to map the fragments of the input to an abstract representation while processing.
-It is possible to augment the tokenization process of pre-trained tokenizers with additional pre and post-processing steps, where inputs can be transformed into semantically similar representations with placeholders replacing the fragments containing content.
-To build such a pipeline, the parser for the given data type has to exist and align with the input parsing rules of PUT.
-Still, the program behavior learning approach is tokenizer-agnostic, and an unsupervised tokenizer like Google SentencePiece (Kudo and Richardson, 2018) is sufficient to quickly start the behavior learning process.
-However, the implementation of a data-specific tokenizer will help to further improve the quality of the learned model.
+Neural machine translation models are artificial intelligence models that use deep neural networks to learn dependencies between source and target language sentences.
+Their success in automated machine translation tasks inspired us to model program behavior as a sequence-to-sequence translation task.
+In principle, current neural machine translation architectures can process arbitrarily complex inputs, but the problem’s computational complexity scales quadratically with the input length.
+So, practical implementations partition data into token sequences and process them in fragments.
+Then, models learn how to generate sequences of tokens in a target language by mapping the cross-token dependencies from a source language to a target language.
+Similarly, the program input and output must be converted into sequences of tokens for processing.
+While text partitioning is well-studied in natural language processing  (Manning et al., 2008; Honnibal and Montani, 2017), tokenizing the program input and output could be a challenging task.
 2.2.1. Masked Tokenization
-While our approach works with generic tokenizers, we obtain better results with custom tokenizers adapted to the languages at hand.
-Our HTMLTokenizer adapted to HTML, for instance, takes an HTML string such as
+While our approach also works with unsupervised tokenizers, such as Google SentencePiece (Kudo and Richardson, 2018), we obtain better results with custom tokenizers adapted for the languages at hand.
+For instance, our HTMLTokenizer handles an HTML string such as:
 <h1>Foobar</h1><p>Foobar is a <em>Python</em> library</p>
-The result of the tokenization looks like following:
+The result of the tokenization looks like the following:
 [ <h1>, TEXT_1, </h1>, <p>, TEXT_2, <em>, TEXT_3, </em>, TEXT_4, </p> ] { TEXT_1: Foobar, TEXT_2: Foobar is a, TEXT_3: Python, TEXT_4: library }
@@
 <h1>TEXT_1</h1><p>TEXT_2<em>TEXT_3</em>TEXT_4</p>
-The reconstruction routine is inverse to tokenization, and for many data formats, it can be performed as a simple string concatenation.
-However, Modelizer provides several implementations for mapped string reconstruction.
-We support both early and late token instantiations, meaning that abstract tokens will be replaced with real-world values before or after the token concatenation.
-Furthermore, we support different mapping policies; see Section 3.2 for details of our implementation, including the pseudocode for HTML tokenization.
-Full control over the tokenization process allows us to specify the tokenization granularity, impacting the prediction accuracy and model learning time.
-The vocabulary size grows corresponding to the number of unique tokens in the training set, which impacts the overall number of model parameters.
-It is useful to combine tokens that sequentially appear in the same context.
+Modelizer supports both early and late token instantiations, meaning that abstract tokens will be replaced with real-world values before or after the token concatenation.
+Furthermore, as well as different mapping policies;.
+See Section 3.2 for details of our implementation, including the pseudocode for HTML tokenization.
+Full control over the tokenization process allows us to control the tokenization granularity, impacting prediction accuracy and model learning time.
+The vocabulary size grows with to the number of unique tokens in the training set, which impacts the overall number of model parameters.
+It is worth combining tokens that sequentially appear in the same context.
 With this measure, we reduce
@@
     the time required for sequence generation.
-How does one implement a domain-specific tokenizer?
-The simplest way is by to extend the parsing functionality, by deriving a parser from the grammar.
-In this case, we can reuse the grammar that was used for synthesizing inputs.
-Still, we cannot infer the output parser this way yet.
-Once we can automatically and correctly extract output grammars from PUTs, we will be able to infer domain-specific parsers and tokenizers for them.
+Modelizer users can implement a domain-specific tokenizer by extending the functionality of parsers that are already available or could be derived from the grammar, for example, with the help of the Earley (Earley, 1970) parser algorithm.
+In this case, one can reuse the grammar that was used for synthesizing inputs.
+Once we can automatically and correctly extract output grammars from PUTs, we will be able to infer parsers and domain-specific tokenizers for them.
+See Section 3.2 for more information on implementing domain-specific tokenizers.
 2.3. Model Learning
-Our Learner is based on a sequence-to-sequence neural machine translation model called Transformer (Vaswani et al., 2017).
+Our Learner is based on the Transformer sequence-to-sequence neural machine translation model (Vaswani et al., 2017).
 It learns a program-specific behavior mocking model from selected input-output pairs.
 We present the high-level representation of the learning pipeline in LABEL:lst:learn and cover the steps taken in detail in the following paragraphs.
-Model Learner accepts the input-output data formats as input and loads the corresponding synthesized training sample pairs from the dataset.
+The Model Learner accepts the input-output data formats as input and loads the corresponding synthesized training sample pairs from the dataset.
 The loaded samples get automatically split into training, validation, and test sets.
@@
 Listing 3: The model learning pipeline in Modelizer (pseudocode)
-Modelizer automates the hyperparameter tuning process.
-We have implemented a Hyperparameter Optimizer that tries to automatically find the best combination of configurable parameter values using a small disjoint set of test samples.
-Such a test set consists of synthesized 10,000 samples.
-The optimizer is built upon the Optuna (Akiba et al., 2019) framework, which is a hyperparameter optimization framework that can be used for any machine learning model.
-Parameter search is performed in two phases.
-The first phase is Model Parameter Search, which looks for the best configuration of the neural network for the fixed number of iterations (Vaswani et al., 2017).
-Using a fixed learning rate, the optimizer computes the value of the cross entropy loss metric for the different combinations of the Linear Layer size, Embedding size, the number of Attention Heads, and the number of Encoder and Decoder layers.
-The combination that has the lowest value of cross-entropy criterion is then used in the Learning Rate Search phase, where the starting learning rate (LR) value, as well as weight decay and learning rate adjustment policy, are determined.
-For our experiments, we have configured the Hyperparameter Optimizer to execute the Model Parameter Search phase for 250 trials and the Learning Rate Search phase for 100 trials.
-Then the Learner gets initialized using the discovered model configuration and the model learning process gets executed for the specified number of iterations, called epochs.
-We use the cross-entropy loss function (Shannon, 1948), which is a generalization of the Logarithmic Loss function for multi-class classification problems, as our training evaluation criterion.
-It measures the dissimilarity of probability distributions between the expected and predicted labels.
-Also, we use the Adaptive Moment Estimation with Decoupled Weight Decay Regularization optimization algorithm (AdamW) (Loshchilov and Hutter, 2019) to automatically adjust the model’s parameters to minimize the training loss value.
+Modelizer additionally automates the hyperparameter tuning process.
+We have implemented an Optimizer that tries to automatically find the best combination of configurable parameter values using a small disjoint set of test samples, which, in our case, typically consists of 10,000 synthesized samples.
+The optimizer is built upon the Optuna (Akiba et al., 2019) framework, which automates hyperparameter search.
+Parameter search is conducted in two phases: Model Parameter Search and Learning Rate Search.
+During Model Parameter Search, the best neural network configuration is sought for a fixed number of iterations (Vaswani et al., 2017).
+While using a fixed learning rate, the Optimizer computes the cross-entropy loss (Shannon, 1948) metric for various combinations of Linear Layer size, Embedding size, number of Attention Heads, and Encoder-Decoder layers.
+The combination with the lowest cross-entropy value proceeds to the Learning Rate Search phase, where the initial learning rate (LR), weight decay, and learning rate adjustment policy are determined.
+In our experiments, the Optimizer executes the Model Parameter Search for 250 trials and, afterward, the Learning Rate Search for 100 trials.
+Automated hyperparameter search facilitates a faster and more effective exploration of the model configuration landscape, which is essential for optimal model performance and training efficiency.
+It saves the time and effort needed for manual tuning, allowing potential Modelizer users to focus on strategic aspects of program behavior extraction.
+After hyperparameter search, the Learner is initialized with the discovered model configuration, and learning runs for the specified epochs.
+We use the cross-entropy loss function to evaluate training, measuring the difference between expected and predicted labels.
+The AdamW optimization algorithm (Loshchilov and Hutter, 2019) adjusts model parameters to minimize training loss.
 The validation split is used to evaluate the model’s performance on unseen data to prevent overfitting.
-With our “lightweight” models, which, for example, require less than 6.3 million parameters to learn the Markdown to HTML conversion behavior of the Pandoc markup converter (MacFarlane, 2006), we achieve high accuracy in mocking program-specific behavior in a short amount of time on a commodity hardware.
-For the rest of the subjects that we used in the evaluation, the forward models contained 3-to-7.7 million parameters and inverse models required 2.8 to 10.3 million parameters to learn the hypothetical inverse behavior.
+Trained models are “lightweight”, which, for example, require less than 6.3 million parameters to learn the Markdown to HTML conversion behavior of the Pandoc markup converter (MacFarlane, 2006).
+Meanwhile, we are able to quickly and accurately mock program-specific behavior on commodity hardware.
 The training time varies from several minutes to a couple of hours, depending on the subject.
+For the subjects from the evaluation set, the forward models contained 3 to 7.7 million parameters, and inverse models required 2.8 to 10.3 million parameters to learn the hypothetical inverse behavior.
+Since leveraging the knowledge embedded in pre-trained LLMs could be beneficial for mocking program behavior, we additionally provide an adaptor for interacting with LLMs as an alternative to training task-specific models from scratch.
+We relied on the Unsloth (Han et al., 2023) framework to support resource-efficient local LLM querying and fine-tuning.
 2.4. Model Deployment
-Figure 4.
-Model Deployment.
-An example scenario of the Learned Behavior Model predicting Program Input given the Program Output.
-The model prediction gets validated for correctness in the optional validation phase.
+Figure 4. Model Deployment. An example scenario of the Behavior model predicting program input given the program output.
+Phases: (1) Input tokenization, (2) Prediction generation, (3) Output reconstruction, (4) Prediction validation.
 […]
@@
 The model prediction gets validated for correctness in the optional validation phase.
-The model Deployment phase, which is summarized in Figure 4, includes the Prediction and optional Validation phases.
+The Model Deployment phase, summarized in Figure 4, includes the Prediction and optional Validation phases.
 Depending on the source data type, which can be either program input or output, the respective model, source tokenizer, and target tokenizer get initialized.
@@
 The source data gets tokenized (1) and passed to the Program Model (2), which predicts a target token sequence.
-Then, the predicted sequence gets lifted from token representation to real-world format by the target reconstructor component (3).
+Then, the target reconstructor component (3) lifts the predicted sequence from token representation to real-world format.
 The translation routine supports an optional max output sequence length parameter.
-If the framework user does not specify this argument, our default implementation dynamically estimates the maximum output sequence length by applying a 1.25 multiplier to the length of the currently processed input sequence.
+If the framework user does not specify this argument, our default implementation dynamically estimates the hypothetical maximum output sequence length by applying a 1.25 multiplier to the length of the currently processed input sequence.
 Lastly, the predicted data can be passed to an optional Validation phase (4), which requires access to the respective PUT.
-It operates according to the model’s operation mode.
-For the behavior-mocking mode (predicting program output from a given program input), predicted output gets compared with computed output, which additionally gets tokenized and is marked as a reference sequence.
-Using both token sequences, we compute the Levenshtein distance (Levenshtein, 1965) and the BLEU (Papineni et al., 2002) score.
-Levenshtein distance, also named the edit distance in literature, is a metric for measuring the difference between two sequences, computed as a minimal number of insertions, deletions, and substitutions required to change the predicted sequence into a reference sequence.
-More on the BLEU score and other sequence similarity metrics will be covered in the next section.
-In the Inverse-mocking mode (finding the program input used to compute the given program output), the computed output collected by passing the predicted input to the PUT gets compared with the initial program output.
-Since the actual program input (ground truth) is unknown, currently we can compute sequence similarity score metrics in cases when the program accepts the predicted input and does not fail while processing it.
-A failure of course will indicate the incorrectness of the predicted input.
-It is also challenging to validate the correctness of the prediction in the inverse-mocking mode if the PUT computes multiple valid inputs to the same output.
+It aligns with the model’s operation mode.
+The predicted and computed outputs are tokenized in behavior-mocking mode (predicting output from input), with the latter treated as a reference sequence.
+Then, we calculate the sequence similarity metrics such as Levenshtein distance (Levenshtein, 1965) and the BLEU (Papineni et al., 2002) score.
+The Levenshtein distance measures the difference between sequences by counting the number of minimal edits needed to transform the predicted sequence into the reference.
+The next section will discuss the BLEU score and sequence similarity metrics in more detail.
 1def deploy(model, grammar, test_case, program_handler=None):
@@
 […]
-20 print("BLEU Score:", bleu_score, "Edit Distance:", edit_distance")
+20 print("BLEU Score:", bleu_score, "Edit Distance:", edit_distance)
 21 else:
-22 print("Model predicted", reconstructed")
+22 print("Model predicted", reconstructed)
 Listing 4: Model deployment with the help of Modelizer (pseudocode)
+In the Inverse-mocking mode (finding the program input used to compute the given program output), the computed output collected by passing the predicted input to the PUT gets compared with the initial program output.
+Since the actual program input (ground truth) is unknown, we can compute sequence similarity score metrics in cases when the program accepts the predicted input and does not fail while processing it.
+A failure, of course, will indicate the incorrectness of the predicted input.
+It is also challenging to validate the correctness of the prediction in the inverse-mocking mode if the PUT computes multiple valid inputs to the same output.
 Such programs have to be modeled by introducing an additional abstraction layer that defines the set of semantically equivalent inputs using formal rules.
-We want to demonstrate the results of the model deployment process depicted in LABEL:lst:deploy using the example from Table 1.
-We used the HTML to Markdown conversion model learned from the Pandoc program, which was trained on 1,000,000 synthesized document pairs.
-The input HTML string is passed to the model, which predicts the corresponding Markdown string.
-At first, the input HTML string is tokenized and the content gets replaced with placeholders (Step 1).
-The model uses the produced input tokens to predict a sequence of Markdown tokens (Step 2).
-During the reconstruction phase (Step 3) the placeholder tokens in the predicted sequence get replaced with the real-world data.
-Since the Pandoc program was locally installed on the same machine, which was used to run the HTML to Markdown conversion model, we can validate the correctness of the predicted output by comparing it with the computed output.
+We demonstrate the results of model deployment shown in LABEL:lst:deploy using the example from Table 1.
+In this example, we used the HTML to Markdown conversion model learned from the Pandoc program, which was trained on 1,000,000 synthesized document pairs.
+First, the input HTML string is tokenized, and the content gets replaced with placeholders (Step 1).
+The model uses the produced input tokens to predict a sequence of Markdown tokens (Step 2), and placeholders are replaced with real data during output reconstruction (Step 3).
+Since the Pandoc program was locally installed on the same machine that ran the HTML to Markdown conversion model, we can validate the correctness of the predicted output by comparing it with the computed output.
 We measured the BLEU score at 1.0 and the Levenshtein distance at 0.0, which indicates that the predicted output is identical to the computed output.
 Table 1. A conversion example from HTML to Markdown.
-Input HTML Reconstructed Markdown
+1. Input HTML 4. Reconstructed Markdown
 <p>Note that <a href="/docs/
@@
 iterate over batches.
-Input Tokens Predicted Tokens
+2. Input Tokens 3. Predicted Tokens
 <p> TEXT_1 <a  href="URL_1" >
@@
 […]
-The conversion was performed by a model of the Pandoc program trained on 1,000,000 synthesized sequence pairs. BLEU score = 1.0, Levenshtein distance = 0.0
+The conversion used a Pandoc model trained on 1,000,000 sequence pairs, achieving a BLEU score of 1.0 and a Levenshtein distance of 0.0.
 3. Implementation
-Modelizer implements grammar-based input generators using the grammar fuzzers from FuzzingBook(Zeller et al., 2023).
-While the Fuzzing Book GrammarFuzzer randomly chooses the expansion rule from the set of available rules, its ProbabilisticGrammarFuzzer allows the developer to control the input generation by assigning probabilities to certain expansion rules.
+Modelizer implements grammar-based input generators using grammar fuzzers from The Fuzzing Book (Zeller et al., 2023).
+While the Fuzzing Book GrammarFuzzer class randomly chooses the expansion rule from the set of available rules, its ProbabilisticGrammarFuzzer variant allows the developer to control the input generation by assigning probabilities to certain expansion rules.
 In this case, the non-attributed expansion rules will be assigned by equally distributed remaining probabilities.
-Such a technique allows us the following:
-  • Automatically generating a dataset of syntactically valid inputs.
-  • Customizing the feature selection procedure in the sample generation process.
-  • Ensuring the representation of specific and corner samples in the data corpus.
+This allows us to (1) automatically generatea dataset of syntactically valid inputs; (2) customize the feature selection procedure in the sample generation process; and (3) ensure the representation of specific and corner samples in the data corpus.
 3.1. System Requirements
-We implemented the model learning functionality of Modelizer in Python using the PyTorch 2.0 Machine Learning Framework (Paszke et al., 2019).
-Our framework currently supports the following hardware backends:
-  • NVIDIA GPU that supports CUDA 11.7 or higher (Fastest)
-  • Apple ARM-based system-on-a-chip with at least 16GB of system memory
-  • System with a x64 processor and 32 GB of system memory (Slowest)
-While model learning is a highly parallelizable task, and it can benefit from the availability of dedicated hardware accelerators, the rest of the tasks can also be scaled according to the number of available CPU cores.
+We implemented Modelizer model learning in Python using the PyTorch 2.0 Machine Learning Framework (Paszke et al., 2019). Our framework currently supports the following hardware backends:
+  • NVIDIA GPUs with support for CUDA 11.7 or higher (Fastest)
+  • Apple Silicon processors with at least 16GB of system memory
+  • Systems with a x64 processor and 32 GB of system memory (Slowest)
+Model learning is a highly parallelizable task, and it can benefit from the availability of dedicated hardware accelerators; however, the remaining tasks can also be scaled according to the number of available CPU cores.
 For example, to reduce the total runtime, our generators support the execution of multiple grammar-based fuzzers with the same configuration in parallel.
-The amount of Fuzzers that got initialized depends on the number of available CPU cores.
+The number of Fuzzers that were initialized depends on the number of available CPU cores.
 Modelizer is responsible for distributing the total workload among all fuzzers by assigning every fuzzer with its own budget.
-Then each fuzzer produces samples in batches until it runs out of the budget.
+Then, each fuzzer produces samples in batches until it runs out of budget.
 Modelizer systematically collects the generated samples and checks for their uniqueness.
-If there was no previously unseen sample generated within a single batch or a configurable parameter of failing attempts is reached, the value of the minimum and maximum non-terminals gets automatically incremented.
+If there was no previously unseen sample generated within a single batch or a configurable parameter of failing attempts is reached, the value of the minimum and maximum number of non-terminal expansions gets automatically incremented.
 To simplify the usability and applicability of our framework, we include a base class generator that can be extended to support new subjects.
 3.2. Tokenization with Placeholders
-Even though our models are trained on synthesized data, they must still be capable of operating with real-world data.
-However, there can be a difference in the representation between real-world and synthesized data due to the latter optionally including placeholders instead of real-world values, which impacts the implementation of our tokenizers.
-As a part of our contribution, we supplied the Modelizer with the base implementation and abstract interfaces for the mapped tokenization and mapped reconstruction routines.
-For every new data type, the framework user can wrap the functionality of unsupervised tokenizers with our mapped tokenization algorithm or implement a format-specific tokenizer in the subclass.
+Even though our models are trained on synthesized data, they must still be able to handle real-world data.
+Since synthesized data may include placeholders instead of actual values, Modelizer provides a base implementation and abstract interfaces for mapped tokenization and reconstruction routines.
+Users can wrap unsupervised tokenizer functions with our mapped tokenization algorithm or implement a format-specific tokenizer in a subclass for new data types.
 3.2.1. Abstract Tokenizers
-By assuming that for both input and output formats, a suitable parser can be obtained, we implemented AbstractTokenizer—a generic mapped tokenization-reconstruction algorithm that can benefit from format structure knowledge and replications of input content fragments in the output.
-While it is not a fully functional tokenizer, it provides a high-level generic and extendible implementation of mapped tokenization-reconstruction routines, which assist in mapping real-world content to abstract placeholders and vice-versa.
-LABEL:lst:tokenize-base contains the fragment of the mapped tokenization routine.
+Assuming the availability of suitable parsers for input and output formats, we implemented AbstractTokenizer—a generic mapped tokenization-reconstruction algorithm leveraging format structure and content fragment replications in both input and output.
+While this is not a fully functional tokenizer, it offers a high-level, extendable implementation for mapped tokenization-reconstruction tasks, aiding the mapping of real-world content to abstract placeholders and vice-versa.
+LABEL:lst:tokenize-base shows a part of the mapped tokenization logic.
+To correctly map real content to abstract placeholders (references as mask tokens in the pseudocode), the parser-tokenizer process calls the mask_token method once it handles the data fragment.
+The developer needs to implement a format-specific tokenize logic in a subclass.
+Then, Modelizer can replace content fragments with placeholders during tokenization and recover the original sequence during reconstruction.
+The tokenizer can learn placeholders from the grammar definition or refer to the user-defined ones, allowing it to handle both synthesized and real-world data.
 1class AbstractTokenizer:
@@
 8
-9 def mask_token(self, token: str, pldr: str) -> str:
-10 tag = pldr + "_" + str(len(self.mappings[pldr]) + 1)
+9 def mask_token(self, token: str, placeholder: str) -> str:
+10 tag = placeholder + "_" + str(len(self.mappings[placeholder]) + 1)
 11 if self.mapping_policy == MappingPolicy.SIMPLIFIED:
-12 self.mappings[pldr][tag] = token
-13 tag = pldr
+12 self.mappings[placeholder][tag] = token
+13 tag = placeholder
 14 elif self.mapping_policy == MappingPolicy.OPTIMIZING:
@@
 17 else:
-18 self.mappings[pldr][tag] = token
+18 self.mappings[placeholder][tag] = token
 19 self.inverse_mapping[token] = tag
 20 elif self.mapping_policy == MappingPolicy.EXHAUSTIVE:
-21 self.mappings[pldr][tag] = token
+21 self.mappings[placeholder][tag] = token
 22 self.inverse_mapping[token] = tag
@@
 Listing 5: Tokenization in Modelizer (pseudocode)
-To be able to correctly map real-world content to abstract placeholders (we use the term mask tokens, which you can see in the pseudocode) whether a parser-tokenizer processes content fragment of data the provided mask_token method must be called.
-Thus, the framework user must provide the format-specific implementation of the tokenize routine in a subclass.
-Then algorithm can correctly replace the content with an abstract placeholder during tokenization and successfully recover the original sequence of characters during the reconstruction.
-During initialization, the tokenizer, can learn the placeholders from the grammar definition automatically or operate with the user-defined collection of placeholders.
-The proposed algorithm allows reusing the same tokenizer for tokenizing both synthesized and real-world data.
 3.2.2. A HTML Tokenizer
-Below we demonstrate a partial implementation of the HTMLTokenizer class, which demonstrates the procedure of masking HTML content by extending the functionality of the AbstractTokenizer class.00footnotetext: The full implementation of HTMLTokenizer, which decomposes data using Python built-in HTMLParser, can be found in our GitHub repository (see Section 7 for details).
+HTMLTokenizer class given below demonstrates the procedure of masking HTML content by extending the functionality of the AbstractTokenizer class.
+Once the mapped_tokenize method is called, the tokenizer clears all internal buffers and starts processing the given string using the selected parser.
+The HTMLParser module from the Python Standard Library, for instance, requires the user to implement handler methods to process HTML tags and document content (handle_data method).
+Depending on the operation mode of the tokenizer, which is determined by the mask_mapping flag, the tokenizer either replaces the document content with a matching placeholder or performs a sequence of format-specific string splitting operations, which are abstracted in this example.00footnotetext: The full implementation of HTMLTokenizer, which decomposes data using Python built-in HTMLParser, can be found in our GitHub repository (see Section 7 for details).
 This repository also provides additional examples of tokenizer implementations that rely on PUT’s functionality to parse inputs or implement a custom parser using regular expressions.
@@
 Listing 6: Tokenization in Modelizer (pseudocode)
-Once the mapped_tokenize method is called, the tokenizer clears all internal buffers and starts processing the given string with the selected Parser.
-In this implementation, the HTMLParser requires the end-user to implement handler methods to process HTML tags together and document content.
-Here we present the simplified implementation of the document content handler handle_data.
-Depending on the operation mode of the tokenizer, which is determined by the mask_mapping flag, the tokenizer either replaces the document content with a matching placeholder or performs a sequence of format-specific string splitting operations, which are abstracted in this example.
 3.2.3. Token Masking Strategies
-The token masking algorithm behaves differently according to one of the selected token mapping policy strategies:
-  • The SIMPLIFIED policy reduces the vocabulary size of the model by removing identifiers from placeholder tokens.
-  • The OPTIMIZING policy, which is the default policy in our implementation, uses a minimal number of placeholder tokens to encode data content.
+The token masking algorithm behaves differently according to the selected token mapping policy strategies:
+  • The SIMPLIFIED policy reduces model vocabulary size by removing identifiers from placeholder tokens.
+  • The OPTIMIZING policy, our default, uses a minimal number of placeholder tokens to encode data.
   • The EXHAUSTIVE policy assigns every content fragment with a unique placeholder token.
-We use the extended version of the HTML document from Figure 1 to illustrate the effect of using different token masking strategies in tokenization:
+We use the extended HTML document from Figure 1 to show the impact of all token masking strategies in tokenization:
 <h1>Foobar</h1><p>Foobar is a <em>Python</em> library. Find more on <a href="127.0.0.1"><i>Foobar</i>.com</a></p>
-When the tokenizer is running in SIMPLIFIED token masking mode, then it removes every unique identifier that had been assigned to a placeholder.
+When the tokenizer is running in SIMPLIFIED token masking mode, it removes every unique identifier assigned to a placeholder.
 It significantly reduces the model’s vocabulary size, but increases the chances of producing errors during the reconstruction phase.
@@
 <a, href="URL">, <i>, TEXT, </i>, TEXT, </a>, </p> ]
-The tokenizer tries to use a minimal number of placeholders to encode the real-world data when it is configured to use OPTIMIZING token masking strategy.
-The repetitive content is encoded by the same placeholder.
-This is the default operation mode and evaluation experiments were arranged only using this policy.
+When configured with the OPTIMIZING token masking strategy, the tokenizer uses a minimal number of placeholders to encode real-world data, reusing the same placeholder for repetitive content.
+This is the default mode, and it was used for the entire evaluation.
 The tokenized sequence looks like the following:
@@
 We evaluated Modelizer by learning the execution behavior of real-world programs.
-For every PUT, we collected two datasets of input-output pairs by executing the program with synthesized and real-world inputs respectively.
+We collected two datasets of input-output pairs for every PUT by executing each program with synthesized and real-world inputs, respectively.
 For every evaluation subject, Modelizer additionally learned an inverse model  $M^{-1}$ that corresponds to a hypothetical program  $\textnormal{PUT}^{-1}$, which is expected to implement an inverse of the behavior of the original program  PUT.
+We also equipped our framework with the ability to fine-tune LLMs to capture the behavior of  PUT and its inverse  $\textnormal{PUT}^{-1}$and compared this approach against custom-trained models.
 4.1. Research Questions
@@
   1. RQ1.
-    How accurate are the Modelizer models that mock the behavior of the  PUT? This question is at the heart of Modelizer: How well do the learned models perform in predicting outputs from inputs?
+    How accurate are the Modelizer models in mocking the behavior of the  PUT and its inverse  $\textnormal{PUT}^{-1}$?
+This question is at the heart of Modelizer: How well do the learned models perform in predicting outputs from inputs and vice-versa?
   2. RQ2.
-    How accurate are the Modelizer models that are trained on synthesized data in processing real-world data?
-This question evaluates the quality of the learned models when tested with real-world data.
+    How well do the synthesized samples represent the real-world data and how much synthetic data is required for mocking program behavior?
+With RQ2, we investigate the quality of the synthesized training set, its correlation to real-world data, and investigate the impact of the size of the synthesized training set on the accuracy of the learned model.
   3. RQ3.
-    How accurate are the inverse models that mock the behavior of  $\textnormal{PUT}^{-1}$?
-With RQ3, we evaluate an important application of Modelizer, namely its ability to predict inputs from outputs.
+    How efficiently can Modelizer learn program models? This question evaluates Modelizer’s input generation and model learning efficiency.
   4. RQ4.
-    How fast Modelizer can synthesize a training dataset? This question evaluates the efficiency of the input generation pipeline implemented in Modelizer.
+    How does Modelizer learn the hyperparameters for each data pair combination?
+This question assesses the impact of hyperparameters (Section 2.3) on the performance of Modelizer.
+We also check the reusability of found hyperparameters with forward models for training inverse models.
   5. RQ5.
-    How fast Modelizer can learn program models? This question evaluates the efficiency of Modelizer core functionality, which is model learning.
+    What is the effect of tokenizer selection on learning performance? With RQ5, we investigate the impact of tokenizers (Section 2.2) on the performance of Modelizer.
   6. RQ6.
-    How well do the synthesized samples represent the real-world data? With RQ6, we investigate the quality of the synthesized training set.
-  7. RQ7.
-    How much synthetic data is required for mocking behavior at which accuracy? With RQ7, we investigate the impact of the size of the synthesized training set on the accuracy of learned models.
-  8. RQ8.
-    How should synthetic data be generated to achieve the highest accuracy of the learned models?
-With RQ8, we investigate the impact of the synthetic data generation strategies on the accuracy of learned models.
-  9. RQ9.
-    How does Modelizer learn the hyperparameters for each data pair combination? This question assesses the impact of hyperparameters (Section 2.3) on the performance of Modelizer.
-  10. RQ10.
-    Should Modelizer re-use learned hyperparameters for inverse modeling? With RQ10, we check the reusability of found hyperparameters (Section 2.3) with forward models for training inverse models.
-  11. RQ11.
-    What is the effect of placeholders on behavior learning? This question assesses the impact of placeholders (Section 2.1.2) on the performance of Modelizer.
-  12. RQ12.
-    What is the effect of tokenizers on learning performance? With RQ12, we investigate the impact of tokenizers (Section 2.2) on the performance of Modelizer.
+    Could LLMs pre-trained on code mock programs?
+With RQ6, we investigate the ability of LLMs pre-trained on code replicating program behavior using zero-shot evaluation (without task-specific model adaptation) or after additional fine-tuning and compare their performance with custom-trained models.
 4.2. Evaluation Subjects
@@
 While the library natively supports the translation in both directions(Markdown to HTML and HTML to Markdown), we have tested it only with Markdown to HTML translation functionality.
-We manually extracted a Markdown Grammar from the CommonMark Spec Version 0.29 (MacFarlane, 2019) specification.
-We used Pandoc’s functionality to convert Markdown to HTML for our case study, first, to validate the feasibility of neural behavior modeling of black-box systems, second, to check the applicability of the same technique for inverse behavior modeling, when a model has to predict the program input given the output.
+We manually extracted a Markdown Grammar from the CommonMark Spec Version 0.29 (MacFarlane, 2014) specification.
+We used Pandoc to convert Markdown to HTML in our case study to validate neural behavior modeling of black-box systems and its applicability for inverse behavior modeling, where the model predicts input from output.
 The promising results of our first experiments encouraged us to test the approach with other subjects.
 Microsoft Threat Intelligence Python Security Tools:
-(msticpy) (Hellen, 2019) is a library for InfoSec investigations, data analysis, and hunting in Jupyter Notebooks authoring for Azure Sentinel, which is a cloud-native security information and event management solution offered by Microsoft for intelligent security analytics and threat intelligence.
-Modelizer was evaluated on the SQL to KQL Conversion feature of MSTIC, which accepts SELECT SQL queries constructed using a subset of ANSI SQL-92 (ISO/IEC 9075:1992, 1992) as input and produces Kusto Query Language (KQL) queries.
-KQL is a proprietary query language developed by Microsoft, which is used for querying and analyzing data in Microsoft Azure services.
-The supported subset of SQL-92 specification was encoded as a CFG grammar by us.
+(msticpy(Hellen, 2019) is a library for InfoSec investigations, data analysis, and hunting in Jupyter Notebooks for Azure Sentinel, which is a cloud-native security information and event management solution offered by Microsoft for intelligent security analytics and threat intelligence.
+Modelizer was evaluated on MSTIC’s SQL to KQL Conversion feature, which converts SELECT SQL queries, using a subset of ANSI SQL-92 (ISO/IEC 9075:1992, 1992), into Kusto Query Language (KQL) - a proprietary query language developed by Microsoft for analyzing data in Azure Cloud.
+We encoded the supported SQL-92 subset as a CFG grammar.
 latexify-py:
@@
 py-asciimath:
-(Belotti, 2020) is a Python library for converting mathematical expressions from AsciiMath representation to Mathematical Markup Language (MathML) and representation.
-The grammar for synthesizing MathML was manually extracted from the MathML 2.0 Specification (Ausbrooks et al., 2003).
-The training pairs were collected by converting the synthesized MathML formulas first into format, and then the collected formulas were translated once more into AsciiMath format.
-This way, we managed to collect three different sets of input-output pairs and test the framework’s capabilities in performing the cross-protocol translations in directions that the given PUT does not currently support.
+(Belotti, 2020) is a Python library for converting AsciiMath expressions to MathML and formats.
+The MathML grammar was manually extracted from MathML 2.0 Specification (Ausbrooks et al., 2003).
+Training pairs were gathered by converting synthesized MathML formulas to and then to AsciiMath.
+This process yielded three sets of input-output pairs, allowing us to test cross-protocol translations in directions unsupported by the original program.
 4.3. Training Setup
@@
 Once configured, the generator has 100,000 attempts to generate unique samples using a fixed configuration.
-If the generator runs out of attempts the values of MinNonTerminals and MaxNonTerminals parameters are increased by 10 and the generator is restarted.
-The generation process will be running till the generator produces the required number of unique samples.
-Also, we configured the data generators to produce a separate set of input samples that had been synthesized with a continuously increasing complexity of the generated samples.
+If the generator runs out of attempts, the values of MinNonTerminals and MaxNonTerminals parameters are increased by 10, and the generator is restarted.
+The generation process runs till the generator produces the required number of unique samples.
+Also, we configured the generators to produce a separate set of input samples that had been synthesized with an increasing complexity of the generated samples.
 It was achieved by implementing the sliding window modification of MinNonTerminals and MaxNonTerminals values.
 In this case, the generator uses at least four different non-terminal expansion configurations to generate the required number of samples.
-Pandoc, msticpy, and py-asciimath subjects were evaluated using real-world data.
-We evaluated the SQL to KQL translation models on 1,000 samples that were collected from the “sql-create-context” open-access dataset (brianm, 2023).
-This dataset contains natural language to SQL SELECT query pairs.
-During testing the msticpy subject, we ensured that it produced a non-empty output for every extracted SQL query extracted from the dataset.
-For the py-asciimath subject, we also extracted 1000 real-world MathML formulas from the NTCIR-12 dataset (Zanibbi, 2016).
-For testing the models of the Pandoc program we extracted 12,783 top-level elements (tree-like structures nesting other markdown definitions) from 100 markdown files that were collected from the open-access Markdown dataset (Schmid, 2023) containing readme documentation files of projects hosted at the HuggingFace Hub (Face, 2024).
-We have not managed to find a dataset containing mathematical expressions written in Python using the math module from the Python Standard Library.
-Instead, we synthesized 1000 samples using our Python Expression grammar and replaced the placeholders with random values of corresponding data types.
-Additionally, we evaluated every model on 10,000 synthesized samples that were not included in the training set.
-The synthesized dataset was generated by fixing the minimal number of non-terminal expansions to 10 and the maximal number of non-terminal expansions to 20.
-While the decoder of Modelizer’s translation routines supports a configurable beam size parameter, we collected all of our experiments with a fixed beam size = 1.
-All models were trained for 10 iterations(epochs) with a context-window size equal to 5000 tokens.
-In every training configuration, 80% of the synthesized data had been used for model training.
-After every training iteration, the training quality was validated using the remaining 20% of the training set.
-Additionally, the learning quality across models was evaluated with 10,000 previously unseen data pairs which were collected using synthesized inputs, that were generated by setting the maximal number of non-terminal expansion to 20.
-We trained all models using a single NVIDIA Geforce RTX 4090 GPU with 24 GB of Memory.
-During the model learning phase Modelizer was allocating from 650 MB to 22 GB of memory depending on the configuration of hyperparameters and the sizes of collected vocabularies.
-Once trained, models can be deployed on any modern computer architecture that is equipped with at least 8 GB of system memory and can execute a Python interpreter version 3.10 or newer.
-In contrast to real-world programs the learned models do not require any external dependencies or permissions for system-wide installation.
-The required Python interpreter and the set of Python modules can be pre-configured and shipped with the model instance.
-A single environment can be used to execute multiple instances of the model in parallel.
-Because the size of trained models is relatively small, we monitored a low memory footprint when they were running in the inference mode.
+Inputs were generated on an Ubuntu Linux 22.04 workstation equipped with a 24-core AMD Ryzen Threadripper 3960X CPU and 128 GB of memory.
+Each model was evaluated on 10,000 unseen synthesized samples generated with a fixed number of non-terminal expansions between 10 and 20.
+Additionally, we evaluated models using the following datasets with real-world inputs:
+Pandoc.:
+We evaluated both Pandoc models with real-world data.
+We extracted 12,783 top-level elements (tree-like structures nesting other Markdown definitions) from 100 Markdown files that were collected from the open-access Markdown dataset (Schmid, 2023) containing README files of projects hosted at the HuggingFace Hub (Face, 2024).
+msticpy.:
+We tested msticpy models on 1,000 samples from the “sql-create-context” dataset (b mc2, 2023), featuring natural language and SQL SELECT query pairs.
+We ensured that  PUTproduced non-empty outputs for every selected SQL query.
+latexify-py.:
+The evaluation of latexify-py was challenging due to the lack of datasets with mathematical expressions written in Python using the math module from the Python Standard Library.
+Thus, we synthesized 1,000 additional samples using our Python Expression grammar and replaced placeholders with random values of the appropriate type.
+py-asciimath.:
+The evaluation of py-asciimath involved 1,000 real-world MathML formulas from the NTCIR-12 dataset (Zanibbi, 2016).
+While Modelizer’s translation routines support a configurable decoder beam size, we fixed the beam size for our experiments to 1.
+All models were trained for 10 epochs with a context window of 5000 tokens.
+In each configuration, 80% of synthesized data was used for training, and the training quality was validated with the remaining 20%.
+All models were trained on a single NVIDIA Geforce RTX 4090 GPU with 24 GB memory.
+During learning, Modelizer allocated 650 MB to 22 GB of memory based on hyperparameters and vocabulary size values.
+All models trained from scratch can be deployed on modern computers with at least 8 GB memory that can execute a Python interpreter version 3.10 or newer.
+Unlike real-world programs, these models require no external dependencies or system-wide installation permissions.
+The necessary Python interpreter and modules can be pre-configured and shipped with the model.
+A single environment can run multiple model instances in parallel due to a lower memory footprint at the inference time.
+Additionally, we experimented with querying and fine-tuning two commonly used LLMs pre-trained on code: CodeLlama (Rozière et al., 2024) and CodeGemma (Team et al., 2024).
+The selected models were instruction tuned by their vendors, meaning that vendors have fine-tuned them on natural language instruction followed by input and expected output pairs.
+This process improves the model’s general utility and performance in practical applications.
+Both models have 7 billion trainable parameters, and their weights were additionally quantized to 4bit to reduce memory requirements.
+Experiments involving the usage of LLMs were performed on the data-center grade NVIDIA A100 Tensor Core GPU with 40GB memory.
+At first, models were queried using the same real-world test samples from previous experiments.
+After that, we performed fine-tuning on synthesized data and repeated the evaluation with test samples.
+We fine-tuned an individual model for 10 epochs for each input-output pair to exclude potential interference between subjects and simplify comparison with custom-trained models.
+During fine-tuning, we reused the synthesized training sets with 10,000 samples generated using the MinNonTerminals) set to 10 and MaxNonTerminals set to 20.
 4.4. Evaluation Metrics
@@
 BLEU score:
-(bilingual evaluation understudy) (Papineni et al., 2002) is a commonly used metric for the evaluation of machine translation systems, that calculates the geometric average precision of single, double, triple, and quadruple token overlap cases between reference and hypothesis token sequences multiplied by brevity penalty.
+(bilingual evaluation understudy) (Papineni et al., 2002) is a commonly used metric for the evaluation of machine translation systems that calculates the geometric average precision of single, double, triple, and quadruple token overlap cases between reference and hypothesis token sequences multiplied by brevity penalty.
 The score output is always a number between 0 and 1, with 1 indicating the perfect translation case.
@@
 BLEU Error:
-stands for Standard Error computed using BLEU score.
+stands for the Standard Error computed using the BLEU score.
 NIST score:
@@
 Word Error Rate:
-(WER) (Morris et al., 2004) is a metric that calculates the accuracy of a model prediction by computing a ratio between the sum of token insertions, deletions, and substitutions required to transform the predicted token sequence into the reference token sequence and the total number of tokens in the reference token sequence.
-The lower the value the better the model performs with a score of 0 indicating that the model has not made any mistakes during the translation process.
-World Information Lost:
-(WIL) (Morris et al., 2004) is a metric that computes the percentage of total tokens that were incorrectly predicted by the model between sets of reference token sequences and sets of hypothesis token sequences.
-The lower the value the better the model performs with a score of 0 indicating that the model has not lost any information during the translation process.
+(WER) (Morris et al., 2004) measures prediction accuracy as a ratio of token insertions, deletions, and substitutions needed to convert the predicted token sequence into the reference sequence and the total reference token count.
+A lower WER indicates better performance, with a score of 0 showing no translation mistakes.
+Word Information Lost:
+(WIL) (Morris et al., 2004) measures the percentage of total tokens incorrectly predicted between reference and hypothesis token sequences.
+The lower the value, the better the model performs, with a score of 0 indicating that the model has not lost any information during the translation process.
 Exact Match:
-is a metric that reflects the percentage of evaluation samples that were translated correctly by the model.
-The higher the value, the better the model performs with a score of 100 indicating that the model has not made any mistakes during the translation process.
-This metric relies on the Levenshtein distance to compare the similarity between the predicted and reference sequences.
+measures the percentage of evaluation samples correctly translated by the model.
+A higher value indicates better performance, with a score of 100 showing no translation errors.
+This metric uses Levenshtein distance to assess the similarity between predicted and reference sequences.
 Close Match:
@@
 1M 0.9986 0.0033 9.9102 0.04 0.05 98.69 99.39
-100k 0.2923 0.3025 4.2689 54.2 72.13 51.72 56.03
+100k 0.2923 0.3025 4.2689 54.20 72.13 51.72 56.03
 Markdown HTML Forward 500k 0.7128 0.2337 7.5700 23.19 35.30 90.57 94.79
@@
 250k 0.2352 0.2319 4.0123 52.62 66.00 55.78 57.51
-Pandoc HTML Markdown Inverse 1M 0.7036 0.1554 7.5202 23.74 34.48 93.27 93.58
+Pandoc HTML Markdown Backward 1M 0.7036 0.1554 7.5202 23.74 34.48 93.27 93.58
 500k 0.9997 0.0035 11.0359 0.03 0.03 99.58 99.69
@@
 50k 0.9566 0.0388 5.0318 2.92 4.51 89.00 93.90
-msticpy KQL SQL Inverse 100k 0.9725 0.0379 5.1913 1.94 2.82 94.60 95.80
+msticpy KQL SQL Backward 100k 0.9725 0.0379 5.1913 1.94 2.82 94.60 95.80
 1M 0.9969 0.0543 12.9138 0.45 0.48 88.47 88.86
@@
 10k 0.8595 0.0908 10.7279 9.43 12.48 31.60 52.40
-latexify-py PyExpres. Inverse 100k 0.8747 0.1240 11.0931 10.20 12.21 44.60 56.00
+latexify-py PyExpres. Backward 100k 0.8747 0.1240 11.0931 10.20 12.21 44.60 56.00
 1M 0.9999 0.0008 11.6252 0.01 0.01 99.82 98.92
@@
 50k 0.3025 0.4990 4.4004 55.25 67.26 0.10 0.40
-AsciiMath Inverse 1M 0.6899 0.3470 8.1353 24.57 32.19 30.40 42.50
+AsciiMath Backward 1M 0.6899 0.3470 8.1353 24.57 32.19 30.40 42.50
 Outliers
@@
 500k 0.0000 0.5871 0.0171 97.16 99.69 0.00 0.20
-MathML Inverse 10k 0.5336 0.5286 4.1001 39.32 50.64 1.00 7.40
+MathML Backward 10k 0.5336 0.5286 4.1001 39.32 50.64 1.00 7.40
 1M 0.9780 0.0101 9.2108 0.92 1.74 60.81 94.01
@@
 5k 0.1028 0.6621 1.9223 74.02 84.91 0.00 0.20
-py-asciimath MathML AsciiMath Hypothet. Inverse 5k 0.2963 0.5779 3.6842 54.06 67.69 3.40 8.80
-Yellow cells indicate an evaluation run with synthetic test data, blue cells indicate evaluation runs with the real-world data, and purple cells indicate the evaluation runs on the real-world data performed after fine-tuning the baseline model on a few real-world data pairs.
-RQ1: Mocking Accuracy
+py-asciimath MathML AsciiMath Hypothet. Backward 5k 0.2963 0.5779 3.6842 54.06 67.69 3.40 8.80
+Yellow cells = evaluation with synthetic data; Blue cells = evaluation with real-world data;
+purple cells = evaluation with real-world data after few-shot fine-tuning.
+RQ1: How accurate are the Modelizer models in mocking the behavior of the  PUT and its inverse  $\textnormal{PUT}^{-1}$?
 We start with addressing RQ1: How well do Modelizer models predict outputs from inputs?
@@
 The results of the evaluation are summarized in Table 3.
-The evaluation results show that the models trained on the synthetic data can achieve a high prediction accuracy for most subjects from the evaluation set.
+They show that the models trained on the synthetic data can achieve a high prediction accuracy for most subjects from the evaluation set.
 For most models and subjects, the BLEU score is above 0.98, the Word Information Lost score is below 0.02 and the Exact Match is above 93%.
+Modelizer models mock the program behavior without making errors in more than 93% of cases.
+Next, we evaluated the accuracy of Modelizer when predicting inputs from outputs.
+The prediction accuracy scores of HTML to Markdown, KQL to SQL, to PyExpression, and AsciiMath to models depicted in Table 3 show that the inverse model achieves a similar Exact Match accuracy as their counterpart forward conversion models.
+The recorded BLEU Score values for both forward and inverse models are high and have insignificant margins between them.
+While the models of the Pandoc and msticpy library achieve more than 90% accuracy in both directions, the latexify-py and py-asciimath can produce fully correct results in half of the cases.
+Between 20% to 35% of predictions require up to 4 edit operations, including token insertions and deletions, to correct the prediction.
+Currently, Modelizer can automatically detect these failures using the prediction validation mechanism discussed in Section 2.4.
+We plan to integrate the automatic prediction repair mechanism to tackle inaccurate prediction cases in our future work.
 Table 4. Evaluation results of Markdown to HTML conversions done with the learned models of Pandoc converter.
@@
 1000k 0.2940 70.49 51.61 55.91 0.7208 36.64 90.35 94.69
-Models were trained on the various sizes of datasets containing synthesized Markdown to HTML structures.
-Blue cells indicate models trained on synthetic data generated with a fixed number of non-terminal expansions, and purple cells indicate application of the synthetic data with an increasing number of non-terminal expansions.
+Blue cells = fixed number of expansions; purple cells = increasing number of expansions.
 Table 5. Evaluation results of HTML to Markdown conversions done with the learned inverse models of Pandoc.
@@
 1000k 0.2408 64.17 55.75 57.31 0.7036 34.48 93.27 93.58
-Models were trained on the various sizes of datasets containing synthesized Markdown to HTML structures.
-Blue cells indicate models trained on synthetic data generated with a fixed number of non-terminal expansions, and purple cells indicate application of the synthetic data with an increasing number of non-terminal expansions.
-According to the prediction accuracy scores depicted in Table 3 and representative results from both Table 4 and Table 5, the medium size of training sets can already achieve a high prediction accuracy for most subjects from the evaluation set.
-Using the given grammar definitions, the data generator can produce the vast majority of data feature combinations within the 100,000–250,000 samples range.
-By further increasing the number of generated samples, while eliminating duplicates, the generator starts to produce a more unnatural combination of features, which tends to negatively affect the model accuracy, makes the learning more costly, and hits the prediction performance.
-Modelizer models trained on synthetic data have a high accuracy in predicting outputs from inputs.
-RQ2: Mocking Performance with Real Data
-To answer this question, we evaluated our models on sets of 1,000 real-world test cases per subject, except Markdown and HTML, where, as already mentioned, we took 100 real-world readme documents and extracted 12,783 top-level elements from them.
-These test sets have no intersection with any of the train sets.
-While the models achieve great accuracy when tested with synthetic data, it is noticeably lower when tested with real-world data.
-For the input generation, we did not learn the feature distribution from the particular set of test cases to avoid overfitting to the test data and check Modelizer’s capabilities in learning the program behavior while having no additional knowledge of the test subject and the data it processes.
-Consecutively, the feature distribution between the real-world and synthesized samples did not align well resulting in a significant performance hit.
-The availability of additional data pairs for the few-shot fine-tuning of the learned models can significantly boost the prediction accuracy and partially compensate for the feature distribution inconsistency at a small cost.
-After the fine-tuning, we recorded the edit distance value between expected and predicted output to be greater than 1 mostly for a small fraction of test samples.
-The inspection of these test cases has shown that they contain features that were not part of the grammar that we used to synthesize input/output pairs, thus samples with similar sets of features were not present in the training data.
-The only way to fix this issue is to populate the grammar definition with new expansion rules.
-In contrast to the majority of test data, a few test cases contained an unusual number of placeholder tokens of a particular type.
-More on this will be discussed in the next research question.
-Modelizer requires that the training data (and hence, the synthesizing grammar) covers the features found in the testing data.
-In this evaluation, all models that had been trained on MathML Formulas used as input or output (MathML to , to MathML, MathML to AsciiMath, AsciiMath to MathML) have shown the lowest prediction accuracy.
-We treated them as outliers and investigated the cause of failure.
-The reason for this is the complexity of the MathML language and the lack of the corresponding test-case features in the grammar we used for input generation.
-The MathML language supports several operators and operands, including the different encodings for the semantically same element.
-The grammar we used for the input generation covers only a fraction of these features.
-Due to this fact, the custom tokenizer also fails to parse these special cases.
-The better mapping of operators and operands during the tokenization will help to resolve this issue.
-While the MathML grammar was used for synthetic input generation and the py-asciimath library was utilized for sequential conversion of first MathML to and then to AsciiMath during the training dataset formation, the accuracy of AsciiMath to and to AsciiMath models were less impacted by the inconsistency of the vocabularies.
-The accuracy drop was only caused by the lack of the specific unseen features, while the encoding variations of the same operators got mapped to the same command by the py-asciimath library in the and AsciiMath representations, thus the model’s vocabulary had already contained them.
-However, as the Table 3 also depicts, the MathML models manage to learn the conversion behavior for the synthesized test cases.
-We believe that alterations to the grammar and the tokenization process will help to improve the accuracy of the learned models for MathML Formulas.
-In the performance evaluation, MathML is an outlier, due to the restricted grammar used for input generation.
-RQ3: Inverse Accuracy
-We turn to RQ3 and evaluate the accuracy of Modelizer when predicting inputs from outputs.
-We again refer to the prediction accuracy scores depicted in Table 3 and representative results from both Table 4 and Table 5.
-In particular, we are interested in HTML to Markdown, KQL to SQL, to PyExpression, and AsciiMath to models.
-The numbers show that the inverse model achieves a similar Exact Match accuracy, as their counterpart forward conversion models.
-The BLEU Score values that we recorded for both forward and inverse models are high and have insignificant margins between them.
-While the models of the Pandoc and msticpy library achieve more than 90% accuracy in both directions, the latexify-py and py-asciimath can produce fully correct results in half of the cases.
-Between 20% to 35% of predictions require up to 4 edit operations, including token insertions and deletions, to correct the prediction.
-Currently, Modelizer can automatically detect these failures using the prediction validation mechanism discussed in Section 2.4.
-We plan to integrate the automatic prediction repair mechanism in our future work.
-Learning inverse models is as performant as learning mock models.
-RQ4: Dataset Generation Time
+Blue cells = fixed number of expansions; purple cells = increasing number of expansions.
+Modelizer models trained on synthetic data are also accurate in predicting inputs from outputs.
+According to the representative results for Pandoc models depicted in Table 4 and Table 5, models trained on the medium training set size can already achieve a high prediction accuracy.
+Using the input generation configuration specified in Section 4.3, the data generator can produce the vast majority of input feature combinations within the 100,000–250,000 samples range.
+By further increasing the number of generated samples while eliminating duplicates, the generator starts to produce more unnatural input feature combinations, which tends to negatively affect the model accuracy by making the learning more costly and hitting the prediction performance.
+Medium datasets with 100,000 to 250,000 synthesized samples suffice to obtain good Modelizer models.
+RQ2: How well do the synthesized samples represent the real-world data and how much synthetic data is required for mocking program behavior?
+To answer this question, we evaluated our models on sets of 1,000 real-world test cases per subject, except Markdown and HTML, where, as already mentioned, we took 100 real-world readme documents and extracted 12,783 top-level elements (basic blocks) from them.
+These test sets do not intersect with any training sets.
+According to Table 3, while all models achieved high accuracy with synthetic test data, they performed noticeably worse when tested with real data.
+A probable cause for this effect is that synthesized and real-world test sets have different distributions of input-output features.
+In input generation, we intentionally avoided learning feature distribution from specific test cases to prevent overfitting and to assess Modelizer’s ability to learn program behavior without extra knowledge of the test subject or processed data.
+As a result, the feature distribution between real-world and synthesized samples did not align, causing a significant performance drop.
+This negative effect can be fixed with few-shot fine-tuning (Brown et al., 2020; Schick and Schütze, 2021), a widely used technique for the adaptation of pre-trained models to domain-specific tasks at a small cost.
+We applied it to models trained on synthetic data, which helped to mitigate feature distribution issues and enhanced prediction accuracy.
+Performance improved significantly after fine-tuning each model with 10 new randomly selected input-output pairs from source datasets.
+Purple cells in Table 3 reflect the performance numbers after few-shot fine-tuning.
 Figure 5.
-Speed of Modelizer training data generation.
-This graph represents the data generation time distribution for different subjects and training set sizes.
-The x-axis represents the input types, and the y-axis represents the time in minutes.
-Every training set size is represented by a different color.
-For every input type, the generator was producing samples with fixed and increasing complexity (partitioned).
+Markdown Sequence Length Frequency Distribution.
+The x-axis represents the length of different token sequences, and the y-axis denotes the frequency count of each token length.
+Every data split is represented by a different color.
 […]
-Speed of Modelizer training data generation.
-This graph represents the data generation time distribution for different subjects and training set sizes.
-The x-axis represents the input types, and the y-axis represents the time in minutes.
-Every training set size is represented by a different color.
-For every input type, the generator was producing samples with fixed and increasing complexity (partitioned).
-We measured the performance of our input generation pipeline in forming a tokenized training set of generated positive samples of data pairs.
-Our generators can be configured to ignore all PUT runs when PUT fails to process the synthesized input.
-In this evaluation, our generators we configured to ensure that the number of requested and produced data pairs was equal.
-Generation of MathML inputs was a special case, where the input generator was executing the PUT twice to first produce from MathML, and then AsciiMath from .
-The dataset generation time plotted on Figure 5 includes input generation with grammar fuzzer, check for the uniqueness of produced input, input processing with PUT, tokenization of both synthesized input and produced output, and flushing the formatted dataset to the persistent storage.
-As Figure 5 highlights, for the majority of subject and input generation strategy combination pairs, it required 2-to-4 minutes to synthesize 10,000 samples, 1.5 to 2 hours to produce 250k samples, and more than 6 to 8 hours to produce a dataset with 1,000,000 samples using an Ubuntu Linux workstation equipped with AMD Ryzen Threadripper 3960X 24-Core processor and 256 GB of RAM.
-The dataset generation time increases almost linearly according to the number of included records.
-We also see that input generation with an increasing number of non-terminal expansions takes from 0.3 to 2.5 times longer, and is on average 44% slower.
-Also, we want to mention that by increasing the complexity of synthesized inputs for SQL and MathML we started to get fewer input samples that PUT can successfully process.
-While this can be seen as a desired effect for stress testing the application, it significantly slows down the dataset generation process if the end goal is to form a dataset containing only positive samples.
-1.5 to 2 hours are required to produce a mid-size dataset containing 250,000 samples using a multicore processor.
-RQ5: Model Learning Time
-Figure 6. Speed of Modelizer learning models from samples with a fixed and increasing complexity (Partitioned). All model learning was done using a single NVIDIA Geforce RTX 4090 GPU.
+Markdown Sequence Length Frequency Distribution.
+The x-axis represents the length of different token sequences, and the y-axis denotes the frequency count of each token length.
+Every data split is represented by a different color.
+Figure 6.
+MathML Sequence Length Frequency Distribution.
+The x-axis represents the length of different token sequences, and the y-axis denotes the frequency count of each token length.
+Every data split is represented by a different color.
+[…]
+MathML Sequence Length Frequency Distribution.
+The x-axis represents the length of different token sequences, and the y-axis denotes the frequency count of each token length.
+Every data split is represented by a different color.
+To explain such a performance boost we examined the token sequence length frequency for data structures used in our experiments.
+As expected, the synthesized data differs from real-world data in terms of the number of tokens that these data structures contain.
+For example, the token sequence length frequency distribution for Markdown that is depicted in Figure 5 shows that the longest recorded synthetic Markdown structure has 368 tokens, with a mean length of 26 tokens.
+When it comes to real-world data, the longest processed sequence has 598 tokens, with a mean length of 12 tokens.
+Also, for MathML, whose token sequence length distribution is depicted in Figure 6, we observed a similar pattern.
+The synthetic data generator produced MathML structures that were tokenized into sequences of up to 159 tokens, while the mean sequence contained 48 tokens.
+However, real-world MathML structures from the test set contained up to 1038 tokens with a mean MathML formula consisting of 151 tokens.
+We conclude that synthetic inputs produced by Modelizer generators are shorter than test samples, which affects the model’s ability to generalize to longer real-world inputs.
+One way to solve this issue is to tweak the input generator’s configuration to produce longer sequences by recursively applying more node expansions from the input grammar.
+However, a larger expansion budget can be wasted on irrelevant expansions due to random grammar expansion picks at the input generation time.
+While this process can be guided, for example, by assigning expansion probabilities to rules of interest, it still requires prior knowledge of input feature distributions inferred from at least a few program inputs and incorporation of this knowledge into grammar(Soremekun et al., 2022).
+The generated dataset will then match the properties of seen inputs, and models trained on such a dataset will still struggle to generalize to dissimilar inputs.
+Alternatively, fine-tuning pre-trained models on real-world samples could be a more generic solution to the problem.
+It does not require the regeneration of training data followed by model re-training.
+Instead, it helps to adapt the capabilities of trained models that contain foundational knowledge about input-output relations to new conditions at a low cost.
+Sequence length discrepancy between synthetic and real-world data can be smoothed using few-shot fine-tuning.
+For a small number of test samples, fine-tuned models still struggled to accurately convert input to output and vice-versa.
+The inspection of these cases has shown that they contain unseen features that are not part of the grammar used in input synthesis, for example, features present in the input syntax extensions such as Markdown Extra (Fortin, 2003) or GitHub Flavored Markdown (GitHub Flavored Markdown Specification, 2019).
+Similarly, all models that had been trained on MathML Formulas used as model input or output (MathML to , to MathML, MathML to AsciiMath, AsciiMath to MathML) have shown the lowest prediction accuracy.
+When tested on real-world data, these models behaved as outliers because the MathML language allows multiple encodings for the semantically equivalent operators and operands, and the grammar we used for the input generation covers only a fraction of these definitions.
+This means that selected grammar does not cover all the capabilities of the  PUT, and thus, samples with such features are not present in the training data.
+If a tokenizer relies on the subject’s syntactical structure learned from grammar, it will also fail to parse these cases.
+We also evaluated the quality of the synthesized samples by checking for the inclusion of tokens collected from real-world samples into the vocabularies of learned models.
+Here are the statistics per data type used in our evaluation:
+Markdown.:
+One top-level element out of 12,783 contained a high number of URL placeholders.
+Our generators did not create sequences with over 51 URL declarations per fuzzing round using the chosen non-terminal expansion configuration.
+Consequently, these placeholders were missing from the model’s vocabulary, leading to their encoding as UNKNOWN tokens and hampering accurate output prediction.
+HTML.:
+In one top-level HTML element, we observed the absence of matching URL placeholders in the model’s vocabulary.
+Additionally, in 9 other top-level elements, we noticed the presence of extra HTML tags, like <head>, </head>, <title>, </title>, and meta attributes like the data-cites class, which were not seen during the synthetic data generation because input generator was removing HTML attributes that do not semantically affect data representation.
+In these cases, the model’s predictions were semantically correct but slightly differed syntactically from the reference sequence.
+SQL.:
+The test data contained three test cases that had been accessing more than three tables in a single query, whereas the synthesized queries lacked such complexity, limiting the model’s ability to generate matching predictions.
+KQL.:
+Two out of 1,000 queries have accessed more than three tables in a single request and used one extra feature not seen during the data generation.
+The model was unable to accurately predict the output for these queries.
+MathML Formulas.:
+In 115 cases out of 1,000 formulas, the token sequence contained the unseen operands and operators.
+Our grammar for synthesizing MathML Formulas covers only a fraction of operands and specifies a single encoding per operand type.
+Thus, the MathML tokenizer also can not parse some operands found in test data due to unseen encoding.
+An abstraction mechanism for operands, similar to numeric values and string literals, was not implemented.
+Another issue is the utilization of more than 18 identifiers in a single formula, which was uncovered by the selected input generation configuration.
+LaTeX and AsciiMath Formulas.:
+We found inconsistencies in placeholder quantities between the train and test sets, similar to MathML, with unseen operands and operators in 70 and 73 test data token sequences, respectively.
+Python and Expressions.:
+No unseen tokens were detected.
+This investigation demonstrates that the quality of learned models depends on the completeness of synthesizing grammar and the similarity of input feature distributions between training and testing sets.
+The more input features are covered in the grammar, the more program behavior is tested and leaked, and the better the model is at accurately mocking program behavior.
+A fully trained model that frequently fails to predict the output acts as a marker of the missing features in the training data and, thus, in the input grammar.
+This means that the grammar that is used in the data generation process is an estimate of the actual protocol specification, and it needs to be refined according to newly discovered features, which could also be mined automatically from the  PUT (see Section 6.10), which also involves the generation of additional samples with new features and the population of the tokenizer vocabulary with new tokens.
+Since grammar refinement is a significant change to the learning process, it also enforces either the model training restart or modifications to the embedding layer of the trained model to support new tokens and post-training on newly synthesized samples.
+Otherwise, if the grammar exactly follows the protocol specification it means that the program implementation does not.
+For example, our case study of the Pandoc markup converter with Modelizer detected deviations from the CommonMark Spec Version 0.29 (MacFarlane, 2014).
+The model accuracy significantly improved once we extended our grammar declaration with extended syntax support.
+Thus, we think that our models can assist in detecting behavior deviations from standard and protocol specifications refinement tasks since the learned model acts as a test oracle.
+Both types of failures can be automatically detected using a mechanism described in Section 2.4.
+To further cover the input feature inconsistency problem, we plotted token frequency distribution for evaluated subjects.
+In contrast to the majority of synthetic data, a few real-world test inputs contained an unusual number of certain placeholder tokens.
+For example, Figure 7 plot shows us that synthetic input generators do not generate more than 65 TEXT placeholders in a single Markdown structure, with a significant fraction of synthesized structures containing only up to 10 TEXT placeholders.
+At the same time, our analysis has shown that there are 2 Markdown structures in the real-world test set with 100 TEXT placeholders and almost 400 structures with 25 TEXT placeholders.
+Also, real-world structures contain from 4.2 to 54.8 times more alignment tokens (depending on the type) than synthetic equivalents, which are instead saturated more with blockquote elements.
+Even worse disproportion is monitored with MathML (Figure 8) where first synthetic structures contain fewer identifiers and numbers per synthesized formula than in real-world formulas, completely omit some operand types found in real-world formulas, and contain an enormous amount of MathML tag declarations mostly shaped as recursive structures consisting of opening and closing <mn>, <mrow>, <mi> tags.
+While tweaks to the number of recursive expansions performed by a grammar fuzzer suffice to improve the quality of synthesized Markdown structures and minimize the gap between synthetic and real-world documents, this will have minimal to zero effect in the case of synthesizing MathML formulas.
+MathML grammar declaration used in our experiments is a significantly more complex structure with multiple mutually-recursive rules.
+As a result, simply increasing the number of expansion cycles will generate abnormal and unrealistic structures.
+Moreover, the given grammar does not define some operators and operands frequently seen in the test samples, which means that both grammar and input generator need adjustments to be able to generate higher-quality inputs.
+These results once more conclude that the quality of learned models depends on the quality and quantity of generated inputs; thus, it is important to have both correct grammar definitions and targeted configurations of input generators.
+Modelizer requires the training data and synthesizing grammar to cover the features in the testing data.
+Figure 7.
+Markdown Token Frequency Distribution.
+The x-axis represents individual tokens, and the y-axis denotes the frequency of each token in the data split.
+Each split is represented by a different color.
+[…]
+Markdown Token Frequency Distribution. The x-axis represents individual tokens, and the y-axis denotes the frequency of each token in the data split. Each split is represented by a different color.
+Figure 8.
+MathML Token Frequency Distribution.
+The x-axis represents individual tokens, and the y-axis denotes the frequency of each token in the data split.
+Each split is represented by a different color.
+[…]
+MathML Token Frequency Distribution. The x-axis represents individual tokens, and the y-axis denotes the frequency of each token in the data split. Each split is represented by a different color.
+RQ3: How efficiently can Modelizer learn program models?
+Figure 9. Speed of Modelizer learning models from samples with a fixed and increasing complexity (Partitioned).
+All model learning experiments were done using a single NVIDIA Geforce RTX 4090 GPU.
 […]
 Speed of Modelizer learning models from samples with a fixed and increasing complexity (Partitioned). All model learning was done using a single NVIDIA Geforce RTX 4090 GPU.
-According to timing records from the model learning log, it takes Modelizer less than half an hour to learn a model from 10,000 synthesized training pairs using NVIDIA RTX 4090 GPU.
+To answer this question, we need to measure Modelizer’s learning speed on commodity hardware by repeating our learning experiments multiple times on each subject with different training set sizes and fixed subject-specific model configurationsTable 6.
+According to timing records from the model learning log, Modelizer needs less than half an hour to learn a model from 10,000 synthesized training pairs for 10 epochs using NVIDIA RTX 4090 GPU.
 The similar models will be trained within 7 to 12 hours using a mid-size dataset with 250,000 training samples.
-Lastly, it takes Modelizer 22 to 49 hours to train a model using a dataset with 1,000,000 records.
-Figure 6 demonstrates the model learning time distribution across subjects from the evaluation set, training set sizes, and input generation strategies.
+Lastly, Modelizer needs 22 to 49 hours to train a model using a dataset with 1,000,000 records.
+Figure 9 shows the model learning time distribution across subjects from the evaluation set, training set sizes, and input generation strategies.
 While an early program mock can be produced already within 30 minutes using a GPU, a significantly more accurate model will be trained within 10 hours.
-We also measured from 1% to 59% slowdown, with an average of 11%, while training models containing inputs generated with fixed or increasing complexity.
-The training time grows almost linearly with the increase of the training set size, given the automatically discovered hyperparameter configurations per subject from the evaluation set.
-Training a Modelizer model can be done on a common PC in a few hours.
-RQ6: Real-World Data
-Our next research question is RQ6: How well do the synthesized samples represent the real-world data?
-The quality of synthesized samples depends on the presence and correctness of the corresponding expansion rules that encode data-relevant features in the grammar.
-We evaluated the quality of the synthesized samples by checking for the inclusion of tokens collected from real-world samples into the vocabularies of learned models.
-We have collected the following statistics per data type used in our evaluation:
-Markdown.:
-One top-level element out of 12,783 contained a high number of URL placeholders.
-Unfortunately, our generators, given the selected non-terminal expansion configuration, have not produced sequences containing more than 51 URL declarations within 1 generation cycle.
-Thus, the corresponding placeholders were not included in the model’s vocabulary, resulting in encoding these tokens as “UNKNOWN” and the model’s inability to fully correctly predict the output.
-HTML.:
-Similar to Markdown, in one of the top-level HTML elements we have observed the absence of the same set of URL placeholders in the model’s vocabulary.
-Additionally, from 9 more top-level elements a few extra HTML tags, like <head> </head> <title> </title>, and meta attributes, like data-cites class, were not seen during the data generation.
-During the data generation, we filter out attributes that do not affect the semantics of the output data.
-But, we have not manipulated the test data.
-Thus the model prediction could be semantically correct but insignificantly differ from the reference sequence.
-SQL.:
-The test data contained 3 test cases that had been accessing more than 3 tables in a single query.
-The synthesized queries did not contain such complex samples, which resulted in the model’s inability to produce a fully matching prediction for these test cases.
-KQL.:
-2 out of 1,000 queries have accessed more than 3 tables in a single request and used 1 additional feature not seen during the data generation. The model cannot predict the output for these queries.
-MathML Formulas.:
-In 115 cases out of 1,000 formulas, the token sequence contained the unseen operands and operators.
-This problem originates from the ability of the custom tokenizer to parse and process these elements due to encoding and lack of the corresponding parser rules.
-The grammar we have defined for the synthesis of MathML Formulas does not cover all possible encodings for operators and contains only a fraction of the special operand space.
-Due to this fact, the custom tokenizer also fails to parse these special cases.
-While our tokenizer already abstracts the numeric values and string literals with the corresponding placeholders, we have not implemented the same abstraction mechanism for the special operands.
-The better encoding and mapping of operators will also help to represent such uncommon inputs without the necessity to store them in the model’s vocabulary.
-The last issue is the utilization of more than 18 identifiers within a single formula.
-To synthesize such patterns the data generation configuration has to be changed.
-and AsciiMath Formulas.:
-Similar to the MathML case we have detected the both larger than synthesized number of enumerated placeholders and the presence of the unseen operands and operators in the test data in 70 and 73 token sequences respectively.
-Python and Expressions.:
-No unseen tokens were detected.
-MathML to AsciiMath behavior mocking and vice-versa, and MathML to  behavior mocking and vice-versa are representative demonstrations of a correlation between the quality of the grammar used for input generation and the accuracy of learned models.
-The more features are covered by the grammar, the more program behavior gets tested and leaked, and the better model is capable of accurately mocking the program behavior.
-A fully trained model continuously failing to predict the output acts as a marker of the missing features in the training data, thus in the input grammar.
-The failure can be detected using a mechanism described in Section 2.4.
-At this point, two scenarios are possible:
-  • First, the grammar that has been used in the data generation process was an estimate of the actual protocol specification, and it has to be refined according to newly discovered features.
-  • Second, the grammar follows the protocol specification, but the program implementation does not.
-Then the learned model can be used as an oracle to detect the program’s behavior deviation from the protocol specification.
-In the case of predicting conversions of MathML Formulas to and AsciiMath representations, we demonstrated the necessity of the population of the grammar with new features and improvements in the tokenization process.
-Our study of the Pandoc markup converter with Modelizer detected a deviation from the CommonMark Spec Version 0.29 (MacFarlane, 2019).
-Once we extended our grammar declaration with Extended Markup Syntax, the model accuracy significantly improved.
-Thus, we think that our models can detect behavior deviations and can assist in the protocol specification refinement.
-Altogether, we can conclude that the synthesizing data is suitable to form a representative model vocabulary for the significant part of the real-world data.
-However, according to the evaluation results that we observed, the absence of feature distribution information encoded as part of grammar or mined from real-world data at the synthesized time cannot guarantee accurate model training.
-Still, the training with synthesized data allows learning basic patterns and relationships between the input and output tokens, which can be further fine-tuned on the real-world data using just a few representative examples (few-shot training) to achieve the desired accuracy.
-The performance of Modelizer depends on the accuracy and completeness of the producing grammar.
-RQ7: Usage of Synthetic Data
-Our next research question is RQ7: How much synthetic data is required for the behavior mocking of black-box programs?
-To answer these questions, we need to distribute the training set sizes into three clusters:
-Small:
-consisting of 5k, 10k, and 50k samples;
-Medium:
-containing 100k, and 250k samples; and
-Large:
-with 500k and 1M samples.
-The results are summarized in Table 4 and Table 5.
-We see that the best results for models trained only on synthetic data were achieved with Medium-sized models.
-They scored the highest number of Exact Matches while also maintaining higher values for Close Matches and BLEU scores.
-However, after a few-shot fine-tuning with real data, the best-performing models were trained with Large datasets.
-Such a pattern is representative of half of the subjects we tested.
-More aggregated statistics for the remaining subjects can be found in Table 3.
-Small and medium data sets suffice to obtain good Modelizer models.
-RQ8: Generating Synthetic Data
-Let us now go to RQ8: How should synthetic data be generated to achieve the highest accuracy of the learned models?
-To answer this question, we are going to explore the impact of the non-terminal expansion strategy on the accuracy of the learned models and refer to the representative results of Markdown to HTML conversion shown in Table 4 and HTML to Markdown conversion shown in Table 5.
-As we can see, increasing the complexity of synthesized tokens does not improve the conversion accuracy when the model is trained only on synthetic data.
-When we fine-tuned the model on a small set number of additional real-world samples that were not part of the test set, we measured an insignificant increase in prediction accuracy.
-Without providing the feature distribution by assigning expansion probabilities, we assume the fuzzing budget gets wasted on randomly chosen expansions that do not align with the feature patterns observed in the real-world data.
-We can conclude that the efficient learning strategy would be to generate the mid-size synthetic dataset with a fixed number of non-terminal expansions, which, according to our results, is sufficient to cover the feature mappings between input and output, and then perform a few-shot fine-tuning on the real-world data, if available.
-This will achieve the highest or almost the highest possible accuracy of the learned models while maintaining the lowest possible time and resource consumption.
-The highest Modelizer performance is obtained by combining a medium-sized synthetic data set with fine-tuning on real-world data.
-RQ9: Hyperparameters
-Our next research question concerns hyperparameters: How does Modelizer learn the hyperparameters for each data pair combination?
-While Modelizer makes no assumptions on the availability of the real-world data at the training time, it still can perform hyperparameter optimization.
-For that, it needs a fixed set of test cases that are used to evaluate the model’s performance by computing the cross-entropy loss function after every search trial.
-The optimizer tries to minimize the value of the loss criterion and prunes non-promising trials to further speed up the search process.
-In our experiments, we reused the same set of test cases for forward and inverse models, which allowed us to study the reusability of discovered hyperparameters for the inverse models.
-Given the behavior mocking accuracy we have observed during the evaluation of real-world data, we think that running a hyperparameter parameter search on synthesized data is a suitable approach for cases when sufficient quality and quantity of real-world inputs are not available at the training time.
-According to the results of the hyperparameter search in Table 6, the optimizer managed to find an individual configuration for every translation pair type and the resulting number of model parameters is small, ranging between 2.7 and 10 million parameters for the given subjects.
-Hyperparameter optimization is automatic and efficient.
+We also measured a slowdown from 1% to 59%, with an average of 11%, while training models containing inputs generated with fixed or increasing complexity.
+The training time grows almost linearly with the increase of the training set size, given the automatically discovered hyperparameter configurations per each subject from the evaluation set.
+Training a Modelizer model takes a few hours on a PC equipped with a commodity GPU.
+While looking for an optimal learning strategy, we studied the impact of the non-terminal expansion strategy on the accuracy of the learned models.
+As we can see in the representative results of Markdown to HTML (Table 4) conversion and HTML to Markdown (Table 5) conversion, continuously increasing the complexity of synthesized tokens does not significantly improve the conversion accuracy neither when the model is trained purely on synthetic data nor after fine-tuning on real-world samples.
+Without providing the feature distribution as grammar rule expansion probabilities, the larger expansion budget still gets wasted on randomly chosen expansions that do not align with patterns observed in the real-world data.
+We conclude that the efficient learning strategy is generating the mid-size synthetic dataset with a fixed number of non-terminal expansions, which, according to our results, is sufficient for learning feature mapping between input and output languages.
+Then, if available, we perform a few-shot fine-tuning on the real-world data.
+It achieves near-optimal conversion accuracy with low resource use–a pattern we saw in half of the subjects we tested.
+More statistics for the remaining subjects can be found in Table 3.
+Optimal Modelizer performance is achieved with a medium-sized synthetic dataset and real-world fine-tuning.
+RQ4: How does Modelizer learn the hyperparameters for each data pair combination?
 Table 6. Automatically configured hyperparameter values using the Hyperparameter Search.
@@
 AsciiMath to MathML 122 174 1 2 256 1024 64 step 0.0001 0.0100 0.0 False 3018158
-RQ10: Reusability of Hyperparameters
-In this research question, we want to check the reusability of the hyperparameters found with forward models for training the inverse models.
-To answer this question, we need to take a look at the results of Hyperparameter Search presented in Table 6.
-While the optimal embedding size value equal to 256 and the number of encoder layers equal to 1 were common for all input-output pairs, the rest of the hyperparameters were different.
-Additionally, we can mention that the best models were mainly trained with the lowest Learning Rate value; Linear Layer Size has to be at least four times larger than the Embedding Size; and, for the same data type pairs, inverse models require mostly two times more decoder layers than forward models.
-We can conclude that the hyperparameters for the forward models can serve as a lower bound for the hyperparameter search for the inverse models.
-Hyperparameters from forward models can be reused for inverse models.
-RQ11: Effect of Placeholders
-Let us now turn to the effect of design decisions, and start with RQ11: How does the inclusion of additional enumerated placeholders into the model’s input/output vocabularies impact behavior learning?
-The inclusion of the additional enumerated tokens into both model’s input/output vocabularies will impact the model’s configuration setup and increase the number of required neurons.
-Since the training data does not contain these injected tokens, the model will not be able to predict them.
-Modelizer will deactivate dead neurons, and they will not be used for the prediction.
-In cases, when the real-world samples contain the injected placeholders, the corresponding neurons still can be revived using the model fine-tuning.
-Such a technique is a cheap workaround to the problem of the incompatible feature distribution between the training and prediction inputs, when the input/output feature distribution is unknown at data generation time, or the protocol specifications that PUT implements get changed over time affecting only the quantity, but not the variety of these features.
-It helps to either completely avoid or delay the necessity to retrain the model, which is a time and resource-consuming process.
-For several subjects that we tested like bidirectional conversion of the AsciiMath formulas to the representation and vice versa, the MathML formulas to the AsciiMath representation and vice versa, MathML formulas to the representation, the quantity of the synthesized placeholders given the used configuration of non-terminal nodes expansion for data generation was insufficient to represent all symbolic values that can be found in the test set.
-The preliminary oversizing of the vocabularies with the additional placeholders helps to avoid the model’s failure to predict test cases that contain unseen placeholders.
-Thus, if the selected hardware used for model learning can tolerate the additional memory allocation caused by the inclusion of additional placeholder tokens into the model’s vocabulary, we recommend using this technique to handle the processing of previously unseen input and output data, which is encoded with additional features, in a combination with fine-tuning.
-Placeholders can significantly improve Modelizer performance.
-RQ12: Effect of Tokenizers
+Our next research question concerns hyperparameters: How does Modelizer learn the hyperparameters for each data pair combination?
+While Modelizer makes no assumptions on the availability of data at the training time, it still can perform hyperparameter optimization.
+For that, it needs a fixed set of test cases that can be generated using grammar, which is used to evaluate different model configurations over multiple trials.
+The optimizer tries to minimize the value of the loss criterion and automatically prunes non-promising configurations to further speed up the search process.
+In our experiments, we reused the same set of test cases to discover optimal configurations for both forward and inverse models, which allowed us to study the reusability of discovered hyperparameters for the inverse models as well.
+While the optimal embedding size value and the number of encoder layers were common for all input-output pairs, the rest of the hyperparameters differed.
+The best models were mainly trained with the smallest Learning Rate value; Linear Layer Size has to be at least four times larger than the Embedding Size; and, for the same data type pairs, inverse models require mostly two times more decoder layers than forward models.
+Thus, we can conclude that the hyperparameters for the forward models can serve as a lower bound for the hyperparameter search for the inverse models.
+Given the behavior mocking accuracy we have observed during the evaluation of real-world data, we think that running a hyperparameter search on synthesized data is a suitable approach for cases when sufficient quality and quantity of real-world inputs are not available at the training time.
+According to the results of the hyperparameter search in Table 6, the optimizer managed to find an individual configuration for every translation pair type, and the resulting number of model parameters is small, ranging between 2.7 and 10 million parameters for the given subjects.
+Hyperparameter optimization performed on synthetic data improves the quality of learned models.
+RQ5: What is the effect of tokenizer selection on learning performance?
 Let us turn to tokenizers.
 How should the model input/output be tokenized to achieve the best learning performance?
-Should the pre-trained tokenizers used instead of domain-specific tokenizers, considering the higher development effort and cost?
-The performance of the model is highly dependent on the quality of the generated tokens.
-The more tokens encode the same input/output data the more complex the model becomes, the more time it requires to produce the output, and the higher the chance it has to fail predicting a correct token sequence.
-Complex models require more powerful resources for functioning.
-At the same time, having a very small vocabulary also limits the model’s capability to learn a mapping between input and output correctly.
-In our work, we suggested using domain-specific tokenizers, which are developed by the domain experts and are capable of tokenizing the input/output data with the matching granularity.
-The custom tokenization pipeline gives full control over the tokenization process and allows to encoding of the input/output data in the most suitable way.
-The custom tokenizers are also capable of handling domain-specific data types and formats, which are not currently supported by the pre-trained tokenizers.
-Table 7. Comparison of the tokenization quality between implemented domain-specific tokenizers and selection of pre-trained tokenizers.
+Considering the higher development costs, should the pre-trained tokenizers be used instead of custom tokenizers?
+The model’s performance heavily relies on token quality.
+Longer token sequences necessitate the use of more complex models, which leads to longer output generation times and a rise in prediction errors.
+Such models also need more processing power, while models trained from small vocabulary have limited learning flexibility.
+We propose using custom semantically aware tokenizers developed by domain experts to achieve optimal token sequence length—vocabulary size ratios by leveraging data structure knowledge.
+Table 7. Comparison of the tokenization quality between custom-implemented tokenizers and selection of pre-trained tokenizers.
 Subject Tokenizer Average Length Vocab Size Length Ratio Vocab Ratio
@@
 AsciiMath Llama3 39 229 1.66 2.66
-The pre-trained tokenizers are shipped with the state-of-the-art NLP models BERT and Llama3.
-Yellow cells indicate an evaluation run with a custom domain-specific tokenizer, blue cells indicate the evaluation runs of the BERT tokenizer, and purple cells indicate the evaluation runs of the Llama3 Tokenizer.
-We compared the tokenization quality of the pre-trained tokenizers that are shipped with the state-of-the-art NLP models BERT (Devlin et al., 2019) and Llama3 (AI@Meta, 2024) with the domain-specific tokenizers that were implemented for each subject.
+Yellow cells = Custom Tokenizer; Blue cells = BERT Tokenizer; purple cells = Llama3 Tokenizer.
+We compared the tokenization quality of the pre-trained tokenizers that are shipped with the state-of-the-art NLP models BERT (Devlin et al., 2019) and Llama3 (AI@Meta, 2024) with the custom structure-aware tokenizers that were implemented for each subject.
 We have computed the average length of the tokenized sequences and the vocabulary size per subject.
-To simplify the comparison we derived additional metrics, the length ratio and vocabulary ratio, which are computed as the ratio of the average length and vocabulary size of the pre-trained tokenizer to the domain-specific tokenizer.
-The results of our experiments are presented in Table 7.
+To simplify the comparison, we derived additional metrics, such as the length ratio and vocabulary ratio, which are computed as the ratio of the average length and vocabulary size of the pre-trained tokenizer to the structure-aware tokenizer.
+The results of our experiments are presented in Table 7 and show that the custom structure-aware tokenizers outperform the pre-trained tokenizers in terms of the average length of the tokenized sequences.
+On the other hand, the pre-trained and unsupervised tokenizers do not require additional development effort, can tokenize multiple subjects, and are easily interchangeable.
 The evaluation was conducted on 1000 real-world samples for each subject, except for Markdown and HTML, where the tokenization was performed on all top-level elements extracted from 100 real-world documents.
 All inputs that were processed by each tokenizer were augmented by the PlaceholderProcessor module, which replaced the numeric and string literals with the matching placeholders.
-While the pre-trained tokenizers managed to formulate a smaller vocabulary for half of the evaluated subjects, the average length of the tokenized sequences was significantly higher than the one produced by the domain-specific tokenizers.
-We demonstrate an example of the tokenization of the HTML element by the domain-specific and pre-trained tokenizers in Table 8.
-In this case, the domain-specific tokenizer produces a sequence consisting of 18 tokens, BERT produces 58 tokens, and Llama3 produces 48 tokens.
-Both BERT and Llama3 split the HTML tags and placeholders into multiple sub-tokens, which increases the length of the tokenized sequence.
-The domain-specific tokenizer, on the other hand, produces a shorter sequence of tokens that are more meaningful and easier to interpret.
-The results of the evaluation in Table 7 show that the domain-specific tokenizers outperform the pre-trained tokenizers in terms of the average length of the tokenized sequences.
-The pre-trained tokenizers, on the other hand, do not require additional development effort, can tokenize multiple subjects, and are easily interchangeable.
-We still think that, currently, the domain-specific tokenizers can achieve better tokenization granularity, thus they are more suitable for better-quality mocking.
-At the same, we think that once a new type of pre-trained tokenizer that focuses on the tokenization of the program-specific data is released, program behavior modeling will become more accessible and easier to implement.
-Still, we currently suggest using the pre-trained tokenizers as a lower bound for the domain-specific tokenizers.
-By repeatedly analyzing the token vocabulary for the test set and using the knowledge of domain experts and feedback from the PUT, one can potentially quicker come up with a domain-specific tokenizer implementation that will outperform the generic pre-trained tokenizers.
-We are currently working on the automation of this process and plan to reduce the development cost and effort of such tokenizers while keeping the tokenization quality high.
+We think that custom structure-aware tokenizers can achieve better tokenization granularity and are thus better suited for program behavior mocking tasks.
+While the pre-trained tokenizers managed to formulate a smaller vocabulary for half of the evaluated subjects, the average length of the tokenized sequences was significantly higher than the one produced by the structure-aware tokenizers.
+We demonstrate an example of the tokenization of the HTML element by the structure-aware and pre-trained tokenizers in Table 8.
+In this case, the structure-aware tokenizer produces a sequence consisting of 18 tokens, BERT produces 58 tokens, and Llama3 produces 48 tokens.
+Both BERT and Llama3 split the HTML tags and placeholders into multiple sub-tokens, increasing the tokenized sequence length.
+On the other hand, the structure-aware tokenizer produces a shorter sequence of tokens that are more meaningful and easier to interpret.
+It is possible to dynamically infer a structure-aware tokenizer for many program input-output pairs by randomly mutating program inputs and analyzing changes in program execution and program output.
+We are working on automating this process to reduce the usage effort of such tokenizers while maintaining high tokenization quality.
 Table 8. Example of the tokenization of the HTML element by the domain-specific and pre-trained tokenizers.
@@
   TEXT   _   6   </   code   >   TEXT   _   7   </   p   >Ċ
-Domain-specific tokenizers (as provided by output grammars) are better than generic tokenizers.
+Structure-aware tokenizers are more suitable for behavior learning than pre-trained or unsupervised tokenizers.
+RQ6: Could LLMs pre-trained on code mock programs?
+In this research question, we examined the ability of LLMs to reproduce program behavior and compared their performance with custom domain-specific models that we have previously trained.
+At first, we decided to perform zero-shot prompting of LLMs pre-trained on code.
+However, selected LLMs, CodeLlama (Rozière et al., 2024) and CodeGemma (Team et al., 2024), were instruction fine-tuned by their vendors, so we prepended every prompt with the following instruction template:
+You are a converter program that translates {SOURCE} to {TARGET}. Translate the following data: {DATA}
+Due to the zero-shot prompting strategy, the instruction does not contain any task-specific examples but, instead, instruments the model with contextual knowledge.
+At the time of querying, {SOURCE}, {TARGET}, and {DATA} placeholders were replaced with actual values.
+We summarized the zero-shot prompting results in Table 10.
+We recorded zero or close to zero Exact Match and Close Match scores for all evaluated subjects.
+It means that selected LLMs were unable to replicate neither actual program behavior nor the behavior of the hypothetical inverse program.
+From BLEU score values, we can infer that the CodeGemma model has a limited understanding of Markdown to HTML and HTML to Markdown conversion tasks, while CodeLlama models struggle to provide any valuable results for any data pair and conversion direction combinations without additional fine-tuning.
+We provide an example of interaction with selected LLMs, including the model outputs, in Table 9.
+Despite the ability to better understand code-specific tasks due to additional pre-training, these models can not act as program stubs without additional fine-tuning.
+Table 9. An example of zero-shot prompting of LLMs pre-trained on code converting Markdown to HTML
+Model Input Expected Output CodeLlama Output CodeGemma Output
+You are a converter program
+that translates Markdown to HTML.
+Translate the following data:
+## BarkProcessor
+<h2>BarkProcessor</h2>
+### BarkProcessor.cs
+“‘csharp
+using System;
+using System.Collections.Generic;
+<h2><strong><em>
+BarkProcessor
+</em></strong></h2>
+Table 10. Evaluation results for models mimicking Pandoc, msticpy, latexify-py, and py-asciimath via LLMs pre-trained on code.
+Program Source Target
+Model
+Type
+Model BLEU
+BLEU
+Error
+NIST WER WIL
+Exact
+Match
+Close
+Match
+CodeLlama 0.1070 0.0010 1.8602 174.23 97.23 0 0
+Markdown HTML Forward CodeGemma 0.4366 0.0029 7.4377 68.24 68.82 0.61 2.17
+CodeLlama 0.0875 0.0009 1.4325 266.98 97.02 0 0
+Pandoc HTML Markdown Backward CodeGemma 0.5348 0.0028 9.1093 49.84 49.67 0 5.70
+CodeLlama 0.0542 0.0009 1.0152 662.57 93.47 0 0
+SQL KQL Forward CodeGemma 0.1546 0.0039 2.5782 250.21 84.32 0 0
+CodeLlama 0.0338 0.0010 0.6302 644.66 96.77 0 0
+msticpy KQL SQL Backward CodeGemma 0.2960 0.0043 5.0848 67.94 52.17 0 0
+CodeLlama 0.0767 0.0011 1.4521 266.39 94.65 0 0
+PyExpres. Forward CodeGemma 0.1855 0.0032 3.4188 108.77 83.83 0 0
+CodeLlama 0.0545 0.0012 0.9835 387.52 94.74 0 0
+latexify-py PyExpres. Backward CodeGemma 0.2428 0.0048 4.1169 100.26 76.04 0 0
+CodeLlama 0.0076 0.0003 0.3041 988.32 98.10 0 0
+AsciiMath Forward CodeGemma 0.0737 0.0047 2.3622 121.12 85.26 0.70 1.20
+CodeLlama 0.0077 0.0004 0.3127 894.36 98.37 0 0
+AsciiMath Backward CodeGemma 0.1231 0.0045 2.9856 117.92 82.96 0.10 0.10
+CodeLlama 0.0023 0.0002 0.2507 1026.46 98.62 0 0
+MathML Forward CodeGemma 0.0853 0.0030 2.5428 109.35 81.96 0 0
+CodeLlama 0.0037 0.0002 0.5105 155.67 99.34 0 0
+MathML Backward CodeGemma 0.0172 0.0006 0.9931 80.27 85.70 0 0
+CodeLlama 0.0078 0.0005 0.5531 154.20 99.26 0 0
+AsciiMath MathML Hypothet. Forward CodeGemma 0.0190 0.0006 1.0907 78.96 86.87 0 0
+CodeLlama 0.0023 0.0002 0.2408 1094.34 98.55 0 0
+py-asciimath MathML AsciiMath Hypothet. Backward CodeGemma 0.0412 0.0030 1.5875 140.33 88.26 0 0.50
+Blue cells = CodeLlama 7B; purple cells = CodeGemma 7B.
+Table 11. Evaluation results for models mocking Pandoc, msticpy, latexify-py, and py-asciimath behaviors queried after fine-tuning.
+Program Source Target
+Model
+Type
+Model BLEU
+BLEU
+Error
+NIST WER WIL
+Exact
+Match
+Close
+Match
+CodeLlama 0.4248 0.0033 6.4448 87.7 71.89 6.78 12.16
+Markdown HTML Forward CodeGemma 0.3463 0.0027 5.6852 82.44 76.61 0 0
+CodeLlama 0.4148 0.0031 6.4789 95.00 64.99 0 13.05
+Pandoc HTML Markdown Backward CodeGemma 0.4701 0.0022 7.9473 51.40 54.69 0 0
+CodeLlama 0.4240 0.0071 6.6428 82.81 56.16 0 3.00
+SQL KQL Forward CodeGemma 0.7035 0.0045 10.3251 22.59 25.89 0 0
+CodeLlama 0.7787 0.0055 11.0942 14.42 21.95 0 15.20
+msticpy KQL SQL Backward CodeGemma 0.0896 0.0096 1.2981 427.72 91.98 0 0
+CodeLlama 0.4270 0.0054 6.1379 59.58 52.22 0 0
+PyExpres. Forward CodeGemma 0.5250 0.0029 7.7616 27.17 39.37 0 0
+CodeLlama 0.4777 0.0103 6.7309 82.56 52.97 10.10 14.90
+latexify-py PyExpres. Backward CodeGemma 0.8204 0.0048 11.3188 16.55 18.91 0 0
+CodeLlama 0.5603 0.0085 8.3336 43.06 41.39 1.40 21.20
+AsciiMath Forward CodeGemma 0.5500 0.0080 8.2532 37.55 39.34 0 0
+CodeLlama 0.2036 0.0084 3.6731 133.84 76.99 2.00 2.60
+AsciiMath Backward CodeGemma 0.3616 0.0076 6.0105 64.78 61.27 0 0
+CodeLlama 0.2441 0.0066 4.9843 76.21 69.62 0.50 0.70
+MathML Forward CodeGemma 0.2991 0.0072 5.4391 65.28 64.26 0 0
+CodeLlama 0.5168 0.0050 4.0192 75.85 69.88 0 0
+MathML Backward CodeGemma 0.5542 0.0043 4.4655 63.05 62.49 0 0
+CodeLlama 0.5176 0.0049 3.9829 77.77 69.13 0 0
+AsciiMath MathML Hypothet. Forward CodeGemma 0.5210 0.0043 4.0966 68.45 64.73 0 0
+CodeLlama 0.2892 0.0071 5.3566 68.68 62.83 0.30 3.30
+py-asciimath MathML AsciiMath Hypothet. Backward CodeGemma 0.3590 0.0073 6.1543 52.99 57.63 0 0
+Outliers
+Pandoc Markdown HTML Forward CodeLlama 0.5645 0.0026 4.9441 73.09 60.17 15.88 71.97
+Pandoc HTML Markdown Backward CodeLlama 0.2424 0.0028 2.3148 188.1 78.01 0 40.18
+latexify-py PyExpres. Backward CodeLlama 0.5306 0.0100 5.065 73.24 46.79 15.4 20.2
+py-asciimath AsciiMath Forward CodeLlama 0.8283 0.0058 7.0132 15.54 18.09 1.9 30.4
+Blue cells = CodeLlama 7B model; purple cells = CodeGemma 7B model;
+yellow cells = CodeLlama 7B with the content part of model input being masked with placeholders.
+Table 12. Comparison of fine-tuning times (in minutes) for CodeLlama 7B and CodeGemma 7B across multiple subjects.
+Subject CodeGemma CodeLlama Custom 10k Custom 250k Custom 1M
+Markdown - HTML 304.6 296.0 26.5 407.4 1359.1
+HTML - Markdown 305.7 297.6 27.0 407.2 1354.3
+SQL - KQL 517.9 506.8 18.1 451.9 1827.4
+KQL - SQL 520.2 508.2 13.4 330.8 1322.3
+PyExpres. - LaTeX 1248.8 1278.6 23.8 619.9 2534.4
+LaTeX - PyExpres. 1244.4 1281.2 14.9 375.3 1488.2
+LaTeX - AsciiMath 113.0 119.9 25.2 642.2 2626.6
+AsciiMath - LaTeX 112.8 120.2 13.1 315.8 1318.7
+MathML - LaTeX 261.6 256.3 28.0 704.6 2904.5
+LaTeX - MathML 263.2 258.5 23.1 570.5 2314.2
+AsciiMath - MathML 258.9 253.8 18.1 446.8 1823.4
+MathML - AsciiMath 259.1 254.6 22.9 579.5 2374.0
+After zero-shot prompting, we independently fine-tuned foundational models for conversion tasks based on every subject-to-conversion direction combination using 10,000 synthesized program input-output pairs to exclude the possibility of cross-subject interference and simplify the comparison with custom subject-specific models.
+We also reused the same instruction template from zero-shot prompting experiments during the model fine-tuning and inference phases.
+We recorded significantly higher BLEU score values for all subjects, meaning that models now better understand the mapping between program input-output features, thus forming better syntactically and semantically structured predictions.
+However, Exact Match scores reflect that pre-trained LLMs cannot serve as in-place mocks for programs.
+Fine-tuning does not completely prevent models from injecting additional tokens, which is intolerable for programs expecting well-structured inputs.
+We assume that iteratively refining incorrect model predictions using additional prompting rounds and fine-tuning on more samples could help fix hallucinations of LLMs in behavior-mocking and reverse engineering tasks.
+Once we additionally masked the content part of the model input with corresponding placeholders in the same way as we did in custom-trained models, we achieved better model performance in half of the subjects with CodeLlama models.
+These cases are marked as outliers in Table 11.
+We also compared the LLM fine-tuning time with training custom models from scratch on different training sets consisting of 10,000 (small), 250,000 (medium), and 1,000,000 (large) entries.
+While custom model training was performed on a widely used Nvidia RTX 4090 high-end gaming level card, fine-tuning of quantized LLMs required a more powerful data-center grade Nvidia A100 Tensor Core GPU.
+Despite the dissimilarity in the computation capabilities of selected hardware accelerators, the results in Table 12 indicate that LLM fine-tuning on a small dataset requires time similar to training custom models from scratch using a medium-size dataset with 250,000 entries.
+In conclusion, despite our early results in the field of automated behavior extraction, we think that custom-trained models are more suitable for program behavior mocking and reverse engineering of evaluated subjects.
+We want to point out several core differences.
+Using LLMs sometimes requires processing data by third parties, which could be undesired or disallowed by terms and conditions, for example, because vendors host models themselves and only provide API access to end-users.
+Typically, proprietary models require significant computation resources, which are inaccessible to individual developers and small-sized companies.
+Depending on usage patterns, the paywall could restrict API access to proprietary models and become costly for long-term use.
+At the same time, our experiments with locally hosted, relatively small by current industry standards LLMs containing 7 billion parameters have demonstrated that such LLMs cannot replace programs in typical usage scenarios.
+Even after fine-tuning, LLMs still struggle to replicate the behavior identical to the actual program.
+In contrast, program-specific models learned by Modelizer
+  • achieve higher forward and inverse mocking accuracy;
+  • have significantly lower computation requirements; and
+  • allow for integration into existing deployment or testing scenarios.
+at a smaller cost. We foresee a big potential for the co-integration of lightweight program-specific models with large language models and large reasoning models.
+Behavior mocking with LLMs is less efficient than training custom models because they are not task-specific, require more resources, and offer similar or lower performance.
 5. Limitations
@@
 Computational Complexity.:
-Modelizer, will not be able to create models for all sorts of programs.
-The complexity of the whole program logic or its individual capabilities can be too high to be modeled by the sequence-to-sequence models.
+Modelizer can not create models for all sorts of programs.
+The program logic or its individual capabilities can be too complex to be modeled by sequence-to-sequence models.
 The transformer models themselves have a quadratic computational complexity.
 For the input sequence length $n$, hidden dimension size $h$, and number of layers in the neural network $L$, we get the following computational complexity:
-Time $\textit{Time Complexity}:O\bigl{(}L\times(n^{2}d+nd^{2})\bigr{)}$
-Space $\textit{Space Complexity}:O\bigl{(}L\times(n^{2}+nd^{2})\bigr{)}$
+[…]
 This means that modeling of extremely long input-output will be challenging and impractical.
-The models that are learned by Modelizer are not intended to replace programs.
-We see Modelizer as a tool that can help to understand the behavior of the program and assist in the program testing and debugging.
+The models that are learned by Modelizer are not intended to ubiquitously replace programs.
+We see Modelizer as a tool that can help to understand the behavior of programs and assist in program testing and debugging.
 Context Window Size.:
@@
 One of the factors that impact translation accuracy is the synthesis of training sequences.
-The grammar-based fuzzers used in the implementation of our Generators, by default, tend to produce short sequences.
-To overcome this problem we have introduced the sliding non-terminal expansion window, which partially resolves this issue.
+By default, the grammar-based fuzzers used in the implementation of our Generators tend to produce short sequences.
+To overcome this problem, we have introduced the sliding non-terminal expansion window, which partially resolves this issue.
 While the total token quantity per input and output sequences reached values similar to real-world samples, it does not guarantee a correlation for the input/output feature distributions between synthesized and real-world samples.
@@
 The implementation of such parsers is a time-consuming and error-prone task.
-The third-party pre-trained tokenizers suboptimally partition the corresponding sequences, by unnecessarily splitting atomic tokens into sub-tokens similar to the constructions that are found in natural language sequences, misinterpreting certain characters like punctuation symbols, which cover part of the format syntax.
+The third-party pre-trained tokenizers suboptimally partition the corresponding sequences by unnecessarily splitting atomic tokens into sub-tokens similar to the constructions that are found in natural language sequences, misinterpreting certain characters like punctuation symbols, which cover part of the format syntax.
 Partially, such an issue can be mitigated by defining custom post-processing routines that will refine the outcome of the tokenization phase, but, such a list of regular expressions and string replacement rules can quickly become very long, create dependencies and conflicts in the order or application, what makes them hard to maintain.
@@
 Semantic Validity of Synthesized Inputs.:
-Given the correct grammar definition, our generators which rely on grammar-based fuzzers, are very efficient in producing syntactically correct inputs.
+Given the correct grammar definition, our generators, which rely on grammar-based fuzzers, are very efficient in producing syntactically correct inputs.
 However, the PUT still can reject these inputs because grammar-based fuzzers cannot guarantee that the generated input will be semantically valid and pass all parsing guards in the PUT.
@@
 Steinhöfel et al.
-(Dominic Steinhöfel and Andreas Zeller, 2022) have proposed a technique and a formal language that allows manual specification of the semantic constraints inferred from the system specification, domain knowledge, and human expertise.
+(Steinhöfel and Zeller, 2022) have proposed a technique and a formal language that allows manual specification of the semantic constraints inferred from the system specification, domain knowledge, and human expertise.
 Their approach, which builds on a modern Z3 SMT solver, can automatically refine the semantically invalid inputs using encoded constraints.
-However, fuzzing with semantic constraints is a computationally expensive task, and it slows down the training dataset formation on a large scale.
+However, fuzzing with semantic constraints is a computationally expensive task, and it slows down the formation of training datasets on a large scale.
 One type of inputs that we synthesized, and which required such a refinement, were SQL queries with a JOIN condition.
@@
 SELECT COLUMN FROM TABLE JOIN TABLE ON TABLE.COLUMN = TABLE.COLUMN
-Once PlaceholderProcessor updates all placeholders with unique identifiers the query starts to look like this:
+Once PlaceholderProcessor updates all placeholders with unique identifiers, the query starts to look like this:
 SELECT COLUMN_1 FROM TABLE_1 JOIN TABLE_2 ON TABLE_3.COLUMN_2 = TABLE_4.COLUMN_3
@@
 To overcome this issue, we implemented an additional string refinement step, which parses the synthesized input and corrects the identifiers.
-While such an additional refinement step requires a bit of manual effort, for simple cases like this, it adds an insignificant performance penalty to the input generation process.
+While such an additional refinement step requires a bit of manual effort, it adds an insignificant performance penalty to the input generation process for simple cases like this.
 Thus, the current implementation of the Modelizer framework does not support fuzzing with constraint solving, but we plan to support this functionality in future releases.
@@
 The evolution of language models and their integration into existing projects and solutions have opened new horizons in automating routine tasks and processes (Chen et al., 2021; Ziegler et al., 2022).
-The so-called general-purpose large language models (LLMs) like Llama 2 (Touvron et al., 2023) or T5 (Raffel et al., 2020) have been adapted to process code-related tasks by fine-tuning on a huge corpus of publicly available source codes and documentation (Code Llama (Rozière et al., 2024) or CodeT5 (Wang et al., 2021)).
+The so-called general-purpose LLMs like Llama 2 (Touvron et al., 2023), Llama 3 (AI@Meta, 2024) or T5 (Raffel et al., 2020) have been adapted to process code-related tasks by fine-tuning on a huge corpus of publicly available source codes and documentation (Code Llama (Rozière et al., 2024) or CodeT5 (Wang et al., 2021)).
+The latter are pre-equipped with a deep understanding of programming syntax, idioms, and semantic nuances.
 Such optimized code-specific models can be further fine-tuned on code generation and understanding downstream tasks.
-Below we discuss the most relevant applications of neural models in the context of program behavior modeling and the challenges they face:
-Behavior Mocking.:
+Using such a transfer learning approach, these models can reduce the computation time and data quality demands compared to training a new model from scratch while enhancing the model’s ability to generalize across different programs.
+However, LLMs generally require significantly more computational resources, leading to higher operational costs and energy consumption, thus limiting their application.
+While our Learner implementation based on a lightweight Transformer model with scaled dot-product attention (Vaswani et al., 2017) mechanism is suitable and sufficient for program behavior modeling and can be trained on a wide range of existing hardware, recent LLMs require access to data-center-grade hardware accelerators.
+Still, LLMs remain the mainstream of modern software engineering, and, below, we discuss the most relevant applications of neural models in program behavior modeling and the challenges they face.
+6.1. Behavior Mocking
 There have already been previous attempts to perform behavior mocking by training sequence-to-sequence neural networks.
-TransCoder (Roziere et al., 2020) represents two sequence-to-sequence models for cross-language translations of code written in C++, Java, and Python with additional support for back translations.
+TransCoder (Rozière et al., 2020) represents two sequence-to-sequence models for cross-language translations of code written in C++, Java, and Python with additional support for back translations.
 Armengol-Estape et al.
@@
 In contrast to previous approaches that rely on sequence-to-sequence models to achieve behavior mocking, Shen.
-et al and their Konure (Shen and Rinard, 2021) system relies on the active learning approach to mine the behavior model of applications that interact with relational databases.
+et al.
+and their Konure (Shen and Rinard, 2021) system relies on the active learning approach to mine the behavior model of applications that interact with relational databases.
 Konure synthesizes inputs with database content simultaneously, monitors the traffic exchange between the database and the application, and inspects the results of the database queries to continuously refine the application behavior hypothesis.
@@
 Similar to Konure, our approach learns the program models from the input-output pairs.
-While Konure is designed to model the database interaction behavior of applications, our approach does not have such a limitation and can be used to either model the behavior of the whole system or its components and individual functionality.
-Reversing Computation.:
-Neural machine translation techniques have also shown promising results in low-level code decompilation tasks.
-Such neural decompilers are either implemented as sequence-to-sequence (Katz et al., 2018; Liang et al., 2021) sequence-to-tree (Fu et al., 2019; Cao et al., 2022) models.
-While existing works focus on converting the program’s binary representation to high-level program code, our framework allows users to reverse the direction of computation (predict the program input that was used to produce a given output) by training a behavioral model of the inverse program.
-This result is achieved by simply swapping training data pairs, i.e., using program output as model input and program input as model output.
-Even though such an inverse program may not exist, our approach still allows framework users to estimate the result of an inverse computation.
-Meanwhile, the approaches proposed by Katz et al.
-in the TRAFIX framework (Katz et al., 2019) and Fu et al., in the N-Bref framework (Fu et al., 2021), suggest training a neural decompiler from generated data.
-We think that these works are the closest to ours.
-In contrast, our framework supports behavior modeling of arbitrary programs, learning both direct and inverse models of the selected PUT, and supports further generalization of the learning process from mapping finite input-output pairs to extracting common input properties and constraints.
-Data and Model Poisoning.:
-Language models trained from crowd-sourced data suffer from data poisoning problems.
-In their research, Pearce et al.
-(Pearce et al., 2022) have studied the impact of training dataset poisoning on the quality of language model predictions.
-Code-generation models still tend to generate buggy or malicious code snippets because they have been trained on very similar corpora of publicly available documents.
+While Konure is designed purely to model the interaction between an application and a database, our approach does not have such a limitation and can be used to either model the behavior of the whole system or its components and individual functionality.
+6.2. Reversing Computation
+Neural machine translation techniques have shown promising results in low-level code decompilation tasks.
+Such neural decompilers are implemented as sequence-to-sequence (Katz et al., 2018; Liang et al., 2021) or sequence-to-tree (Fu et al., 2019; Cao et al., 2022) models.
+While existing works focus on converting a program’s binary representation to high-level code, our framework lets users reverse computation direction (predict the program input used to produce a given output) by training a model of the inverse program.
+This is done by simply swapping training data pairs, i.e., using program output as model input and program input as output.
+Even when an exact inverse program may not exist, our approach allows users to estimate the result of an inverse computation.
+Approaches by Katz et al.
+in the TRAFIX framework (Katz et al., 2019) and Fu et al.
+in the N-Bref framework (Fu et al., 2021) suggest training a neural decompiler from generated data, which we find closest to our work.
+In contrast, our framework supports behavior modeling of arbitrary programs, learning both direct and inverse models of the selected PUT.
+It also generalizes the learning process from mapping finite input-output pairs to extracting common input properties and constraints.
+6.3. Data and Model Poisoning
+Language models trained on crowd-sourced data suffer from poisoning problems.
+Pearce et al.
+(Pearce et al., 2022) studied the impact of dataset poisoning on language model predictions’ quality.
+Code-generation models still tend to generate buggy or malicious code snippets because they are trained on very similar corpora of publicly available documents.
 Schuster et al.
-(Schuster et al., 2021) have demonstrated two types of attacks on generative models: data poisoning and model poisoning.
-In the first scenario, only a few specially crafted samples, which are added to the training corpus, are sufficient to enforce code generation models to produce malicious code snippets.
+(Schuster et al., 2021) demonstrated two types of attacks on generative models: data poisoning and model poisoning.
+In the first scenario, only a few specially crafted samples added to the training corpus are enough to make code generation models produce malicious snippets.
 Alternatively, such models can be attacked by manipulating the training parameters (weights) or fine-tuning them on a few malicious samples.
-Both attack strategies will result in a deviation of generated sequences, and existing defense measures against dataset poisoning attacks cannot mitigate the problem completely (Wan et al., 2022; He and Vechev, 2023).
-Due to the possibility of manipulating model behavior, the model outcome cannot be trusted, and their application in mission-critical systems is still questionable.
-In contrast to other approaches that rely on crowd-sourced training data, our framework allows full control over the data collection and model training processes through the specification of input synthesis rules, testing of generated samples with the PUT, and filtration or sanitization of faulty, undesired, and dangerous samples.
-Domain-specific Adaptations.:
-Due to the reason that such domain-specific models were fine-tuned only on the program source codes, they cannot capture the program behavior patterns without further fine-tuning.
-Such fine-tuning could be an expensive process due to the sizes of models and specific hardware requirements.
+Both attack strategies result in a deviation of generated sequences, and existing defensive measures against dataset poisoning cannot mitigate the problem completely (Wan et al., 2022; He and Vechev, 2023).
+Due to possible manipulation of model behavior, the model outcome cannot be trusted, and their application in mission-critical systems remains questionable.
+In contrast to other approaches relying on crowd-sourced training data, our framework allows full control over data collection and model training processes by specifying input synthesis rules, testing generated samples with the PUT, and filtering or sanitizing faulty, undesired, and dangerous samples.
+6.4. Pre-trained and Unsupervised Tokenizers
+Instead of implementing tokenizers from scratch, one can potentially rely on existing generic pre-trained tokenizers.
+For example, the HuggingFace repository (Wolf et al., 2020) provides access to many community-developed natural language tokenizers through the HuggingFace API.
+Such tokenizers cannot properly split the program data for several reasons.
+First, they ignore the data-specific or computation-specific features since they cannot distinguish structure from content, which impacts the tokenization granularity.
+Second, some tokenizers, like Llama3 (AI@Meta, 2024) tokenizer, modify the input by prepending or appending additional characters to the produced tokens.
+This can break the syntactical correctness of program data and impact the quality of learned models.
+To control the tokenization process while maintaining low development effort for end users of Modelizer, we provided support to train the Google SentencePiece (Kudo and Richardson, 2018) unsupervised tokenizer for the given program program data.
+SentencePiece learns a subword vocabulary from raw input text using a specified vocabulary size; it treats the text as a continuous stream of characters.
+Once the training is complete, the tokenizer uses this learned vocabulary to break new input data into token sequences.
+By setting the vocabulary size during training, one can control both the lengths of the resulting token sequences and the number of trainable parameters in the model.
+Similar to other existing unsupervised tokenizers, SentencePiece also ignores the structural representation of data at the tokenization time.
+To reduce the size of token vocabularies in our experiments, we used custom tokenizers adapted for the subjects we were experimenting with.
+6.5. Domain-specific Adaptations
+Because such domain-specific models were fine-tuned only on the program source codes, they cannot capture the program behavior patterns without further fine-tuning.
+Our evaluation shows that LLMs pre-trained on code do not behave as drop-in replacements for existing programs.
+While humans can tolerate a certain degree of noise and errors in the LLM output, other programs cannot interpret cannot interpret them without failure.
+Additionally, their operation is significantly more costly than interactions with custom-trained models with fewer parameters and is incomparable in terms of cost-efficiency with original programs.
+Fine-tuning could be an expensive process as well due to the sizes of models and specific hardware requirements.
 Zheng et al.
 and their LlamaFactory citezheng2024llamafactory framework provide estimated memory requirements implied in fine-tuning open-source LLMs.
-For example, the small versions of commonly used LLMs, like Llama2, Llama3, and Mixtral, with 7-to-8 billion parameters require more than 60GB of memory for full-precision fine-tuning.
-Even though models are open-source, it is mostly impossible to fine-tune them locally without access to data-center-grade hardware.
-Instead, our framework can learn “lightweight” program behavior models consisting of only several million parameters and can be trained on commodity hardware.
-Neural Program Induction.:
+For example, the small versions of commonly used LLMs, like Llama2, Llama3, and Mixtral, with 7-to-8 billion parameters, require more than 60GB of memory for full-precision fine-tuning.
+Even though models are open-source, it is mainly impossible to fine-tune them locally without access to data-center-grade hardware.
+To mitigate these memory requirements problems, one can use a post-training model weights quantization(Krishnamoorthi, 2018; Banner et al., 2019) technique that reduces the precision of model weights, thus requiring a lower number of bits to store them.
+Also, the low-rank adaptors(Hu et al., 2021) technique can further optimize computation and storage efficiency by introducing low-rank matrices into specific layers of LLMs like the attention mechanism to capture task-specific information by updating only these matrices during fine-tuning.
+However, model weight quantization can tamper with the prediction quality due to quantization errors, which may significantly impact performance in learning uncommon behavior patterns.
+Similarly, LoRA adapters might limit the model’s expressiveness and adaptability by constraining updates only to low-rank matrices.
+Instead, our framework can learn “lightweight” program behavior models consisting of only several million parameters and can be trained on commodity hardware at a low cost.
+6.6. Neural Program Induction
 We find that the neural program induction and neural program synthesis fields are also closely related to our research.
@@
 Recent advances in sequence-to-sequence prediction algorithms (Transformer architecture and its variations) make it possible to automatically infer complex data patterns from input-output pairs.
-Moreover, our approach can also automatically learn the inverse patterns, which allows the revision of the computation results of certain programs at a small cost.
-AI-driven Testing.:
-AI-driven, or specifically LLM-driven software testing is becoming more and more popular.
-We picked a few representative works from myriads of publications in this field to highlight the s TOGA (Dinella et al., 2022) project explores neural test oracle generation problem by fine-tuning the CodeBERT (Feng et al., 2020) model on automated inference of exceptional and assertion oracles from developer-written tests.
+Moreover, our approach can also automatically learn the inverse patterns, which allows for the revision of certain programs’ computation results at a low cost.
+6.7. AI-driven Testing
+AI-driven, or specifically LLM-driven software testing, is becoming more and more popular.
+We picked a few representative works from myriads of publications in this field to highlight common use case scenarios.
+For example, the TOGA (Dinella et al., 2022) project explores neural test oracle generation problem by fine-tuning the CodeBERT (Feng et al., 2020) model on automated inference of exceptional and assertion oracles from developer-written tests.
 Haluptzok et al.
@@
 The input sequence is represented by a sequence of characters in PDF objects, while the output sequence is obtained by shifting the corresponding input sequence by one character.
-While every mentioned work follows its unique approach, they all require a significant amount of human effort for problem-solving.
+While each work mentioned follows its unique approach, they all require a significant amount of human effort to solve problems.
 In contrast, Modelizer tries to automate the process of data collection (which is one of the most time-consuming parts of model preparation) and model training.
 The minimalistic set of requirements, in particular the ability to run a Python interpreter locally, simplifies the integration of the learned models into the existing projects and testing tools.
-AI-generated Synthetic Data:
+6.8. AI-generated Synthetic Data
 The capabilities of LLMs revealed several approaches of their application for synthetic data generation (Li et al., 2023).
@@
 Such AI-generated synthetic data is mostly used for training classifiers or question-answering systems.
-For example, general-purpose LLMs like OpenAI ChatGPT or Google Bard have shown promising results (İrfan AYGÜN, 2023) in generating synthetic data in a medical text production task.
+For example, general-purpose LLMs like OpenAI ChatGPT or Google Bard have shown promising results (İrfan Aygün and Kaya, 2023) in generating synthetic data in a medical text production task.
 However, in certain tasks, like hateful speech detection and classification, the inclusion of syntactically generated samples to the training set does not improve the performance of classifiers (Kruschwitz and Schmidhuber, 2024).
-Still, it is an active and promising field of research, and industry giants like Nvidia invest in the development of LLMs, like Nemotron-4 (Nvidia et al., 2024), specifically designed for synthetic data generation.
-While synthetic data generation with LLMs implies specific hardware requirements to run models locally or lead to a privacy compromise while processing the data with a third-party cloud provider at an additional service cost, our approach that relies on synthesizing data with grammar fuzzers only requires a system with an ability to run a Python interpreter locally.
-Data Generation by Mutating Inputs:
+Still, it is an active and promising field of research, and industry giants like Nvidia invest in the development of LLMs, like Nemotron-4 (Adler et al., 2024), specifically designed for synthetic data generation.
+While synthetic data generation with LLMs implies specific hardware requirements to run models locally or leads to a privacy compromise while processing the data with a third-party cloud provider at an additional service cost, our approach, which relies on synthesizing data with grammar fuzzers, only requires a system with the ability to run a Python interpreter locally.
+6.9. Data Generation by Mutating Inputs
 We also want to address the alternatives to grammar-based and AI-based input generation strategies.
-Potentially, an input set for the given PUT can be automatically constructed with the help of mutation-based fuzz-testing (Barton P.
-Miller et al., 1990).
-A mutation fuzzer, like AFL++ (Andrea Fioraldi et al., 2020), can accept one input for a PUT as an initial seed and produce infinitely many new inputs by applying pre-defined mutation strategies, like bit flips, addition, or subtraction of integers, on different parts of the given seed.
+Potentially, an input set for the given PUT can be automatically constructed with the help of mutation-based fuzz-testing (Miller et al., 1990).
+A mutation fuzzer, like AFL++ (Fioraldi et al., 2020), can accept one input for a PUT as an initial seed and produce infinitely many new inputs by applying pre-defined mutation strategies, like bit flips, addition, or subtraction of integers, on different parts of the given seed.
 Such a technique produces two types of mutants.
 The first type of mutants has a broken syntactical or semantic structure.
-These mutants will either crash the PUT if it contains a bug in its implementation, or the input parser guards will reject such input, and program execution will be terminated.
+These mutants will either crash the PUT if it contains a bug in its implementation, or the input parser guards will reject such input, terminating program execution.
 The collected input-output pairs will only cover the behavior of the input parser rather than the whole program.
@@
 Thus, testing the program with this kind of input will unlikely produce a diverse set of outputs.
-While fuzz testing is helpful for bug discovery in the implementations of input parsers, we do not think the mutation-based input generation technique is applicable to automate program behavior modeling.
-Mining Input Grammars:
+While fuzz testing is helpful for bug discovery in input parser implementations, we do not think the mutation-based input generation technique is applicable to automate program behavior modeling.
+6.10. Mining Input Grammars
 In certain cases, input grammars can also be automatically mined from a PUT.
 The Mimid approach by Gopinath et al.
-(Rahul Gopinath et al., 2020) relies on the program instrumentation in combination with string wrappers to identify the respective control flow nodes that access the relevant portions of a program input.
+(Gopinath et al., 2020b) uses program instrumentation with string wrappers to identify control flow nodes accessing relevant portions of a program input.
 Kulkarni et al.
-presented the Arvada algorithm (Kulkarni et al., 2021) that can automatically learn a highly recursive CFG from a given program using a set of positive examples and a Boolean-valued oracle.
-Currently, if input specifications are not publicly available or are not recoverable from the PUT, then the user of our framework has to encode it in CFG format manually.
+presented the Arvada algorithm (Kulkarni et al., 2021) that can automatically learn a highly recursive grammar from a given program using positive examples and a Boolean-valued oracle.
+The GDBMiner (Eisele et al., 2025) tool by Eisele et al.
+can also automatically extract input grammars from binaries and executables written in any compiled language on any system architecture that can execute a GDB debugger.
+Finally, symbolic parsing (Bettscheider and Zeller, 2024) by Bettscheider et al.
+promises to statically extract accurate input grammars from parser code, without requiring any sample inputs.
+Currently, if input formats are not publicly available or are not recoverable from the PUT, the user of our framework must specify them as a grammar.
+Once an input grammar is available, Modelizer can generate inputs and learn the program behavior.
 7. Conclusion and Future Work
 With Modelizer, we provide a principled approach to turning regular programs into models.
-This opens up the manifold capabilities of machine learning models for several Software Engineering tasks such as mocking, testing, and debugging.
-The maybe most promising feature of the learned translation models is their ability to predict inputs for given outputs, not only reversing executions but also offering new opportunities for test input generation.
+This opens up the manifold capabilities of machine learning models for several Software Engineering tasks such as mocking, testing, monitoring, and debugging.
+Program models learned from synthetic inputs effectively represent real-world program execution scenarios by learning basic patterns from program input-output relations.
+While the lack of knowledge of the end data feature distribution could limit accurate model training from scratch, few-shot fine-tuning on real-world examples efficiently enhances prediction accuracy.
+The most promising feature of the learned translation models is their ability to predict inputs for given outputs, not only reversing executions but also offering new opportunities for test input generation.
 We see Modelizer as a pioneer in project-specific learning—rather than trying to generate expensive LLMs that may or may not evolve into a one-size-fits-all, Modelizer demonstrates that it is possible to learn small models for nontrivial programs that can successfully replicate their behavior—and their reverse.
@@
 By treating traces or coverage as output of the PUT (and supplying an output grammar that can parse trace logs), Modelizer can learn which input features correlate with which traces or coverage features; and can then again predict which input would be required to achieve a particular trace or coverage.
-These inputs could then be tested automatically, and if the trace or coverage deviates, such feedback could be used to fine-tune the model (see below).
+These inputs could then be tested automatically, and if the trace or coverage deviates, such feedback could be used to fine-tune the model.
 Monitoring.:
-Once Modelizer has extracted a model, one can also use it for monitoring the PUT, continuously comparing the program behavior against the extracted model.
+Once Modelizer has extracted a model from the reference program, one can also use it for monitoring revisions of PUT by continuously comparing the program behavior against the extracted model.
 This could be used to detect and prevent behavior not seen during testing.
@@
 Smarter test generation.:
-In our evaluation, we have found again and again that the performance of Modelizer very much depends on the quality and diversity of the synthesized training data, and thus on the quality of the test generator.
+In our evaluation, we repeatedly found evidence that Modelizer’s performance greatly depends on the quality and diversity of synthesized training data and, thus, on the quality of the test generator.
 In our future work, we will experiment with smarter black-box test generators, covering grammar features and combinations thereof (Havrikov and Zeller, 2020), but also smarter grey-box fuzzing, using code coverage in the PUT as guidance (Pham et al., 2019; Aschermann et al., 2019).
 Stateful programs.:
-At this point, the models extracted by Modelizer do not consider the internal state, as, say, a server could maintain.
+At this point, the models extracted by Modelizer do not consider internal state, as, say, a server might maintain.
 Let us assume the PUT starts in a known initial state, and the input to the PUT is a sequence of commands.
-Then, having Modelizer learn the effects of command sequences should also give it the ability to detect that certain outputs depend on specific subsequences of commands (implying a specific state).
+Then, having Modelizer learn the effects of command sequences should enable it to detect that certain outputs depend on specific subsequences of commands (implying a specific state).
 Such reasoning will likely require larger models; we plan to evaluate Modelizer on more complex and stateful programs.
@@
 Leveraging program feedback.:
-We estimate that the model’s performance can be further improved by utilizing the feedback from the PUT, when available.
+We estimate that the model’s performance can be further improved by utilizing the feedback from the PUT when available.
 In contrast to fundamental LLMs, Modelizer framework allows relatively cheap continuous re-training of the learned models.
-If the specifications of the target system have insignificantly changed, only affecting the system behavior, but not extending the population of the input-output vocabularies with new elements, the learned models can be fine-tuned with a few samples collected from system executions.
+If the specifications of the target system have insignificantly changed, only affecting the system behavior but not extending the population of the input-output vocabularies with new elements, the learned models can be fine-tuned with a few samples collected from system executions.
 In Section 2.4, we have represented a scenario when the PUT serves as an oracle to validate the model’s predictions.
 Once the model prediction does not align with execution results, the collected feedback can be passed to the model as additional training data to adjust the model’s parameters.
-The model has to be repeatedly re-evaluated and fine-tuned till the model’s predictions match the PUT’s execution results.
+The model has to be repeatedly re-evaluated and fine-tuned until its predictions match the PUT’s execution results.
 We plan to study the capabilities of Modelizer and respective learned models to serve as an oracle for automatic detecting of the behavior deviation cases between various revisions and implementations of the same protocol.
-Also, we hope to explore the possibility of forging feedback-driven tokenizers, which will be able to automatically adjust the tokenization granularity and possibly mine the input-output grammars from black-box systems.
+We also hope to explore the possibility of forging feedback-driven tokenizers, which will be able to automatically adjust the tokenization granularity and possibly mine the input-output grammars from black-box systems.
 They will rely not only on the feedback from the PUT but also leverage the feedback from the model evaluation.
-1-bit models.:
-Lastly, we are excited by the idea of replacing 16-bit parameter weights with a 1-bit representation (Ma et al., 2024), which significantly reduces the memory requirements, and thus allows training more cost-effective models.
-We plan to integrate this feature into Modelizer soon.
-Modelizer and all experimental data are available at
+Modelizer and all experimental data, including interactive plots, are available at
 https://github.com/2ral/Modelizer
@@
 This work is funded by the European Union (ERC S3, 101093186).
-Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council.
+Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council.
 Neither the European Union nor the granting authority can be held responsible for them.